The GNU Mach Reference Manual

The GNU Mach Reference Manual

Marcus Brinkmann

with

Gordon Matzigkeit, Gibran Hasnaoui,

Robert V. Baron, Richard P. Draves, Mary R. Thompson, Joseph S. Barrera

Edition 0.3

last updated 2001-09-01

for version 1.2 Copyright \begin{matrix}c \\ ◯ \end{matrix} 2001 Free Software Foundation, Inc. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with the Invariant Sections being "Free Software Needs Free Documentation" and "GNU Lesser General Public License", the Front-Cover texts being (a) (see below), and with the Back-Cover Texts being (b) (see below). A copy of the license is included in the section entitled "GNU Free Documentation License".

(a) The FSF’s Front-Cover Text is: A GNU Manual

(b) The FSF’s Back-Cover Text is: You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development. This work is based on manual pages under the following copyright and license: Mach Operating System Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991,1990 Carnegie Mellon University All Rights Reserved. Permission to use, copy, modify and distribute this software and its documentation is hereby granted, provided that both the copyright notice and this permission notice appear in all copies of the software, derivative works or modified versions, and any portions thereof, and that both notices appear in supporting documentation. CARNEGIE MELLON ALLOWS FREE USE OF THIS SOFTWARE IN ITS "AS IS" CONDITION. CARNEGIE MELLON DISCLAIMS ANY LIABILITY OF ANY KIND FOR ANY DAMAGES WHATSOEVER RESULTING FROM THE USE OF THIS SOFT-WARE.

1 Introduction

GNU Mach is the microkernel of the GNU Project. It is the base of the operating system, and provides its functionality to the Hurd servers, the GNU C Library and all user applications. The microkernel itself does not provide much functionality of the system, just enough to make it possible for the Hurd servers and the C library to implement the missing features you would expect from a POSIX compatible operating system.

1.1 Audience

This manual is designed to be useful to everybody who is interested in using, administering, or programming the Mach microkernel.

If you are an end-user and you are looking for help on running the Mach kernel, the first few chapters of this manual describe the essential parts of installing and using the kernel in the GNU operating system.

The rest of this manual is a technical discussion of the Mach programming interface and its implementation, and would not be helpful until you want to learn how to extend the system or modify the kernel.

This manual is organized according to the subsystems of Mach, and each chapter begins with descriptions of conceptual ideas that are related to that subsystem. If you are a programmer and want to learn more about, say, the Mach IPC subsystem, you can skip to the IPC chapter (see Chapter 4 [Inter Process Communication], page 13), and read about the related concepts and interface definitions.

1.2 Features

GNU Mach is not the most advanced microkernel known to the planet, nor is it the fastest or smallest, but it has a rich set of interfaces and some features which make it useful as the base of the Hurd system.

it’s free software Anybody can use, modify, and redistribute it under the terms of the GNU General Public License (see Appendix A [Copying], page 111). GNU Mach is part of the GNU system, which is a complete operating system licensed under the GPL.

it’s built to survive As a microkernel, GNU Mach doesn’t implement a lot of the features commonly found in an operating system, but only the bare minimum that is required to implement a full operating system on top of it. This means that a lot of the operating system code is maintained outside of GNU Mach, and while this code may go through a complete redesign, the code of the microkernel can remain comparatively stable.

it’s scalable Mach is particularly well suited for SMP and network cluster techniques. Thread support is provided at the kernel level, and the kernel itself takes advantage of that. Network transparency at the IPC level makes resources of the system available across machine boundaries (with NORMA IPC, currently not available in GNU Mach).

it exists The Mach microkernel is real software that works Right Now. It is not a research or a proposal. You don’t have to wait at all before you can start using and developing it. Mach has been used in many operating systems in the past, usually as the base for a single UNIX server. In the GNU system, Mach is the base of a functional multi-server operating system, the Hurd.

1.3 Overview

An operating system kernel provides a framework for programs to share a computer’s hardware resources securely and efficiently. This requires that the programs are seperated and protected from each other. To make running multiple programs in parallel useful, there also needs to be a facility for programs to exchange information by communication.

The Mach microkernel provides abstractions of the underlying hardware ressources like devices and memory. It organizes the running programs in tasks and manages the threads (points of execution in the tasks). In addition, Mach provides a rich interface for interprocess communication.

What Mach does not provide is a POSIX compatible programming interface. In fact, it has no understanding of file systems, POSIX process semantics, network protocols and many more. All this is implemented in tasks running on top of the microkernel. In the GNU operating system, the Hurd servers and the C library share the responsibility to implement the POSIX interface, and the additional interfaces which are specific to the GNU system.

1.4 History

XXX History of Mach here.

2 Installing

Before you can use the Mach microkernel in your system you’ll need to install it and all components you want to use with it, e.g. the rest of the operating system. You also need a bootloader to load the kernel from the storage medium and run it when the computer is started.

GNU Mach is only available for Intel i386-compatible architectures (such as the Pentium) currently. If you have a different architecture and want to run the GNU Mach microkernel, you will need to port the kernel and all other software of the system to your machine’s architecture. Porting is an involved process which requires considerable programming skills, and it is not recommended for the faint-of-heart. If you have the talent and desire to do a port, contact bug-hurd@gnu.org in order to coordinate the effort.

2.1 Binary Distributions

By far the easiest and best way to install GNU Mach and the operating system is to obtain a GNU binary distribution. The GNU operating system consists of GNU Mach, the Hurd, the C library and many applications. Without the GNU operating system, you will only have a microkernel, which is not very useful by itself, without the other programs.

Building the whole operating system takes a huge effort, and you are well advised to not do it yourself, but to get a binary distribution of the GNU operating system. The distribution also includes a binary of the GNU Mach microkernel.

Information on how to obtain the GNU system can be found in the Hurd info manual.

2.2 Compilation

If you already have a running GNU system, and only want to recompile the kernel, for example to select a different set of included hardware drivers, you can easily do this. You need the GNU C compiler and MiG, the Mach interface generator, which both come in their own packages.

Building and installing the kernel is as easy as with any other GNU software package. The configure script is used to configure the source and set the compile time options. The compilation is done by running:

     make

To install the kernel and its header files, just enter the command:

     make install

This will install the kernel into $(prefix)/boot/gnumach and the header files into $(prefix)/include. You can also only install the kernel or the header files. For this, the two targets install-kernel and install-headers are provided.

2.3 Configuration

The following options can be passed to the configure script as command line arguments and control what components are built into the kernel, or where it is installed.

The default for an option is to be disabled, unless otherwise noted.

--prefix prefix

Sets the prefix to PREFIX. The default prefix is the empty string, which is the correct value for the GNU system. The prefix is prepended to all path names at installation time.

--enable-kdb

Enables the in-kernel debugger. This is only useful if you actually anticipate debugging the kernel. It is not enabled by default because it adds considerably to the unpageable memory footprint of the kernel.

--enable-kmsg

Enables the kernel message device kmsg.

--enable-lpr

Enables the parallel port devices lpr%d.

--enable-floppy

Enables the PC floppy disk controller devices fd%d.

--enable-ide

Enables the IDE controller devices hd%d, hd%ds%d.

The following options enable drivers for various SCSI controller. SCSI devices are named sd%d (disks) or cd%d (CD ROMs).

--enable-advansys

Enables the AdvanSys SCSI controller devices sd%d, cd%d.

--enable-buslogic

Enables the BusLogic SCSI controller devices sd%d, cd%d.

--disable-flashpoint

Only meaningful in conjunction with ‘--enable-buslogic’. Omits the Flsh-Point support. This option is enabled by default if ‘--enable-buslogic’ is specified.

--enable-u1434f

Enables the UltraStor 14F/34F SCSI controller devices sd%d, cd%d.

--enable-ultrastor

Enables the UltraStor SCSI controller devices sd%d, cd%d.

--enable-aha152x
--enable-aha2825

Enables the Adaptec AHA-152x/2825 SCSI controller devices sd%d, cd%d.

--enable-aha1542

Enables the Adaptec AHA-1542 SCSI controller devices sd%d, cd%d.

--enable-aha1740

Enables the Adaptec AHA-1740 SCSI controller devices sd%d, cd%d.

--enable-aic7xxx

Enables the Adaptec AIC7xxx SCSI controller devices sd%d, cd%d.

--enable-futuredomain

Enables the Future Domain 16xx SCSI controller devices sd%d, cd%d.

--enable-in2000

Enables the Always IN 2000 SCSI controller devices sd%d, cd%d.

--enable-ncr5380
--enable-ncr53c400

Enables the generic NCR5380/53c400 SCSI controller devices sd%d, cd%d.

--enable-ncr53c406a

Enables the NCR53c406a SCSI controller devices sd%d, cd%d.

--enable-pas16

Enables the PAS16 SCSI controller devices sd%d, cd%d.

--enable-seagate

Enables the Seagate ST02 and Future Domain TMC-8xx SCSI controller devices sd%d, cd%d.

--enable-t128
--enable-t128f
--enable-t228

Enables the Trantor T128/T128F/T228 SCSI controller devices sd%d, cd%d.

--enable-ncr53c7xx

Enables the NCR53C7,8xx SCSI controller devices sd%d, cd%d.

--enable-eatadma

Enables the EATA-DMA (DPT, NEC, AT&T, SNI, AST, Olivetti, Alphatronix) SCSI controller devices sd%d, cd%d.

--enable-eatapio

Enables the EATA-PIO (old DPT PM2001, PM2012A) SCSI controller devices sd%d, cd%d.

--enable-wd7000

Enables the WD 7000 SCSI controller devices sd%d, cd%d.

--enable-eata

Enables the EATA ISA/EISA/PCI (DPT and generic EATA/DMA-compliant boards) SCSI controller devices sd%d, cd%d.

--enable-am53c974
--enable-am79c974

Enables the AM53/79C974 SCSI controller devices sd%d, cd%d.

--enable-dtc3280
--enable-dtc3180

Enables the DTC3180/3280 SCSI controller devices sd%d, cd%d.

--enable-ncr53c8xx
--enable-dc390w
--enable-dc390u
--enable-dc390f

Enables the NCR53C8XX SCSI controller devices sd%d, cd%d.

--enable-dc390t
--enable-dc390

Enables the Tekram DC-390(T) SCSI controller devices sd%d, cd%d.

--enable-ppa

Enables the IOMEGA Parallel Port ZIP drive device sd%d.

--enable-qlogicfas

Enables the Qlogic FAS SCSI controller devices sd%d, cd%d.

--enable-qlogicisp

Enables the Qlogic ISP SCSI controller devices sd%d, cd%d.

--enable-gdth

Enables the GDT SCSI Disk Array controller devices sd%d, cd%d.

The following options enable drivers for various ethernet cards. NIC device names are usually eth%d, except for the pocket adaptors.

GNU Mach does only autodetect one ethernet card. To enable any further cards, the source code has to be edited.

--enable-ne2000
--enable-ne1000

Enables the NE2000/NE1000 ISA netword card devices eth%d.

--enable-3c503
--enable-el2

Enables the 3Com 503 (Etherlink II) netword card devices eth%d.

--enable-3c509
--enable-3c579
--enable-el3

Enables the 3Com 509/579 (Etherlink III) netword card devices eth%d.

--enable-wd80x3

Enables the WD80X3 netword card devices eth%d.

--enable-3c501
--enable-el1

Enables the 3COM 501 netword card devices eth%d.

--enable-ul

Enables the SMC Ultra netword card devices eth%d.

--enable-ul32

Enables the SMC Ultra 32 netword card devices eth%d.

--enable-hplanplus

Enables the HP PCLAN+ (27247B and 27252A) netword card devices eth%d.

--enable-hplan

Enables the HP PCLAN (27245 and other 27xxx series) netword card devices eth%d.

--enable-3c59x
--enable-3c90x
--enable-vortex

Enables the 3Com 590/900 series (592/595/597/900/905) "Vortex/Boomerang" netword card devices eth%d.

--enable-seeq8005

Enables the Seeq8005 netword card devices eth%d.

--enable-hp100
--enable-hpj2577
--enable-hpj2573
--enable-hp27248b
--enable-hp2585

Enables the HP 10/100VG PCLAN (ISA, EISA, PCI) netword card devices eth%d.

--enable-ac3200

Enables the Ansel Communications EISA 3200 netword card devices eth%d.

--enable-e2100

Enables the Cabletron E21xx netword card devices eth%d.

--enable-at1700

Enables the AT1700 (Fujitsu 86965) netword card devices eth%d.

--enable-eth16i
--enable-eth32

Enables the ICL EtherTeam 16i/32 netword card devices eth%d.

--enable-znet
--enable-znote

Enables the Zenith Z-Note netword card devices eth%d.

--enable-eexpress

Enables the EtherExpress 16 netword card devices eth%d.

--enable-eexpresspro

Enables the EtherExpressPro netword card devices eth%d.

--enable-eexpresspro100

Enables the Intel EtherExpressPro PCI 10+/100B/100+ netword card devices eth%d.

--enable-depca
--enable-de100
--enable-de101
--enable-de200
--enable-de201
--enable-de202
--enable-de210
--enable-de422

Enables the DEPCA, DE10x, DE200, DE201, DE202, DE210, DE422 netword card devices eth%d.

--enable-ewrk3
--enable-de203
--enable-de204
--enable-de205

Enables the EtherWORKS 3 (DE203, DE204, DE205) netword card devices eth%d.

--enable-de4x5
--enable-de425
--enable-de434
--enable-435
--enable-de450
--enable-500

Enables the DE425, DE434, DE435, DE450, DE500 netword card devices eth%d.

--enable-apricot

Enables the Apricot XEN-II on board ethernet netword card devices eth%d.

--enable-wavelan

Enables the AT&T WaveLAN & DEC RoamAbout DS netword card devices eth%d.

--enable-3c507
--enable-el16

Enables the 3Com 507 netword card devices eth%d.

--enable-3c505
--enable-elplus

Enables the 3Com 505 netword card devices eth%d.

--enable-de600

Enables the D-Link DE-600 netword card devices eth%d.

--enable-de620

Enables the D-Link DE-620 netword card devices eth%d.

--enable-skg16

Enables the Schneider & Koch G16 netword card devices eth%d.

--enable-ni52

Enables the NI5210 netword card devices eth%d.

--enable-ni65

Enables the NI6510 netword card devices eth%d.

--enable-atp

Enables the AT-LAN-TEC/RealTek pocket adaptor netword card devices atp%d.

--enable-lance
--enable-at1500
--enable-ne2100

Enables the AMD LANCE and PCnet (AT1500 and NE2100) netword card devices eth%d.

--enable-elcp
--enable-tulip

Enables the DECchip Tulip (dc21x4x) PCI netword card devices eth%d.

--enable-fmv18x

Enables the FMV-181/182/183/184 netword card devices eth%d.

--enable-3c515

Enables the 3Com 515 ISA Fast EtherLink netword card devices eth%d.

--enable-pcnet32

Enables the AMD PCI PCnet32 (PCI bus NE2100 cards) netword card devices eth%d.

--enable-ne2kpci

Enables the PCI NE2000 netword card devices eth%d.

--enable-yellowfin

Enables the Packet Engines Yellowfin Gigabit-NIC netword card devices eth%d.

--enable-rtl8139
--enable-rtl8129

Enables the RealTek 8129/8139 (not 8019/8029!) netword card devices eth%d.

--enable-epic
--enable-epic100

Enables the SMC 83c170/175 EPIC/100 (EtherPower II) netword card devices eth%d.

--enable-tlan

Enables the TI ThunderLAN netword card devices eth%d.

--enable-viarhine

Enables the VIA Rhine netword card devices eth%d.

2.4 Cross-Compilation

Another way to install the kernel is to use an existing operating system in order to compile the kernel binary. This is called cross-compiling, because it is done between two different platforms. If the pre-built kernels are not working for you, and you can’t ask someone to compile a custom kernel for your machine, this is your last chance to get a kernel that boots on your hardware.

Luckily, the kernel does have light dependencies. You don’t even need a cross compiler if your build machine has a compiler and is the same architecture as the system you want to run GNU Mach on.

You need a cross-mig, though.

10 The GNU Mach Reference Manual

3 Bootstrap

Bootstrapping^{1} is the procedure by which your machine loads the microkernel and transfers control to the operating system.

3.1 Bootloader

The bootloader is the first software that runs on your machine. Many hardware architectures have a very simple startup routine which reads a very simple bootloader from the beginning of the internal hard disk, then transfers control to it. Other architectures have startup routines which are able to understand more of the contents of the hard disk, and directly start a more advanced bootloader.

Currently, GRUB^{2} is the preferred GNU bootloader. GRUB provides advanced functionality, and is capable of loading several different kernels (such as Mach, Linux, DOS, and the *BSD family).

GNU Mach conforms to the Multiboot specification which defines an interface between the bootloader and the components that run very early at startup. GNU Mach can be started by any bootloader which supports the multiboot standard. After the bootloader loaded the kernel image to a designated address in the system memory, it jumps into the startup code of the kernel. This code initializes the kernel and detects the available hardware devices. Afterwards, the first system task is started.

3.2 Modules

Because the microkernel does not provide filesystem support and other features necessary to load the first system task from a storage medium, the first task is loaded by the bootloader as a module to a specified address. In the GNU system, this first program is the serverboot executable. GNU Mach inserts the host control port and the device master port into this task and appends the port numbers to the command line before executing it.

The serverboot program is responsible for loading and executing the rest of the Hurd servers. Rather than containing specific instructions for starting the Hurd, it follows general steps given in a user-supplied boot script.

XXX finish

1 The term bootstrapping refers to a Dutch legend about a boy who was able to fly by pulling himself up by his bootstraps. In computers, this term refers to any process where a simple system activates a more complicated system.

2 The GRand Unified Bootloader, available from http://www.uruk.org/grub/.

12 The GNU Mach Reference Manual

4 Inter Process Communication

This chapter describes the details of the Mach IPC system. Only the actual calls concerned with sending and receiving messages are discussed here. The details of the port system are described in the next chapter.

4.1 Major Concepts

The Mach kernel provides message-oriented, capability-based interprocess communication. The interprocess communication (IPC) primitives efficiently support many different styles of interaction, including remote procedure calls, object-oriented distributed programming, streaming of data, and sending very large amounts of data.

The IPC primitives operate on three abstractions: messages, ports, and port sets. User tasks access all other kernel services and abstractions via the IPC primitives.

The message primitives let tasks send and receive messages. Tasks send messages to ports. Messages sent to a port are delivered reliably (messages may not be lost) and are received in the order in which they were sent. Messages contain a fixed-size header and a variable amount of typed data following the header. The header describes the destination and size of the message.

The IPC implementation makes use of the VM system to efficiently transfer large amounts of data. The message body can contain the address of a region in the sender’s address space which should be transferred as part of the message. When a task receives a message containing an out-of-line region of data, the data appears in an unused portion of the receiver’s address space. This transmission of out-of-line data is optimized so that sender and receiver share the physical pages of data copy-on-write, and no actual data copy occurs unless the pages are written. Regions of memory up to the size of a full address space may be sent in this manner.

Ports hold a queue of messages. Tasks operate on a port to send and receive messages by exercising capabilities for the port. Multiple tasks can hold send capabilities, or rights, for a port. Tasks can also hold send-once rights, which grant the ability to send a single message. Only one task can hold the receive capability, or receive right, for a port. Port rights can be transferred between tasks via messages. The sender of a message can specify in the message body that the message contains a port right. If a message contains a receive right for a port, then the receive right is removed from the sender of the message and the right is transferred to the receiver of the message. While the receive right is in transit, tasks holding send rights can still send messages to the port, and they are queued until a task acquires the receive right and uses it to receive the messages.

Tasks can receive messages from ports and port sets. The port set abstraction allows a single thread to wait for a message from any of several ports. Tasks manipulate port sets with a capability, or port-set right, which is taken from the same space as the port capabilities. The port-set right may not be transferred in a message. A port set holds receive rights, and a receive operation on a port set blocks waiting for a message sent to any of the constituent ports. A port may not belong to more than one port set, and if a port is a member of a port set, the holder of the receive right can’t receive directly from the port.

Port rights are a secure, location-independent way of naming ports. The port queue is a protected data structure, only accessible via the kernel’s exported message primitives. Rights are also protected by the kernel; there is no way for a malicious user task to guess a port name and send a message to a port to which it shouldn’t have access. Port rights do not carry any location information. When a receive right for a port moves from task to task, and even between tasks on different machines, the send rights for the port remain unchanged and continue to function.

4.2 Messaging Interface

This section describes how messages are composed, sent and received within the Mach IPC system.

4.2.1 Mach Message Call

To use the mach_msg call, you can include the header files ‘mach/port.h’ and ‘mach/message.h’.

Function mach_msg_return_t mach msg (mach_msg_header_t *msg,
    mach_msg_option_t option, mach_msg_size_t send size,
    mach_msg_size_t rcv size, mach_port_t rcv name,
    mach_msg_timeout_t timeout, mach_port_t notify)

The mach_msg function is used to send and receive messages. Mach messages contain typed data, which can include port rights and references to large regions of memory. msg is the address of a buffer in the caller’s address space. Message buffers should be aligned on long-word boundaries. The message options option are bit values, combined with bitwise-or. One or both of MACH_SEND_MSG and MACH_RCV_MSG should be used. Other options act as modifiers. When sending a message, send size specifies the size of the message buffer. Otherwise zero should be supplied. When receiving a message, rcv size specifies the size of the message buffer. Otherwise zero should be supplied. When receiving a message, rcv name specifies the port or port set. Otherwise MACH_PORT_NULL should be supplied. When using the MACH_SEND_TIMEOUT and MACH_RCV_TIMEOUT options, timeout specifies the time in milliseconds to wait before giving up. Otherwise MACH_MSG_TIMEOUT_NONE should be supplied. When

using the MACH_SEND_NOTIFY, MACH_SEND_CANCEL, and MACH_RCV_NOTIFY options,

notify specifies the port used for the notification. Otherwise MACH_PORT_NULL should be supplied. If the option argument is MACH_SEND_MSG, it sends a message. The send size argument specifies the size of the message to send. The msgh_remote_port field of the message header specifies the destination of the message. If the option argument is MACH_RCV_MSG, it receives a message. The rcv size argument specifies the size of the message buffer that will receive the message; messages larger than rcv size are not received. The rcv name argument specifies the port or port set from which to receive.

If the option argument is MACH_SEND_MSG|MACH_RCV_MSG, then mach_msg does both

send and receive operations. If the send operation encounters an error (any return code other than MACH_MSG_SUCCESS), then the call returns immediately without attempting the receive operation. Semantically the combined call is equivalent to separate send and receive calls, but it saves a system call and enables other internal optimizations. If the option argument specifies neither MACH_SEND_MSG nor MACH_RCV_MSG, then mach_msg does nothing. Some options, like MACH_SEND_TIMEOUT and MACH_RCV_TIMEOUT, share a supporting argument. If these options are used together, they make independent use of the supporting argument’s value.

Data type natural t mach msg timeout t The type used by the timeout mechanism. The units are milliseconds. The value to be used when there is no timeout is MACH_MSG_TIMEOUT_NONE.

4.2.2 Message Format

A Mach message consists of a fixed size message header, a mach_msg_header_t, followed by zero or more data items. Data items are typed. Each item has a type descriptor followed by the actual data (or the address of the data, for out-of-line memory regions).

The following data types are related to Mach ports:

Data type mach port t

The mach_port_t data type is an unsigned integer type which represents a port name in the task’s port name space. In GNU Mach, this is an unsigned int.

The following data types are related to Mach messages:

Data type mach msg bits t

The mach_msg_bits_t data type is an unsigned int used to store various flags for a message.

Data type mach msg size t

The mach_msg_size_t data type is an unsigned int used to store the size of a message.

Data type mach msg id t

The mach_msg_id_t data type is an integer_t typically used to convey a function or operation id for the receiver.

Data type mach msg header t

This structure is the start of every message in the Mach IPC system. It has the following members:

mach_msg_bits_t msgh_bits

The msgh_bits field has the following bits defined, all other bits should be zero:

MACH_MSGH_BITS_REMOTE_MASK
MACH_MSGH_BITS_LOCAL_MASK

The remote and local bits encode mach_msg_type_name_t values that specify the port rights in the msgh_remote_port and msgh_local_port fields. The remote value must specify a send or send-once right for the destination of the message. If the local value doesn’t specify a send or send-once right for the message’s reply port, it must be zero and msgh local port

          must be MACH_PORT_NULL.
MACH_MSGH_BITS_COMPLEX

The complex bit must be specified if the message body contains port rights or out-of-line memory regions. If it is not specified, then the message body carries no port rights or memory, no matter what the type descriptors may seem to indicate.

MACH_MSGH_BITS_REMOTE and MACH_MSGH_BITS_LOCAL macros return the

appropriate mach_msg_type_name_t values, given a msgh_bits value. The MACH_MSGH_BITS macro constructs a value for msgh_bits, given two

mach_msg_type_name_t values.
mach_msg_size_t msgh_size

The msgh_size field in the header of a received message contains the message’s size. The message size, a byte quantity, includes the message header, type descriptors, and in-line data. For out-of-line memory regions, the message size includes the size of the in-line address, not the size of the actual memory region. There are no arbitrary limits on the size of a Mach message, the number of data items in a message, or the size of the data items.

mach_port_t msgh_remote_port

The msgh_remote_port field specifies the destination port of the message. The field must carry a legitimate send or send-once right for a port.

mach_port_t msgh_local_port

The msgh_local_port field specifies an auxiliary port right, which is conventionally used as a reply port by the recipient of the message. The field must carry a send right, a send-once right, MACH_PORT_NULL, or

MACH_PORT_DEAD.
mach_port_seqno_t msgh_seqno

The msgh_seqno field provides a sequence number for the message. It is only valid in received messages; its value in sent messages is overwritten.

mach_msg_id_t msgh_id

The mach_msg call doesn’t use the msgh_id field, but it conventionally conveys an operation or function id.

Macro mach_msg_bits_t MACH MSGH BITS

    (mach_msg_type_name_t remote, mach_msg_type_name_t local)

This macro composes two mach_msg_type_name_t values that specify the port rights

in the msgh_remote_port and msgh_local_port fields of a mach_msg call into an

appropriate mach_msg_bits_t value.

Macro mach_msg_type_name_t MACH MSGH BITS REMOTE

    (mach_msg_bits_t bits)

This macro extracts the mach_msg_type_name_t value for the remote port right in a

mach_msg_bits_t value.

Macro mach_msg_type_name_t MACH MSGH BITS LOCAL

    (mach_msg_bits_t bits)

This macro extracts the mach_msg_type_name_t value for the local port right in a

mach_msg_bits_t value.

Macro mach_msg_bits_t MACH MSGH BITS PORTS

    (mach_msg_bits_t bits)

This macro extracts the mach_msg_bits_t component consisting of the mach_msg_ type_name_t values for the remote and local port right in a mach_msg_bits_t value.

Macro mach_msg_bits_t MACH MSGH BITS OTHER

    (mach_msg_bits_t bits)

This macro extracts the mach_msg_bits_t component consisting of everything except the mach_msg_type_name_t values for the remote and local port right in a mach_msg_

bits_t value.

Each data item has a type descriptor, a mach_msg_type_t or a mach_msg_type_long_ t. The mach_msg_type_long_t type descriptor allows larger values for some fields. The msgtl_header field in the long descriptor is only used for its inline, longform, and deallocate bits.

Data type mach msg type name t

This is an unsigned int and can be used to hold the msgt_name component of the

mach_msg_type_t and mach_msg_type_long_t structure.

Data type mach msg type size t

This is an unsigned int and can be used to hold the msgt_size component of the

mach_msg_type_t and mach_msg_type_long_t structure.

Data type mach msg type number t

This is an natural_t and can be used to hold the msgt_number component of the

mach_msg_type_t and mach_msg_type_long_t structure.

Data type mach msg type t

This structure has the following members:

unsigned int msgt_name : 8

The msgt_name field specifies the data’s type. The following types are predefined:

MACH_MSG_TYPE_UNSTRUCTURED
MACH_MSG_TYPE_BIT
MACH_MSG_TYPE_BOOLEAN
MACH_MSG_TYPE_INTEGER_16
MACH_MSG_TYPE_INTEGER_32
MACH_MSG_TYPE_CHAR
MACH_MSG_TYPE_BYTE
MACH_MSG_TYPE_INTEGER_8
MACH_MSG_TYPE_REAL
MACH_MSG_TYPE_STRING
MACH_MSG_TYPE_STRING_C
MACH_MSG_TYPE_PORT_NAME

The following predefined types specify port rights, and receive special treatment. The next section discusses these types in detail. The type MACH_MSG_TYPE_PORT_NAME describes port right names, when no rights are being transferred, but just names. For this purpose, it should be used

in preference to MACH_MSG_TYPE_INTEGER_32.
MACH_MSG_TYPE_MOVE_RECEIVE
MACH_MSG_TYPE_MOVE_SEND
MACH_MSG_TYPE_MOVE_SEND_ONCE
MACH_MSG_TYPE_COPY_SEND
MACH_MSG_TYPE_MAKE_SEND
MACH_MSG_TYPE_MAKE_SEND_ONCE
msgt_size : 8

The msgt_size field specifies the size of each datum, in bits. For example,

the msgt size of MACH_MSG_TYPE_INTEGER_32 data is 32.
msgt_number : 12

The msgt_number field specifies how many data elements comprise the data item. Zero is a legitimate number. The total length specified by a type descriptor is (msgt_size * msgt_number), rounded up to an integral number of bytes. In-line data is then padded to an integral number of long-words. This ensures that type descriptors always start on long-word boundaries. It implies that message sizes are always an integral multiple of a long-word’s size.

msgt_inline : 1

The msgt_inline bit specifies, when FALSE, that the data actually resides in an out-of-line region. The address of the memory region (a vm_offset_ t or vm_address_t) follows the type descriptor in the message body.

The msgt_name, msgt_size, and msgt_number fields describe the memory region, not the address.

msgt_longform : 1

The msgt_longform bit specifies, when TRUE, that this type descriptor is a

mach_msg_type_long_t instead of a mach_msg_type_t. The msgt_name,

msgt_size, and msgt_number fields should be zero. Instead, mach_msg

uses the following msgtl_name, msgtl_size, and msgtl_number fields.
msgt_deallocate : 1

The msgt_deallocate bit is used with out-of-line regions. When TRUE, it specifies that the memory region should be deallocated from the sender’s address space (as if with vm_deallocate) when the message is sent.

msgt_unused : 1

The msgt_unused bit should be zero.

Macro boolean_t MACH MSG TYPE PORT ANY

(mach_msg_type_name_t type)

This macro returns TRUE if the given type name specifies a port type, otherwise it returns FALSE.

Macro boolean_t MACH MSG TYPE PORT ANY SEND

(mach_msg_type_name_t type)

This macro returns TRUE if the given type name specifies a port type with a send or send-once right, otherwise it returns FALSE.

Macro boolean_t MACH MSG TYPE PORT ANY RIGHT

(mach_msg_type_name_t type)

This macro returns TRUE if the given type name specifies a port right type which is moved, otherwise it returns FALSE.

Data type mach msg type long t

This structure has the following members:

mach_msg_type_t msgtl_header

Same meaning as msgt_header.

unsigned short msgtl_name

Same meaning as msgt_name.

unsigned short msgtl_size

Same meaning as msgt_size.

unsigned int msgtl_number

Same meaning as msgt_number.

4.2.3 Exchanging Port Rights

Each task has its own space of port rights. Port rights are named with positive integers. Except for the reserved values MACH_PORT_NULL (0)^{1} and MACH_PORT_DEAD (~0), this is a full 32-bit name space. When the kernel chooses a name for a new right, it is free to pick any unused name (one which denotes no right) in the space.

There are five basic kinds of rights: receive rights, send rights, send-once rights, port-set rights, and dead names. Dead names are not capabilities. They act as place-holders to prevent a name from being otherwise used.

A port is destroyed, or dies, when its receive right is deallocated. When a port dies, send and send-once rights for the port turn into dead names. Any messages queued at the port are destroyed, which deallocates the port rights and out-of-line memory in the messages.

Tasks may hold multiple user-references for send rights and dead names. When a task receives a send right which it already holds, the kernel increments the right’s user-reference count. When a task deallocates a send right, the kernel decrements its user-reference count, and the task only loses the send right when the count goes to zero.

Send-once rights always have a user-reference count of one, although a port can have multiple send-once rights, because each send-once right held by a task has a different name. In contrast, when a task holds send rights or a receive right for a port, the rights share a single name.

A message body can carry port rights; the msgt_name (msgtl_name) field in a type descriptor specifies the type of port right and how the port right is to be extracted from the caller. The values MACH_PORT_NULL and MACH_PORT_DEAD are always valid in place of a port right in a message body. In a sent message, the following msgt_name values denote port rights:

MACH_MSG_TYPE_MAKE_SEND

The message will carry a send right, but the caller must supply a receive right. The send right is created from the receive right, and the receive right’s makesend count is incremented.

MACH_MSG_TYPE_COPY_SEND

The message will carry a send right, and the caller should supply a send right. The user reference count for the supplied send right is not changed. The caller may also supply a dead name and the receiving task will get MACH_PORT_DEAD.

MACH_MSG_TYPE_MOVE_SEND

The message will carry a send right, and the caller should supply a send right. The user reference count for the supplied send right is decremented, and the right is destroyed if the count becomes zero. Unless a receive right remains, the name becomes available for recycling. The caller may also supply a dead name, which loses a user reference, and the receiving task will get MACH_PORT_DEAD.

MACH_MSG_TYPE_MAKE_SEND_ONCE

The message will carry a send-once right, but the caller must supply a receive right. The send-once right is created from the receive right.

MACH_MSG_TYPE_MOVE_SEND_ONCE

The message will carry a send-once right, and the caller should supply a sendonce right. The caller loses the supplied send-once right. The caller may also supply a dead name, which loses a user reference, and the receiving task will

get MACH_PORT_DEAD.
MACH_MSG_TYPE_MOVE_RECEIVE

The message will carry a receive right, and the caller should supply a receive right. The caller loses the supplied receive right, but retains any send rights with the same name.

If a message carries a send or send-once right, and the port dies while the message is in transit, then the receiving task will get MACH_PORT_DEAD instead of a right. The following msgt_name values in a received message indicate that it carries port rights:

MACH_MSG_TYPE_PORT_SEND

This name is an alias for MACH_MSG_TYPE_MOVE_SEND. The message carried a send right. If the receiving task already has send and/or receive rights for the port, then that name for the port will be reused. Otherwise, the new right will have a new name. If the task already has send rights, it gains a user reference for the right (unless this would cause the user-reference count to overflow). Otherwise, it acquires the send right, with a user-reference count of one.

MACH_MSG_TYPE_PORT_SEND_ONCE

This name is an alias for MACH_MSG_TYPE_MOVE_SEND_ONCE. The message carried a send-once right. The right will have a new name.

MACH_MSG_TYPE_PORT_RECEIVE

This name is an alias for MACH_MSG_TYPE_MOVE_RECEIVE. The message carried a receive right. If the receiving task already has send rights for the port, then that name for the port will be reused. Otherwise, the right will have a new name. The make-send count of the receive right is reset to zero, but the port retains other attributes like queued messages, extant send and send-once rights, and requests for port-destroyed and no-senders notifications.

When the kernel chooses a new name for a port right, it can choose any name, other than MACH_PORT_NULL and MACH_PORT_DEAD, which is not currently being used for a port right or dead name. It might choose a name which at some previous time denoted a port right, but is currently unused.

4.2.4 Memory

A message body can contain the address of a region in the sender’s address space which should be transferred as part of the message. The message carries a logical copy of the memory, but the kernel uses VM techniques to defer any actual page copies. Unless the sender or the receiver modifies the data, the physical pages remain shared.

An out-of-line transfer occurs when the data’s type descriptor specifies msgt_inline as FALSE. The address of the memory region (a vm_offset_t or vm_address_t) should follow the type descriptor in the message body. The type descriptor and the address contribute to the message’s size (send_size, msgh_size). The out-of-line data does not contribute to the message’s size.

The name, size, and number fields in the type descriptor describe the type and length of the out-of-line data, not the in-line address. Out-of-line memory frequently requires long type descriptors (mach_msg_type_long_t), because the msgt_number field is too small to describe a page of 4K bytes.

Out-of-line memory arrives somewhere in the receiver’s address space as new memory. It has the same inheritance and protection attributes as newly vm_allocate’d memory. The receiver has the responsibility of deallocating (with vm_deallocate) the memory when it is no longer needed. Security-conscious receivers should exercise caution when using outof-line memory from untrustworthy sources, because the memory may be backed by an unreliable memory manager.

Null out-of-line memory is legal. If the out-of-line region size is zero (for example, because msgtl_number is zero), then the region’s specified address is ignored. A received null out-of-line memory region always has a zero address.

Unaligned addresses and region sizes that are not page multiples are legal. A received message can also contain memory with unaligned addresses and funny sizes. In the general case, the first and last pages in the new memory region in the receiver do not contain only data from the sender, but are partly zero.^{2} The received address points to the start of the data in the first page. This possibility doesn’t complicate deallocation, because vm_ deallocate does the right thing, rounding the start address down and the end address up to deallocate all arrived pages.

Out-of-line memory has a deallocate option, controlled by the msgt_deallocate bit. If it is TRUE and the out-of-line memory region is not null, then the region is implicitly deallocated from the sender, as if by vm_deallocate. In particular, the start and end addresses are rounded so that every page overlapped by the memory region is deallocated. The use of msgt_deallocate effectively changes the memory copy into a memory movement. In a received message, msgt_deallocate is TRUE in type descriptors for out-of-line memory.

Out-of-line memory can carry port rights.

4.2.5 Message Send

The send operation queues a message to a port. The message carries a copy of the caller’s data. After the send, the caller can freely modify the message buffer or the out-of-line memory regions and the message contents will remain unchanged.

Message delivery is reliable and sequenced. Messages are not lost, and messages sent to a port, from a single thread, are received in the order in which they were sent.

If the destination port’s queue is full, then several things can happen. If the message is sent to a send-once right (msgh_remote_port carries a send-once right), then the kernel ignores the queue limit and delivers the message. Otherwise the caller blocks until there is room in the queue, unless the MACH_SEND_TIMEOUT or MACH_SEND_NOTIFY options are used. If a port has several blocked senders, then any of them may queue the next message when space in the queue becomes available, with the proviso that a blocked sender will not be indefinitely starved.

2 Sending out-of-line memory with a non-page-aligned address, or a size which is not a page multiple, works but with a caveat. The extra bytes in the first and last page of the received memory are not zeroed, so the receiver can peek at more data than the sender intended to transfer. This might be a security problem for the sender.

These options modify MACH_SEND_MSG. If MACH_SEND_MSG is not also specified, they are ignored.

MACH_SEND_TIMEOUT

The timeout argument should specify a maximum time (in milliseconds) for the call to block before giving up.^{3} If the message can’t be queued before the timeout interval elapses, then the call returns MACH_SEND_TIMED_OUT. A zero timeout is legitimate.

MACH_SEND_NOTIFY

The notify argument should specify a receive right for a notify port. If the send were to block, then instead the message is queued, MACH_SEND_WILL_NOTIFY is returned, and a msg-accepted notification is requested. If MACH_SEND_TIMEOUT is also specified, then MACH_SEND_NOTIFY doesn’t take effect until the timeout interval elapses. With MACH_SEND_NOTIFY, a task can forcibly queue to a send right one message at a time. A msg-accepted notification is sent to the the notify port when another message can be forcibly queued. If an attempt is made to use MACH_

SEND_NOTIFY before then, the call returns a MACH_SEND_NOTIFY_IN_PROGRESS

error. The msg-accepted notification carries the name of the send right. If the send right is deallocated before the msg-accepted notification is generated, then the msg-accepted notification carries the value MACH_PORT_NULL. If the destination port is destroyed before the notification is generated, then a send-once notification is generated instead.

MACH_SEND_INTERRUPT

If specified, the mach_msg call will return MACH_SEND_INTERRUPTED if a software interrupt aborts the call. Otherwise, the send operation will be retried.

MACH_SEND_CANCEL

The notify argument should specify a receive right for a notify port. If the send operation removes the destination port right from the caller, and the removed right had a dead-name request registered for it, and notify is the notify port for the dead-name request, then the dead-name request may be silently canceled (instead of resulting in a port-deleted notification). This option is typically used to cancel a dead-name request made with the MACH_RCV_NOTIFY option. It should only be used as an optimization.

The send operation can generate the following return codes. These return codes imply that the call did nothing:

MACH_SEND_MSG_TOO_SMALL

The specified send size was smaller than the minimum size for a message.

3 If MACH SEND TIMEOUT is used without MACH SEND INTERRUPT, then the timeout duration might not be accurate. When the call is interrupted and automatically retried, the original timeout is used. If interrupts occur frequently enough, the timeout interval might never expire.

MACH_SEND_NO_BUFFER

A resource shortage prevented the kernel from allocating a message buffer.

MACH_SEND_INVALID_DATA

The supplied message buffer was not readable.

MACH_SEND_INVALID_HEADER

The msgh_bits value was invalid.

MACH_SEND_INVALID_DEST

The msgh_remote_port value was invalid.

MACH_SEND_INVALID_REPLY

The msgh_local_port value was invalid.

MACH_SEND_INVALID_NOTIFY

When using MACH_SEND_CANCEL, the notify argument did not denote a valid receive right.

These return codes imply that some or all of the message was destroyed:

MACH_SEND_INVALID_MEMORY

The message body specified out-of-line data that was not readable.

MACH_SEND_INVALID_RIGHT

The message body specified a port right which the caller didn’t possess.

MACH_SEND_INVALID_TYPE

A type descriptor was invalid.

MACH_SEND_MSG_TOO_SMALL

The last data item in the message ran over the end of the message.

These return codes imply that the message was returned to the caller with a pseudoreceive operation:

MACH_SEND_TIMED_OUT

The timeout interval expired.

MACH_SEND_INTERRUPTED

A software interrupt occurred.

MACH_SEND_INVALID_NOTIFY

When using MACH_SEND_NOTIFY, the notify argument did not denote a valid receive right.

MACH_SEND_NO_NOTIFY

A resource shortage prevented the kernel from setting up a msg-accepted notification.

MACH_SEND_NOTIFY_IN_PROGRESS

A msg-accepted notification was already requested, and hasn’t yet been generated.

These return codes imply that the message was queued:

MACH_SEND_WILL_NOTIFY

The message was forcibly queued, and a msg-accepted notification was requested.

MACH_MSG_SUCCESS

The message was queued.

Some return codes, like MACH_SEND_TIMED_OUT, imply that the message was almost sent, but could not be queued. In these situations, the kernel tries to return the message contents to the caller with a pseudo-receive operation. This prevents the loss of port rights or memory which only exist in the message. For example, a receive right which was moved into the message, or out-of-line memory sent with the deallocate bit.

The pseudo-receive operation is very similar to a normal receive operation. The pseudoreceive handles the port rights in the message header as if they were in the message body. They are not reversed. After the pseudo-receive, the message is ready to be resent. If the message is not resent, note that out-of-line memory regions may have moved and some port rights may have changed names.

The pseudo-receive operation may encounter resource shortages. This is similar to a MACH_RCV_BODY_ERROR return code from a receive operation. When this happens, the normal send return codes are augmented with the MACH_MSG_IPC_SPACE, MACH_MSG_VM_SPACE, MACH_MSG_IPC_KERNEL, and MACH_MSG_VM_KERNEL bits to indicate the nature of the resource shortage.

The queueing of a message carrying receive rights may create a circular loop of receive rights and messages, which can never be received. For example, a message carrying a receive right can be sent to that receive right. This situation is not an error, but the kernel will garbage-collect such loops, destroying the messages and ports involved.

4.2.6 Message Receive

The receive operation dequeues a message from a port. The receiving task acquires the port rights and out-of-line memory regions carried in the message.

The rcv_name argument specifies a port or port set from which to receive. If a port is specified, the caller must possess the receive right for the port and the port must not be a member of a port set. If no message is present, then the call blocks, subject to the

MACH_RCV_TIMEOUT option.

If a port set is specified, the call will receive a message sent to any of the member ports. It is permissible for the port set to have no member ports, and ports may be added and removed while a receive from the port set is in progress. The received message can come from any of the member ports which have messages, with the proviso that a member port with messages will not be indefinitely starved. The msgh_local_port field in the received message header specifies from which port in the port set the message came.

The rcv_size argument specifies the size of the caller’s message buffer. The mach_ msg call will not receive a message larger than rcv_size. Messages that are too large are destroyed, unless the MACH_RCV_LARGE option is used.

The destination and reply ports are reversed in a received message header. The msgh_ local_port field names the destination port, from which the message was received, and the msgh_remote_port field names the reply port right. The bits in msgh_bits are also

reversed. The MACH_MSGH_BITS_LOCAL bits have the value MACH_MSG_TYPE_PORT_SEND if

the message was sent to a send right, and the value MACH_MSG_TYPE_PORT_SEND_ONCE if was sent to a send-once right. The MACH_MSGH_BITS_REMOTE bits describe the reply port right.

A received message can contain port rights and out-of-line memory. The msgh_local_ port field does not receive a port right; the act of receiving the message destroys the send or send-once right for the destination port. The msgh remote port field does name a received port right, the reply port right, and the message body can carry port rights and memory if MACH_MSGH_BITS_COMPLEX is present in msgh bits. Received port rights and memory should be consumed or deallocated in some fashion.

In almost all cases, msgh_local_port will specify the name of a receive right, either rcv_ name or if rcv_name is a port set, a member of rcv_name. If other threads are concurrently manipulating the receive right, the situation is more complicated. If the receive right is renamed during the call, then msgh_local_port specifies the right’s new name. If the caller loses the receive right after the message was dequeued from it, then mach_msg will proceed instead of returning MACH_RCV_PORT_DIED. If the receive right was destroyed, then msgh_local_port specifies MACH_PORT_DEAD. If the receive right still exists, but isn’t held

by the caller, then msgh_local_port specifies MACH_PORT_NULL.

Received messages are stamped with a sequence number, taken from the port from which the message was received. (Messages received from a port set are stamped with a sequence number from the appropriate member port.) Newly created ports start with a zero sequence number, and the sequence number is reset to zero whenever the port’s receive right moves between tasks. When a message is dequeued from the port, it is stamped with the port’s sequence number and the port’s sequence number is then incremented. The dequeue and increment operations are atomic, so that multiple threads receiving messages from a port can use the msgh_seqno field to reconstruct the original order of the messages.

These options modify MACH_RCV_MSG. If MACH_RCV_MSG is not also specified, they are ignored.

MACH_RCV_TIMEOUT

The timeout argument should specify a maximum time (in milliseconds) for the call to block before giving up.^{4} If no message arrives before the timeout interval elapses, then the call returns MACH_RCV_TIMED_OUT. A zero timeout is legitimate.

MACH_RCV_NOTIFY

The notify argument should specify a receive right for a notify port. If receiving the reply port creates a new port right in the caller, then the notify port is used to request a dead-name notification for the new port right.

MACH_RCV_INTERRUPT

If specified, the mach_msg call will return MACH_RCV_INTERRUPTED if a software interrupt aborts the call. Otherwise, the receive operation will be retried.

4 If MACH RCV TIMEOUT is used without MACH RCV INTERRUPT, then the timeout duration might not be accurate. When the call is interrupted and automatically retried, the original timeout is used. If interrupts occur frequently enough, the timeout interval might never expire.

MACH_RCV_LARGE

If the message is larger than rcv_size, then the message remains queued instead of being destroyed. The call returns MACH_RCV_TOO_LARGE and the actual size of the message is returned in the msgh_size field of the message header.

The receive operation can generate the following return codes. These return codes imply that the call did not dequeue a message:

MACH_RCV_INVALID_NAME

The specified rcv_name was invalid.

MACH_RCV_IN_SET

The specified port was a member of a port set.

MACH_RCV_TIMED_OUT

The timeout interval expired.

MACH_RCV_INTERRUPTED

A software interrupt occurred.

MACH_RCV_PORT_DIED

The caller lost the rights specified by rcv_name.

MACH_RCV_PORT_CHANGED

rcv_name specified a receive right which was moved into a port set during the call.

MACH_RCV_TOO_LARGE

When using MACH_RCV_LARGE, and the message was larger than rcv_size. The message is left queued, and its actual size is returned in the msgh_size field of the message buffer.

These return codes imply that a message was dequeued and destroyed:

MACH_RCV_HEADER_ERROR

A resource shortage prevented the reception of the port rights in the message header.

MACH_RCV_INVALID_NOTIFY

When using MACH_RCV_NOTIFY, the notify argument did not denote a valid receive right.

MACH_RCV_TOO_LARGE

When not using MACH_RCV_LARGE, a message larger than rcv_size was dequeued and destroyed.

In these situations, when a message is dequeued and then destroyed, the reply port and all port rights and memory in the message body are destroyed. However, the caller receives the message’s header, with all fields correct, including the destination port but excepting the reply port, which is MACH_PORT_NULL.

These return codes imply that a message was received:

MACH_RCV_BODY_ERROR

A resource shortage prevented the reception of a port right or out-of-line memory region in the message body. The message header, including the reply port, is correct. The kernel attempts to transfer all port rights and memory regions in the body, and only destroys those that can’t be transferred.

MACH_RCV_INVALID_DATA

The specified message buffer was not writable. The calling task did successfully receive the port rights and out-of-line memory regions in the message.

MACH_MSG_SUCCESS

A message was received.

Resource shortages can occur after a message is dequeued, while transferring port rights and out-of-line memory regions to the receiving task. The mach_msg call returns MACH_RCV_ HEADER_ERROR or MACH_RCV_BODY_ERROR in this situation. These return codes always carry extra bits (bitwise-ored) that indicate the nature of the resource shortage:

MACH_MSG_IPC_SPACE

There was no room in the task’s IPC name space for another port name.

MACH_MSG_VM_SPACE

There was no room in the task’s VM address space for an out-of-line memory region.

MACH_MSG_IPC_KERNEL

A kernel resource shortage prevented the reception of a port right.

MACH_MSG_VM_KERNEL

A kernel resource shortage prevented the reception of an out-of-line memory region.

If a resource shortage prevents the reception of a port right, the port right is destroyed and the caller sees the name MACH_PORT_NULL. If a resource shortage prevents the reception of an out-of-line memory region, the region is destroyed and the caller receives a zero address. In addition, the msgt_size (msgtl_size) field in the data’s type descriptor is changed to zero. If a resource shortage prevents the reception of out-of-line memory carrying port rights, then the port rights are always destroyed if the memory region can not be received. A task never receives port rights or memory regions that it isn’t told about.

4.2.7 Atomicity

The mach_msg call handles port rights in a message header atomically. Port rights and out-of-line memory in a message body do not enjoy this atomicity guarantee. The message body may be processed front-to-back, back-to-front, first out-of-line memory then port rights, in some random order, or even atomically.

For example, consider sending a message with the destination port specified as MACH_

MSG_TYPE_MOVE_SEND and the reply port specified as MACH_MSG_TYPE_COPY_SEND. The

same send right, with one user-reference, is supplied for both the msgh_remote_port and msgh_local_port fields. Because mach_msg processes the message header atomically, this

succeeds. If msgh_remote_port were processed before msgh_local_port, then mach_msg

would return MACH_SEND_INVALID_REPLY in this situation.

On the other hand, suppose the destination and reply port are both specified as MACH_ MSG_TYPE_MOVE_SEND, and again the same send right with one user-reference is supplied for both. Now the send operation fails, but because it processes the header atomically,

mach msg can return either MACH_SEND_INVALID_DEST or MACH_SEND_INVALID_REPLY.

For example, consider receiving a message at the same time another thread is deallocating the destination receive right. Suppose the reply port field carries a send right for the destination port. If the deallocation happens before the dequeuing, then the receiver gets MACH_RCV_PORT_DIED. If the deallocation happens after the receive, then the msgh_local_ port and the msgh_remote_port fields both specify the same right, which becomes a dead name when the receive right is deallocated. If the deallocation happens between the dequeue and the receive, then the msgh_local_port and msgh_remote_port fields both specify MACH_PORT_DEAD. Because the header is processed atomically, it is not possible for just one of the two fields to hold MACH_PORT_DEAD.

The MACH_RCV_NOTIFY option provides a more likely example. Suppose a message carrying a send-once right reply port is received with MACH_RCV_NOTIFY at the same time the reply port is destroyed. If the reply port is destroyed first, then msgh_remote_port specifies MACH_PORT_DEAD and the kernel does not generate a dead-name notification. If the reply port is destroyed after it is received, then msgh_remote_port specifies a dead name for which the kernel generates a dead-name notification. It is not possible to receive the reply port right and have it turn into a dead name before the dead-name notification is requested; as part of the message header the reply port is received atomically.

4.3 Port Manipulation Interface

This section describes the interface to create, destroy and manipulate ports and port sets.

4.3.1 Port Creation

Function kern_return_t mach port allocate (mach_port_t task,

         mach_port_right_t right, mach_port_t *name)

The mach_port_allocate function creates a new right in the specified task. The new right’s name is returned in name, which may be any name that wasn’t in use. The right argument takes the following values:

     MACH_PORT_RIGHT_RECEIVE

mach_port_allocate creates a port. The new port is not a member of any port set. It doesn’t have any extant send or send-once rights. Its make-send count is zero, its sequence number is zero, its queue limit is MACH_PORT_QLIMIT_DEFAULT, and it has no queued messages. name denotes the receive right for the new port. task does not hold send rights for the new port, only the receive right.

               mach_port_insert_right and mach_port_extract_right can be used

to convert the receive right into a combined send/receive right.

     MACH_PORT_RIGHT_PORT_SET

mach_port_allocate creates a port set. The new port set has no members.

MACH_PORT_RIGHT_DEAD_NAME

mach_port_allocate creates a dead name. The new dead name has one user reference.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_VALUE if right was invalid, KERN_NO_SPACE if there was no room in task’s IPC name space for another right and KERN_RESOURCE_SHORTAGE if the kernel ran out of memory. The mach_port_allocate call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function mach_port_t mach reply port ()

The mach_reply_port system call creates a reply port in the calling task. mach_reply_port creates a port, giving the calling task the receive right for the port. The call returns the name of the new receive right. This is very much like creating a receive right with the mach_port_allocate call, with two differences. First, mach_reply_port is a system call and not an RPC (which requires a reply port). Second, the port created by mach_reply_port may be optimized for use as a reply port. The function returns MACH_PORT_NULL if a resource shortage prevented the creation of the receive right.

Function kern_return_t mach port allocate name (mach_port_t task,

    mach_port_right_t right, mach_port_t name)

The function mach_port_allocate_name creates a new right in the specified task, with a specified name for the new right. name must not already be in use for some right, and it can’t be the reserved values MACH_PORT_NULL and MACH_PORT_DEAD. The right argument takes the following values:

MACH_PORT_RIGHT_RECEIVE

mach_port_allocate_name creates a port. The new port is not a member of any port set. It doesn’t have any extant send or send-once rights. Its make-send count is zero, its sequence number is zero, its queue limit is MACH_PORT_QLIMIT_DEFAULT, and it has no queued messages. name denotes the receive right for the new port. task does not hold send rights for the new port, only the receive right.

          mach_port_insert_right and mach_port_extract_right can be used

to convert the receive right into a combined send/receive right.

MACH_PORT_RIGHT_PORT_SET

mach_port_allocate_name creates a port set. The new port set has no members.

MACH_PORT_RIGHT_DEAD_NAME

mach_port_allocate_name creates a new dead name. The new dead name has one user reference.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_VALUE if right was invalid or name was MACH_PORT_NULL or MACH_PORT_DEAD, KERN_NAME_EXISTS if name was already in use for a port right and KERN_RESOURCE_SHORTAGE if the kernel ran out of memory.

The mach_port_allocate_name call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.2 Port Destruction

Function kern_return_t mach port deallocate (mach_port_t task,

    mach_port_t name)

The function mach_port_deallocate releases a user reference for a right in task’s IPC name space. It allows a task to release a user reference for a send or send-once right without failing if the port has died and the right is now actually a dead name.

If name denotes a dead name, send right, or send-once right, then the right loses one user reference. If it only had one user reference, then the right is destroyed.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right and KERN_INVALID_ RIGHT if name denoted an invalid right.

The mach_port_deallocate call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port destroy (mach_port_t task,

    mach_port_t name)

The function mach_port_destroy deallocates all rights denoted by a name. The name becomes immediately available for reuse.

For most purposes, mach_port_mod_refs and mach_port_deallocate are preferable.

If name denotes a port set, then all members of the port set are implicitly removed from the port set.

If name denotes a receive right that is a member of a port set, the receive right is implicitly removed from the port set. If there is a port-destroyed request registered for the port, then the receive right is not actually destroyed, but instead is sent in a port-destroyed notification to the backup port. If there is no registered portdestroyed request, remaining messages queued to the port are destroyed and extant send and send-once rights turn into dead names. If those send and send-once rights have dead-name requests registered, then dead-name notifications are generated for them.

If name denotes a send-once right, then the send-once right is used to produce a send-once notification for the port.

If name denotes a send-once, send, and/or receive right, and it has a dead-name request registered, then the registered send-once right is used to produce a portdeleted notification for the name. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right. The mach_port_destroy call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.3 Port Names

Function kern_return_t mach port names (mach_port_t task,

    mach_port_name_array_t *names, mach_msg_type_number_t *ncount,
    mach_port_type_array_t *types, mach_msg_type_number_t *tcount)

The function mach_port_names returns information about task’s port name space. For each name, it also returns what type of rights task holds. (The same information returned by mach_port_type.) names and types are arrays that are automatically allocated when the reply message is received. The user should vm_deallocate them when the data is no longer needed. mach_port_names will return in names the names of the ports, port sets, and dead names in the task’s port name space, in no particular order and in ncount the number of names returned. It will return in types the type of each corresponding name, which indicates what kind of rights the task holds with that name. tcount should be the same as ncount. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_RESOURCE_SHORTAGE if the kernel ran out of memory. The mach_port_names call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port type (mach_port_t task,
    mach_port_t name, mach_port_type_t *ptype)

The function mach_port_type returns information about task’s rights for a specific name in its port name space. The returned ptype is a bitmask indicating what rights task holds for the port, port set or dead name. The bitmask is composed of the following bits:

MACH_PORT_TYPE_SEND

The name denotes a send right.

MACH_PORT_TYPE_RECEIVE

The name denotes a receive right.

MACH_PORT_TYPE_SEND_ONCE

The name denotes a send-once right.

MACH_PORT_TYPE_PORT_SET

The name denotes a port set.

MACH_PORT_TYPE_DEAD_NAME

The name is a dead name.

MACH_PORT_TYPE_DNREQUEST

A dead-name request has been registered for the right.

MACH_PORT_TYPE_MAREQUEST

A msg-accepted request for the right is pending.

MACH_PORT_TYPE_COMPAT

The port right was created in the compatibility mode.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid and KERN_INVALID_NAME if name did not denote a right. The mach_port_type call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port rename (mach_port_t task,

    mach_port_t old name, mach_port_t new name)

The function mach_port_rename changes the name by which a port, port set, or dead name is known to task. old name is the original name and new name the new name for the port right. new name must not already be in use, and it can’t be the

distinguished values MACH_PORT_NULL and MACH_PORT_DEAD.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if old name did not denote a right, KERN_INVALID_

VALUE if new name was MACH_PORT_NULL or MACH_PORT_DEAD, KERN_NAME_EXISTS if

new_name already denoted a right and KERN_RESOURCE_SHORTAGE if the kernel ran out of memory. The mach_port_rename call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.4 Port Rights

Function kern_return_t mach port get refs (mach_port_t task,

    mach_port_t name, mach_port_right_t right, mach_port_urefs_t *refs)

The function mach_port_get_refs returns the number of user references a task has for a right. The right argument takes the following values:

  • MACH_PORT_RIGHT_SEND
  • MACH_PORT_RIGHT_RECEIVE

  • MACH_PORT_RIGHT_SEND_ONCE
  • MACH_PORT_RIGHT_PORT_SET
  • MACH_PORT_RIGHT_DEAD_NAME

If name denotes a right, but not the type of right specified, then zero is returned. Otherwise a positive number of user references is returned. Note that a name may simultaneously denote send and receive rights.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task

was invalid, KERN_INVALID_VALUE if right was invalid and KERN_INVALID_NAME if

name did not denote a right.

The mach_port_get_refs call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port mod refs (mach_port_t task,

    mach_port_t name, mach_port_right_t right, mach_port_delta_t delta)

The function mach_port_mod_refs requests that the number of user references a task has for a right be changed. This results in the right being destroyed, if the number of user references is changed to zero. The task holding the right is task, name should denote the specified right. right denotes the type of right being modified. delta is the signed change to the number of user references.

The right argument takes the following values:

  • MACH_PORT_RIGHT_SEND
  • MACH_PORT_RIGHT_RECEIVE
  • MACH_PORT_RIGHT_SEND_ONCE
  • MACH_PORT_RIGHT_PORT_SET
  • MACH_PORT_RIGHT_DEAD_NAME

The number of user references for the right is changed by the amount delta, subject to the following restrictions: port sets, receive rights, and send-once rights may only have one user reference. The resulting number of user references can’t be negative. If the resulting number of user references is zero, the effect is to deallocate the right. For dead names and send rights, there is an implementation-defined maximum number of user references.

If the call destroys the right, then the effect is as described for mach_port_destroy, with the exception that mach_port_destroy simultaneously destroys all the rights denoted by a name, while mach_port_mod_refs can only destroy one right. The name will be available for reuse if it only denoted the one right.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_VALUE if right was invalid or the user-reference count would become negative, KERN_INVALID_NAME if name did not denote a right, KERN_ INVALID_RIGHT if name denoted a right, but not the specified right and KERN_UREFS_ OVERFLOW if the user-reference count would overflow.

The mach_port_mod_refs call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.5 Ports and other Tasks

Function kern_return_t mach port insert right (mach_port_t task,

    mach_port_t name, mach_port_t right, mach_msg_type_name_t right type)

The function mach port insert right inserts into task the caller’s right for a port, using a specified name for the right in the target task.

The specified name can’t be one of the reserved values MACH_PORT_NULL or MACH_

PORT_DEAD. The right can’t be MACH_PORT_NULL or MACH_PORT_DEAD.

The argument right type specifies a right to be inserted and how that right should be extracted from the caller. It should be a value appropriate for msgt name; see

mach_msg.
If right type is MACH_MSG_TYPE_MAKE_SEND, MACH_MSG_TYPE_MOVE_SEND, or MACH_

MSG_TYPE_COPY_SEND, then a send right is inserted. If the target already holds send or receive rights for the port, then name should denote those rights in the target. Otherwise, name should be unused in the target. If the target already has send rights, then those send rights gain an additional user reference. Otherwise, the target gains a send right, with a user reference count of one.

If right type is MACH_MSG_TYPE_MAKE_SEND_ONCE or MACH_MSG_TYPE_MOVE_SEND_

ONCE, then a send-once right is inserted. The name should be unused in the target. The target gains a send-once right.

If right type is MACH_MSG_TYPE_MOVE_RECEIVE, then a receive right is inserted. If the target already holds send rights for the port, then name should denote those rights in the target. Otherwise, name should be unused in the target. The receive right is moved into the target task.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_VALUE if right was not a port right or name was

MACH_PORT_NULL or MACH_PORT_DEAD, KERN_NAME_EXISTS if name already denoted
a right, KERN_INVALID_CAPABILITY if right was MACH_PORT_NULL or MACH_PORT_

DEAD KERN_RIGHT_EXISTS if task already had rights for the port, with a different name, KERN_UREFS_OVERFLOW if the user-reference count would overflow and KERN_ RESOURCE_SHORTAGE if the kernel ran out of memory.

The mach_port_insert_right call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port extract right (mach_port_t task,

    mach_port_t name, mach_msg_type_name_t desired type,
    mach_port_t *right, mach_msg_type_name_t *acquired type)

The function mach port extract right extracts a port right from the target task and returns it to the caller as if the task sent the right voluntarily, using desired type as the value of msgt name. See mach_msg. The returned value of acquired type will be MACH_MSG_TYPE_PORT_SEND if a send right is extracted, MACH_MSG_TYPE_PORT_RECEIVE if a receive right is extracted, and MACH_MSG_TYPE_PORT_SEND_ONCE if a send-once right is extracted. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right, KERN_INVALID_RIGHT if name denoted a right, but an invalid one, KERN_INVALID_VALUE if desired type was invalid. The mach_port_extract_right call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.6 Receive Rights

Data type mach port seqno t

The mach_port_seqno_t data type is an unsigned int which contains the sequence number of a port.

Data type mach port mscount t

The mach_port_mscount_t data type is an unsigned int which contains the makesend count for a port.

Data type mach port msgcount t

The mach_port_msgcount_t data type is an unsigned int which contains a number of messages.

Data type mach port rights t

The mach_port_rights_t data type is an unsigned int which contains a number of rights for a port.

Data type mach port status t

This structure contains some status information about a port, which can be queried with mach_port_get_receive_status. It has the following members:

mach_port_t mps_pset

The containing port set.

mach_port_seqno_t mps_seqno

The sequence number.

mach_port_mscount_t mps_mscount

The make-send count.

mach_port_msgcount_t mps_qlimit

The maximum number of messages in the queue.

mach_port_msgcount_t mps_msgcount

The number of messages in the queue.

mach_port_rights_t mps_sorights

The number of send-once rights that exist.

boolean_t mps_srights

TRUE when send rights exist.

boolean_t mps_pdrequest

TRUE if port-deleted notification is requested.

boolean_t mps_nsrequest

TRUE if no-senders notification is requested.

Function kern_return_t mach port get receive status

    (mach_port_t task, mach_port_t name, mach_port_status_t *status)

The function mach_port_get_receive_status returns the current status of the specified receive right. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right and KERN_INVALID_ RIGHT if name denoted a right, but not a receive right. The mach_port_get_receive_status call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port set mscount (mach_port_t task,

    mach_port_t name, mach_port_mscount_t mscount)

The function mach_port_set_mscount changes the make-send count of task’s receive right named name to mscount. All values for mscount are valid. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right and KERN_INVALID_ RIGHT if name denoted a right, but not a receive right. The mach_port_set_mscount call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port set qlimit (mach_port_t task,

    mach_port_t name, mach_port_msgcount_t qlimit)

The function mach_port_set_qlimit changes the queue limit task’s receive right named name to qlimit. Valid values for qlimit are between zero and MACH_PORT_ QLIMIT_MAX, inclusive.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right, KERN_INVALID_ RIGHT if name denoted a right, but not a receive right and KERN_INVALID_VALUE if qlimit was invalid. The mach_port_set_qlimit call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port set seqno (mach_port_t task,

    mach_port_t name, mach_port_seqno_t seqno)

The function mach_port_set_seqno changes the sequence number task’s receive right named name to seqno. All sequence number values are valid. The next message received from the port will be stamped with the specified sequence number. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right and KERN_INVALID_ RIGHT if name denoted a right, but not a receive right. The mach_port_set_seqno call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.7 Port Sets

Function kern_return_t mach port get set status (mach_port_t task,

    mach_port_t name, mach_port_array_t *members,
    mach_msg_type_number_t *count)

The function mach_port_get_set_status returns the members of a port set. members is an array that is automatically allocated when the reply message is received. The user should vm_deallocate it when the data is no longer needed. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if name did not denote a right, KERN_INVALID_ RIGHT if name denoted a right, but not a receive right and KERN_RESOURCE_SHORTAGE if the kernel ran out of memory. The mach_port_get_set_status call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t mach port move member (mach_port_t task,

    mach_port_t member, mach_port_t after)

The function mach port move member moves the receive right member into the port set after. If the receive right is already a member of another port set, it is removed from that set first (the whole operation is atomic). If the port set is MACH_PORT_NULL, then the receive right is not put into a port set, but removed from its current port set. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_NAME if member or after did not denote a right, KERN_ INVALID_RIGHT if member denoted a right, but not a receive right or after denoted a right, but not a port set, and KERN_NOT_IN_SET if after was MACH_PORT_NULL, but member wasn’t currently in a port set. The mach_port_move_member call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

4.3.8 Request Notifications

Function kern_return_t mach port request notification

    (mach_port_t task, mach_port_t name, mach_msg_id_t variant,
    mach_port_mscount_t sync, mach_port_t notify,
    mach_msg_type_name_t notify type, mach_port_t *previous)

The function mach_port_request_notification registers a request for a notification and supplies the send-once right notify to which the notification will be sent. The notify type denotes the IPC type for the send-once right, which can be MACH_MSG_

TYPE_MAKE_SEND_ONCE or MACH_MSG_TYPE_MOVE_SEND_ONCE. It is an atomic swap,

returning the previously registered send-once right (or MACH_PORT_NULL for none) in previous. A previous notification request may be cancelled by providing MACH_PORT_ NULL for notify. The variant argument takes the following values:

MACH_NOTIFY_PORT_DESTROYED

sync must be zero. The name must specify a receive right, and the call requests a port-destroyed notification for the receive right. If the receive right were to have been destroyed, say by mach_port_destroy, then instead the receive right will be sent in a port-destroyed notification to the registered send-once right.

MACH_NOTIFY_DEAD_NAME

The call requests a dead-name notification. name specifies send, receive, or send-once rights for a port. If the port is destroyed (and the right remains, becoming a dead name), then a dead-name notification which carries the name of the right will be sent to the registered send-once right. If notify is not null and sync is non-zero, the name may specify a dead name, and a dead-name notification is immediately generated. Whenever a dead-name notification is generated, the user reference count of the dead name is incremented. For example, a send right with two user refs has a registered dead-name request. If the port is destroyed, the send right turns into a dead name with three user refs (instead of two), and a dead-name notification is generated.

If the name is made available for reuse, perhaps because of mach_port_ destroy or mach_port_mod_refs, or the name denotes a send-once right which has a message sent to it, then the registered send-once right is used to generate a port-deleted notification.

MACH_NOTIFY_NO_SENDERS

The call requests a no-senders notification. name must specify a receive right. If notify is not null, and the receive right’s make-send count is greater than or equal to the sync value, and it has no extant send rights, than an immediate no-senders notification is generated. Otherwise the notification is generated when the receive right next loses its last extant send right. In either case, any previously registered send-once right is returned. The no-senders notification carries the value the port’s make-send count had when it was generated. The make-send count is incremented whenever MACH_MSG_TYPE_MAKE_SEND is used to create a new send right from the receive right. The make-send count is reset to zero when the receive right is carried in a message.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_TASK if task was invalid, KERN_INVALID_VALUE if variant was invalid, KERN_INVALID_NAME if name did not denote a right, KERN_INVALID_RIGHT if name denoted an invalid right and

KERN_INVALID_CAPABILITY if notify was invalid.
When using MACH_NOTIFY_PORT_DESTROYED, the function returns KERN_INVALID_

VALUE if sync wasn’t zero.

When using MACH_NOTIFY_DEAD_NAME, the function returns KERN_RESOURCE_SHORTAGE

if the kernel ran out of memory, KERN_INVALID_ARGUMENT if name denotes a dead name, but sync is zero or notify is MACH_PORT_NULL, and KERN_UREFS_OVERFLOW if name denotes a dead name, but generating an immediate dead-name notification would overflow the name’s user-reference count. The mach_port_request_notification call is actually an RPC to task, normally a send right for a task port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

5 Virtual Memory Interface

5.1 Memory Allocation

Function kern_return_t vm allocate (mach_port_t target task,

    vm_address_t *address, vm_size_t size, boolean_t anywhere)

The function vm_allocate allocates a region of virtual memory, placing it in the specified task’s address space. The starting address is address. If the anywhere option is false, an attempt is made to allocate virtual memory starting at this virtual address. If this address is not at the beginning of a virtual page, it will be rounded down to one. If there is not enough space at this address, no memory will be allocated. If the anywhere option is true, the input value of this address will be ignored, and the space will be allocated wherever it is available. In either case, the address at which memory was actually allocated will be returned in address. size is the number of bytes to allocate (rounded by the system in a machine dependent way to an integral number of virtual pages). If anywhere is true, the kernel should find and allocate any region of the specified size, and return the address of the resulting region in address address, rounded to a virtual page boundary if there is sufficient space. The physical memory is not actually allocated until the new virtual memory is referenced. By default, the kernel rounds all addresses down to the nearest page boundary and all memory sizes up to the nearest page size. The global variable vm_page_size contains the page size. mach_task_self returns the value of the current task port which should be used as the target task argument in order to allocate memory in the caller’s address space. For languages other than C, these values can be obtained by the calls vm_statistics and mach_task_self. Initially, the pages of allocated memory will be protected to allow all forms of access, and will be inherited in child tasks as a copy. Subsequent calls to vm_protect and vm_inherit may be used to change these properties. The allocated region is always zero-filled. The function returns KERN_SUCCESS if the memory was successfully allocated, KERN_ INVALID_ADDRESS if an illegal address was specified and KERN_NO_SPACE if there was not enough space left to satisfy the request.

5.2 Memory Deallocation

Function kern_return_t vm deallocate (mach_port_t target task,

    vm_address_t address, vm_size_t size)

vm_deallocate relinquishes access to a region of a task’s address space, causing further access to that memory to fail. This address range will be available for reallocation. address is the starting address, which will be rounded down to a page boundary. size is the number of bytes to deallocate, which will be rounded up to give a page boundary. Note, that because of the rounding to virtual page boundaries, more than size bytes may be deallocated. Use vm_page_size or vm_statistics to find out the current virtual page size.

This call may be used to deallocte memory that was passed to a task in a message (via out of line data). In that case, the rounding should cause no trouble, since the region of memory was allocated as a set of pages.

The vm_deallocate call affects only the task specified by the target task. Other tasks which may have access to this memory may continue to reference it.

The function returns KERN_SUCCESS if the memory was successfully deallocated and KERN_INVALID_ADDRESS if an illegal or non-allocated address was specified.

5.3 Data Transfer

Function kern_return_t vm read (mach_port_t target task,
    vm_address_t address, vm_size_t size, vm_offset_t *data,
    mach_msg_type_number_t *data count)

The function vm_read allows one task’s virtual memory to be read by another task. The target task is the task whose memory is to be read. address is the first address to be read and must be on a page boundary. size is the number of bytes of data to be read and must be an integral number of pages. data is the array of data copied from the given task, and data count is the size of the data array in bytes (will be an integral number of pages).

Note that the data array is returned in a newly allocated region; the task reading the data should vm_deallocate this region when it is done with the data.

The function returns KERN_SUCCESS if the memory was successfully read, KERN_ INVALID_ADDRESS if an illegal or non-allocated address was specified or there was not size bytes of data following the address, KERN_INVALID_ARGUMENT if the address does not start on a page boundary or the size is not an integral number of pages, KERN_ PROTECTION_FAILURE if the address region in the target task is protected against reading and KERN_NO_SPACE if there was not enough room in the callers virtual memory to allocate space for the data to be returned.

Function kern_return_t vm write (mach_port_t target task,

    vm_address_t address, vm_offset_t data,
    mach_msg_type_number_t data count)

The function vm_write allows a task to write to the vrtual memory of target task. address is the starting address in task to be affected. data is an array of bytes to be written, and data count the size of the data array.

The current implementation requires that address, data and data count all be pagealigned. Otherwise, KERN_INVALID_ARGUMENT is returned.

The function returns KERN_SUCCESS if the memory was successfully written, KERN_ INVALID_ADDRESS if an illegal or non-allocated address was specified or there was not data count bytes of allocated memory starting at address and KERN_PROTECTION_ FAILURE if the address region in the target task is protected against writing.

Function kern_return_t vm copy (mach_port_t target task,
    vm_address_t source address, vm_size_t count, vm_offset_t dest address)

The function vm_copy causes the source memory range to be copied to the destination address. The source and destination memory ranges may overlap. The destination address range must already be allocated and writable; the source range must be readable. vm_copy is equivalent to vm_read followed by vm_write. The current implementation requires that address, data and data count all be pagealigned. Otherwise, KERN_INVALID_ARGUMENT is returned. The function returns KERN_SUCCESS if the memory was successfully written, KERN_ INVALID_ADDRESS if an illegal or non-allocated address was specified or there was insufficient memory allocated at one of the addresses and KERN_PROTECTION_FAILURE if the destination region was not writable or the source region was not readable.

5.4 Memory Attributes

Function kern_return_t vm region (mach_port_t target task,

    vm_address_t *address, vm_size_t *size, vm_prot_t *protection,
    vm_prot_t *max protection, vm_inherit_t *inheritance, boolean_t *shared,
    mach_port_t *object name, vm_offset_t *offset)

The function vm_region returns a description of the specified region of target task’s virtual address space. vm_region begins at address and looks forward through memory until it comes to an allocated region. If address is within a region, then that region is used. Various bits of information about the region are returned. If address was not within a region, then address is set to the start of the first region which follows the incoming value. In this way an entire address space can be scanned. The size returned is the size of the located region in bytes. protection is the current protection of the region, max protection is the maximum allowable protection for this region. inheritance is the inheritance attribute for this region. shared tells if the region is shared or not. The port object name identifies the memory object associated with this region, and offset is the offset into the pager object that this region begins at. The function returns KERN_SUCCESS if the memory region was successfully located and the information returned and KERN_NO_SPACE if there is no region at or above address in the specified task.

Function kern_return_t vm protect (mach_port_t target task,

    vm_address_t address, vm_size_t size, boolean_t set maximum,

vm_prot_t new protection) The function vm_protect sets the virtual memory access privileges for a range of allocated addresses in target task’s virtual address space. The protection argument describes a combination of read, write, and execute accesses that should be permitted. address is the starting address, which will be rounded down to a page boundary. size is the size in bytes of the region for which protection is to change, and will be rounded up to give a page boundary. If set maximum is set, make the protection change apply to the maximum protection associated with this address range; otherwise, the current protection on this range is changed. If the maximum protection is reduced below the current protection, both will be changed to reflect the new maximum. new protection is the new protection value for this region; a set of: VM_PROT_READ, VM_PROT_WRITE,

VM_PROT_EXECUTE.

The enforcement of virtual memory protection is machine-dependent. Nominally read access requires VM_PROT_READ permission, write access requires VM_PROT_WRITE permission, and execute access requires VM_PROT_EXECUTE permission. However, some combinations of access rights may not be supported. In particular, the kernel interface allows write access to require VM_PROT_READ and VM_PROT_WRITE permission and execute access to require VM_PROT_READ permission. The function returns KERN_SUCCESS if the memory was successfully protected, KERN_ INVALID_ADDRESS if an illegal or non-allocated address was specified and KERN_ PROTECTION_FAILURE if an attempt was made to increase the current or maximum protection beyond the existing maximum protection value.

Function kern_return_t vm inherit (mach_port_t target task,

    vm_address_t address, vm_size_t size, vm_inherit_t new inheritance)

The function vm_inherit specifies how a region of target task’s address space is to be passed to child tasks at the time of task creation. Inheritance is an attribute of virtual pages, so address to start from will be rounded down to a page boundary and size, the size in bytes of the region for wihch inheritance is to change, will be rounded up to give a page boundary. How this memory is to be inherited in child tasks is specified by new inheritance. Inheritance is specified by using one of these following three values:

VM_INHERIT_SHARE

Child tasks will share this memory with this task.

VM_INHERIT_COPY

Child tasks will receive a copy of this region.

VM_INHERIT_NONE

This region will be absent from child tasks.

Setting vm_inherit to VM_INHERIT_SHARE and forking a child task is the only way two Mach tasks can share physical memory. Remember that all the theads of a given task share all the same memory. The function returns KERN_SUCCESS if the memory inheritance was successfully set and KERN_INVALID_ADDRESS if an illegal or non-allocated address was specified.

Function kern_return_t vm wire (mach_port_t host priv,
    mach_port_t target task, vm_address_t address, vm_size_t size,
    vm_prot_t access)

The function vm_wire allows privileged applications to control memory pageability. host priv is the privileged host port for the host on which target task resides. address is the starting address, which will be rounded down to a page boundary. size is the size in bytes of the region for which protection is to change, and will be rounded up to give a page boundary. access specifies the types of accesses that must not cause page faults.

The semantics of a successful vm_wire operation are that memory in the specified range will not cause page faults for any accesses included in access. Data memory can be made non-pageable (wired) with a access argument of VM_PROT_READ | VM_ PROT_WRITE. A special case is that VM_PROT_NONE makes the memory pageable.

The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_HOST if host priv was not the privileged host port, KERN_INVALID_TASK if task was not a valid task, KERN_INVALID_VALUE if access specified an invalid access mode, KERN_ FAILURE if some memory in the specified range is not present or has an inappropriate protection value, and KERN_INVALID_ARGUMENT if unwiring (access is VM_PROT_NONE) and the memory is not already wired.

The vm_wire call is actually an RPC to host priv, normally a send right for a privileged host port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

Function kern_return_t vm machine attribute (mach_port_t task,

    vm_address_t address, vm_size_t size, vm_prot_t access,
    vm_machine_attribute_t attribute, vm_machine_attribute_val_t value)

The function vm_machine_attribute specifies machine-specific attributes for a VM mapping, such as cachability, migrability, replicability. This is used on machines that allow the user control over the cache (this is the case for MIPS architectures) or placement of memory pages as in NUMA architectures (Non-Uniform Memory Access time) such as the IBM ACE multiprocessor.

Machine-specific attributes can be consider additions to the machine-independent ones such as protection and inheritance, but they are not guaranteed to be supported by any given machine. Moreover, implementations of Mach on new architectures might find the need for new attribute types and or values besides the ones defined in the initial implementation.

The types currently defined are

MATTR_CACHE

Controls caching of memory pages

MATTR_MIGRATE

Controls migrability of memory pages

MATTR_REPLICATE

Controls replication of memory pages

Corresponding values, and meaning of a specific call to vm_machine_attribute

MATTR_VAL_ON

Enables the attribute. Being enabled is the default value for any applicable attribute.

MATTR_VAL_OFF

Disables the attribute, making memory non-cached, or non-migratable, or non-replicatable.

MATTR_VAL_GET

Returns the current value of the attribute for the memory segment. If the attribute does not apply uniformly to the given range the value returned applies to the initial portion of the segment only.

MATTR_VAL_CACHE_FLUSH

Flush the memory pages from the Cache. The size value in this case might be meaningful even if not a multiple of the page size, depending on the implementation.

MATTR_VAL_ICACHE_FLUSH

Same as above, applied to the Instruction Cache alone.

MATTR_VAL_DCACHE_FLUSH

Same as above, applied to the Data Cache alone.

The function returns KERN_SUCCESS if call succeeded, and KERN_INVALID_ARGUMENT if task is not a task, or address and size do not define a valid address range in task, or attribute is not a valid attribute type, or it is not implemented, or value is not a permissible value for attribute.

5.5 Mapping Memory Objects

Function kern_return_t vm map (mach_port_t target task,
    vm_address_t *address, vm_size_t size, vm_address_t mask,
    boolean_t anywhere, mach_port_t memory object, vm_offset_t offset,
    boolean_t copy, vm_prot_t cur protection, vm_prot_t max protection,
    vm_inherit_t inheritance)

The function vm_map maps a region of virtual memory at the specified address, for which data is to be supplied by the given memory object, starting at the given offset within that object. In addition to the arguments used in vm_allocate, the vm_map call allows the specification of an address alignment parameter, and of the initial protection and inheritance values. If the memory object in question is not currently in use, the kernel will perform a memory_object_init call at this time. If the copy parameter is asserted, the specified region of the memory object will be copied to this address space; changes made to this object by other tasks will not be visible in this mapping, and changes made in this mapping will not be visible to others (or returned to the memory object). The vm_map call returns once the mapping is established. Completion of the call does not require any action on the part of the memory manager. Warning: Only memory objects that are provided by bona fide memory managers should be used in the vm_map call. A memory manager must implement the memory object interface described elsewhere in this manual. If other ports are used, a thread that accesses the mapped virtual memory may become permanently hung or may receive a memory exception. target task is the task to be affected. The starting address is address. If the anywhere option is used, this address is ignored. The address actually allocated will be returned in address. size is the number of bytes to allocate (rounded by the system in a machine dependent way). The alignment restriction is specified by mask. Bits asserted in this mask must not be asserted in the address returned. If anywhere is set, the kernel should find and allocate any region of the specified size, and return the address of the resulting region in address. memory object is the port that represents the memory object: used by user tasks in vm_map; used by the make requests for data or other management actions. If this port is MEMORY_OBJECT_NULL, then zero-filled memory is allocated instead. Within a memory object, offset specifes an offset in bytes. This must be page aligned. If copy is set, the range of the memory object should be copied to the target task, rather than mapped read-write. The function returns KERN_SUCCESS if the object is mapped, KERN_NO_SPACE if no unused region of the task’s virtual address space that meets the address, size, and alignment criteria could be found, and KERN_INVALID_ARGUMENT if an illegal argument was provided.

5.6 Memory Statistics

Data type vm statistics data t

This structure is returned in vm stats by the vm_statistics function and provides virtual memory statistics for the system. It has the following members:

long pagesize

The page size in bytes.

long free_count

The number of free pages.

long active_count

The umber of active pages.

long inactive_count

The number of inactive pages.

long wire_count

The number of pages wired down.

long zero_fill_count

The number of zero filled pages.

long reactivations

The number of reactivated pages.

long pageins

The number of pageins.

long pageouts

The number of pageouts.

long faults

The number of faults.

long cow_faults

The number of copy-on-writes.

long lookups

The number of object cache lookups.

long hits The number of object cache hits.

Function kern_return_t vm statistics (mach_port_t target task,

    vm_statistics_data_t *vm stats)

The function vm_statistics returns the statistics about the kernel’s use of virtual memory since the kernel was booted. pagesize can also be found as a global variable vm_page_size which is set at task initialization and remains constant for the life of the task.

6 External Memory Management

6.1 Memory Object Server

Function boolean_t memory object server (msg_header_t *in msg,

    msg_header_t *out msg)

Function boolean_t memory object default server

    (msg_header_t *in msg, msg_header_t *out msg)

Function boolean_t seqnos memory object server

    (msg_header_t *in msg, msg_header_t *out msg)

Function boolean_t seqnos memory object default server

    (msg_header_t *in msg, msg_header_t *out msg)

A memory manager is a server task that responds to specific messages from the kernel in order to handle memory management functions for the kernel.

In order to isolate the memory manager from the specifics of message formatting, the remote procedure call generator produces a procedure, memory_object_server, to handle a received message. This function does all necessary argument handling, and actually calls one of the following functions: memory_object_init,

memory_object_data_return, memory_object_data_request, memory_object_
data_unlock, memory_object_lock_completed, memory_object_copy, memory_

object_terminate. The default memory manager may get two additional requests

from the kernel: memory_object_create and memory_object_data_initialize.

The remote procedure call generator produces a procedure memory_object_default_ server to handle those functions specific to the default memory manager.

The seqnos_memory_object_server and seqnos_memory_object_default_server
differ from memory_object_server and memory_object_default_server in that

they supply message sequence numbers to the server interfaces. They call the seqnos_ memory_object_* functions, which complement the memory_object_* set of functions.

The return value from the memory_object_server function indicates that the message was appropriate to the memory management interface (returning TRUE), or that it could not handle this message (returning FALSE).

The in msg argument is the message that has been received from the kernel. The out msg is a reply message, but this is not used for this server.

The function returns TRUE to indicate that the message in question was applicable to this interface, and that the appropriate routine was called to interpret the message. It returns FALSE to indicate that the message did not apply to this interface, and that no other action was taken.

6.2 Memory Object Creation

Function kern_return_t memory object init (mach_port_t memory object,

mach_port_t memory control, mach_port_t memory object name, vm_size_t memory object page size)

Function kern_return_t seqnos memory object init

    (mach_port_t memory object, mach_port_seqno_t seqno,

mach_port_t memory control, mach_port_t memory object name, vm_size_t memory object page size)

The function memory_object_init serves as a notification that the kernel has been asked to map the given memory object into a task’s virtual address space. Additionally, it provides a port on which the memory manager may issue cache management requests, and a port which the kernel will use to name this data region. In the event that different each will perform a memory_object_init call with new request and name ports. The virtual page size that is used by the calling kernel is included for planning purposes.

When the memory manager is prepared to accept requests for data for this object, it must call memory_object_ready with the attribute. Otherwise the kernel will not process requests on this object. To reject all mappings of this object, the memory

manager may use memory_object_destroy.

The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) memory object name is a port used by the kernel to refer to the memory object data in reponse to vm_region calls. memory_object_page_size is the page size to be used by this kernel. All data sizes in calls involving this kernel must be an integral multiple of the page size. Note that different kernels, indicated by different memory_controls, may have different page sizes.

The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object ready

(mach_port_t memory control, boolean_t may cache object,

    memory_object_copy_strategy_t copy strategy)

The function memory_object_ready informs the kernel that the memory manager is ready to receive data or unlock requests on behalf of the clients. The argument memory control is the port, provided by the kernel in a memory_object_init call, to which cache management requests may be issued. If may cache object is set, the kernel may keep data associated with this memory object, even after virtual memory references to it are gone.

copy strategy tells how the kernel should copy regions of the associated memory object. There are three possible caching strategies: MEMORY_OBJECT_COPY_NONE which specifies that nothing special should be done when data in the object is copied; MEMORY_OBJECT_COPY_CALL which specifies that the memory manager should be notified via a memory_object_copy call before any part of the object is copied; and MEMORY_OBJECT_COPY_DELAY which guarantees that the memory manager does not externally modify the data so that the kernel can use its normal copy-on-write algorithms. MEMORY_OBJECT_COPY_DELAY is the strategy most commonly used.

This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

6.3 Memory Object Termination

Function kern_return_t memory object terminate

(mach_port_t memory object, mach_port_t memory control, mach_port_t memory object name)

Function kern_return_t seqnos memory object terminate

    (mach_port_t memory object, mach_port_seqno_t seqno,

mach_port_t memory control, mach_port_t memory object name) The function memory_object_terminate indicates that the kernel has completed its use of the given memory object. All rights to the memory object control and name ports are included, so that the memory manager can destroy them (using mach_ port_deallocate) after doing appropriate bookkeeping. The kernel will terminate a memory object only after all address space mappings of that memory object have been deallocated, or upon explicit request by the memory manager.

The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) memory object name is a port used by the kernel to refer to the memory object data in reponse to vm_region calls.

The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object destroy

    (mach_port_t memory control, kern_return_t reason)

The function memory_object_destroy tells the kernel to shut down the memory object. As a result of this call the kernel will no longer support paging activity or any memory_object calls on this object, and all rights to the memory object port, the memory control port and the memory name port will be returned to the memory manager in a memory object terminate call. If the memory manager is concerned that any modified cached data be returned to it before the object is terminated, it should call memory_object_lock_request with should flush set and a lock value of VM_PROT_WRITE before making this call.

The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. reason is an error code indicating why the object must be destroyed.

This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

6.4 Memory Objects and Data

Function kern_return_t memory object data return

(mach_port_t memory object, mach_port_t memory control,

    vm_offset_t offset, vm_offset_t data, vm_size_t data count,
    boolean_t dirty, boolean_t kernel copy)

Function kern_return_t seqnos memory object data return

    (mach_port_t memory object, mach_port_seqno_t seqno,
    mach_port_t memory control, vm_offset_t offset, vm_offset_t data,
    vm_size_t data count, boolean_t dirty, boolean_t kernel copy)

The function memory_object_data_return provides the memory manager with data that has been modified while cached in physical memory. Once the memory manager no longer needs this data (e.g., it has been written to another storage medium), it should be deallocated using vm_deallocate. The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) offset is the offset within a memory object to which this call refers. This will be page aligned. data is the data which has been modified while cached in physical memory. data count is the amount of data to be written, in bytes. This will be an integral number of memory object pages. The kernel will also use this call to return precious pages. If an unmodified precious age is returned, dirty is set to FALSE, otherwise it is TRUE. If kernel copy is TRUE, the kernel kept a copy of the page. Precious data remains precious if the kernel keeps a copy. The indication that the kernel kept a copy is only a hint if the data is not precious; the cleaned copy may be discarded without further notifying the manager. The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object data request

(mach_port_t memory object, mach_port_t memory control,

    vm_offset_t offset, vm_offset_t length, vm_prot_t desired access)

Function kern_return_t seqnos memory object data request

    (mach_port_t memory object, mach_port_seqno_t seqno,
    mach_port_t memory control, vm_offset_t offset, vm_offset_t length,

vm_prot_t desired access) The function memory_object_data_request is a request for data from the specified memory object, for at least the access specified. The memory manager is expected to return at least the specified data, with as much access as it can allow, using memory_ object_data_supply. If the memory manager is unable to provide the data (for example, because of a hardware error), it may use the memory_object_data_error call. The memory_object_data_unavailable call may be used to tell the kernel to supply zero-filled memory for this region. The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) offset is the offset within a memory object to which this call refers. This will be page aligned. length is the number of bytes of data, starting at offset, to which this call refers. This will be an integral number of memory object pages. desired access is a protection value describing the memory access modes which must be permitted on the specified cached data. One or

more of: VM_PROT_READ, VM_PROT_WRITE or VM_PROT_EXECUTE.

The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object data supply

    (mach_port_t memory control, vm_offset_t offset, vm_offset_t data,
    vm_size_t data count, vm_prot_t lock value, boolean_t precious,
    mach_port_t reply)

The function memory_object_data_supply supplies the kernel with data for the specified memory object. Ordinarily, memory managers should only provide data in reponse to memory_object_data_request calls from the kernel (but they may provide data in advance as desired). When data already held by this kernel is provided again, the new data is ignored. The kernel may not provide any data (or protection) consistency among pages with different virtual page alignments within the same object.

The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. offset is an offset within a memory object in bytes. This must be page aligned. data is the data that is being provided to the kernel. This is a pointer to the data. data count is the amount of data to be provided. Only whole virtual pages of data can be accepted; partial pages will be discarded.

lock value is a protection value indicating those forms of access that should not be permitted to the specified cached data. The lock values must be one or more of the set:

VM_PROT_NONE, VM_PROT_READ, VM_PROT_WRITE, VM_PROT_EXECUTE and VM_PROT_ALL
as defined in ‘mach/vm_prot.h’.

If precious is FALSE, the kernel treats the data as a temporary and may throw it away if it hasn’t been changed. If the precious value is TRUE, the kernel treats its copy as a data repository and promises to return it to the manager; the manager may tell the kernel to throw it away instead by flushing and not cleaning the data (see

memory_object_lock_request).

If reply to is not MACH_PORT_NULL, the kernel will send a completion message to the

provided port (see memory_object_supply_completed).

This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

Function kern_return_t memory object supply completed

(mach_port_t memory object, mach_port_t memory control,

    vm_offset_t offset, vm_size_t length, kern_return_t result,
    vm_offset_t error offset)

Function kern_return_t seqnos memory object supply completed

    (mach_port_t memory object, mach_port_seqno_t seqno,
    mach_port_t memory control, vm_offset_t offset, vm_size_t length,
    kern_return_t result, vm_offset_t error offset)
The function memory_object_supply_completed indicates that a previous memory_

object_data_supply has been completed. Note that this call is made on whatever port was specified in the memory_object_data_supply call; that port need not be the memory object port itself. No reply is expected after this call. The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) offset is the offset within a memory object to which this call refers. length is the length of the data covered by the lock request. The result parameter indicates what happened during the supply. If it is not KERN_SUCCESS, then error offset identifies the first offset at which a problem occurred. The pagein operation stopped at this point. Note that the only failures reported by this mechanism are KERN_MEMORY_PRESENT. All other failures (invalid argument, error on pagein of supplied data in manager’s address space) cause the entire operation to fail.

Function kern_return_t memory object data error

    (mach_port_t memory control, vm_offset_t offset, vm_size_t size,
    kern_return_t reason)

The function memory_object_data_error indicates that the memory manager cannot return the data requested for the given region, specifying a reason for the error. This is typically used when a hardware error is encountered. The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. offset is an offset within a memory object in bytes. This must be page aligned. data is the data that is being provided to the kernel. This is a pointer to the data. size is the amount of cached data (starting at offset) to be handled. This must be an integral number of the memory object page size. reason is an error code indicating what type of error occured. This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

Function kern_return_t memory object data unavailable

    (mach_port_t memory control, vm_offset_t offset, vm_size_t size,
    kern_return_t reason)

The function memory_object_data_unavailable indicates that the memory object does not have data for the given region and that the kernel should provide the data for this range. The memory manager may use this call in three different situations.

1. The object was created by memory_object_create and the kernel has not yet provided data for this range (either via a memory_object_data_initialize or

    a memory_object_data_return for the object.

2. The object was created by an memory_object_data_copy and the kernel should copy this region from the original memory object.

3. The object is a normal user-created memory object and the kernel should supply unlocked zero-filled pages for the range.

The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. offset is an offset within a memory object, in bytes. This must be page aligned. size is the amount of cached data (starting at offset) to be handled. This must be an integral number of the memory object page size.

This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

Function kern_return_t memory object copy

(mach_port_t old memory object, mach_port_t old memory control,

    vm_offset_t offset, vm_size_t length, mach_port_t new memory object)

Function kern_return_t seqnos memory object copy

    (mach_port_t old memory object, mach_port_seqno_t seqno,
    mach_port_t old memory control, vm_offset_t offset, vm_size_t length,

mach_port_t new memory object) The function memory_object_copy indicates that a copy has been made of the specified range of the given original memory object. This call includes only the new memory object itself; a memory_object_init call will be made on the new memory object after the currently cached pages of the original object are prepared. After the memory manager receives the init call, it must reply with the memory_object_ready call to assert the "ready" attribute. The kernel will use the new memory object, control and name ports to refer to the new copy.

This call is made when the original memory object had the caching parameter set to MEMORY_OBJECT_COPY_CALL and a user of the object has asked the kernel to copy it.

Cached pages from the original memory object at the time of the copy operation are handled as follows: Readable pages may be silently copied to the new memory object (with all access permissions). Pages not copied are locked to prevent write access.

The new memory object is temporary, meaning that the memory manager should not change its contents or allow the memory object to be mapped in another client. The memory manager may use the memory_object_data_unavailable call to indicate that the appropriate pages of the original memory object may be used to fulfill the data request.

The argument old memory object is the port that represents the old memory object data. old memory control is the kernel port for the old object. offset is the offset within a memory object to which this call refers. This will be page aligned. length is the number of bytes of data, starting at offset, to which this call refers. This will be an integral number of memory object pages. new memory object is a new memory object created by the kernel; see synopsis for further description. Note that all port rights (including receive rights) are included for the new memory object. The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object data provided

    (mach_port_t memory control, vm_offset_t offset, vm_offset_t data,
    vm_size_t data count, vm_prot_t lock value)

The function memory_object_data_provided supplies the kernel with data for the specified memory object. It is the old form of memory_object_data_supply. Ordinarily, memory managers should only provide data in reponse to memory_object_ data_request calls from the kernel. The lock value specifies what type of access will not be allowed to the data range. The lock values must be one or more of the set:

VM_PROT_NONE, VM_PROT_READ, VM_PROT_WRITE, VM_PROT_EXECUTE and VM_PROT_ALL
as defined in ‘mach/vm_prot.h’.

The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. offset is an offset within a memory object in bytes. This must be page aligned. data is the data that is being provided to the kernel. This is a pointer to the data. data count is the amount of data to be provided. This must be an integral number of memory object pages. lock value is a protection value indicating those forms of access that should not be permitted to the specified cached data. This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

6.5 Memory Object Locking

Function kern_return_t memory object lock request

    (mach_port_t memory control, vm_offset_t offset, vm_size_t size,
    memory_object_return_t should clean, boolean_t should flush,
    vm_prot_t lock value, mach_port_t reply to)

The function memory_object_lock_request allows a memory manager to make cache management requests. As specified in arguments to the call, the kernel will:

clean (i.e., write back using memory_object_data_supply) any cached data which has been modified since the last time it was written flush (i.e., remove any uses of) that data from memory lock (i.e., prohibit the specified uses of) the cached data

Locks applied to cached data are not cumulative; new lock values override previous ones. Thus, data may also be unlocked using this primitive. The lock values must be

one or more of the following values: VM_PROT_NONE, VM_PROT_READ, VM_PROT_WRITE,
VM_PROT_EXECUTE and VM_PROT_ALL as defined in ‘mach/vm_prot.h’.

Only data which is cached at the time of this call is affected. When a running thread requires a prohibited access to cached data, the kernel will issue a memory_object_ data_unlock call specifying the forms of access required.

Once all of the actions requested by this call have been completed, the kernel issues

a memory_object_lock_completed call on the specified reply port.

The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. offset is an offset within a memory object, in bytes. This must be page aligned. size is the amount of cached data (starting at offset) to be handled. This must be an integral number of the memory object page size. If should clean is set, modified data should be written back to the memory manager. If should flush is set, the specified cached data should be invalidated, and all uses of that data should be revoked. lock value is a protection value indicating those forms of access that should not be permitted to the specified cached data. reply to is a port on which a memory_object_lock_comleted call should be issued, or MACH_PORT_NULL if no acknowledgement is desired. This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

Function kern_return_t memory object lock completed

(mach_port_t memory object, mach_port_t memory control,

    vm_offset_t offset, vm_size_t length)

Function kern_return_t seqnos memory object lock completed

    (mach_port_t memory object, mach_port_seqno_t seqno,
    mach_port_t memory control, vm_offset_t offset, vm_size_t length)
The function memory_object_lock_completed indicates that a previous memory_

object_lock_request has been completed. Note that this call is made on whatever port was specified in the memory_object_lock_request call; that port need not be the memory object port itself. No reply is expected after this call. The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) offset is the offset within a memory object to which this call refers. length is the length of the data covered by the lock request. The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object data unlock

(mach_port_t memory object, mach_port_t memory control,

    vm_offset_t offset, vm_size_t length, vm_prot_t desired access)

Function kern_return_t seqnos memory object data unlock

    (mach_port_t memory object, mach_port_seqno_t seqno,
    mach_port_t memory control, vm_offset_t offset, vm_size_t length,

vm_prot_t desired access) The function memory_object_data_unlock is a request that the memory manager permit at least the desired access to the specified data cached by the kernel. A call

to memory_object_lock_request is expected in response.

The argument memory object is the port that represents the memory object data, as supplied to the kernel in a vm_map call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) offset is the offset within a memory object to which this call refers. This will be page aligned. length is the number of bytes of data, starting at offset, to which this call refers. This will be an integral number of memory object pages. desired access a protection value describing the memory access modes which must be permitted on the specified cached data. One

or more of: VM_PROT_READ, VM_PROT_WRITE or VM_PROT_EXECUTE.

The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

6.6 Memory Object Attributes

Function kern_return_t memory object get attributes

(mach_port_t memory control, boolean_t *object ready,

boolean_t *may cache object,

    memory_object_copy_strategy_t *copy strategy)

The function memory_object_get_attribute retrieves the current attributes associated with the memory object. The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. If object ready is set, the kernel may issue new data and unlock requests on the associated memory object. If may cache object is set, the kernel may keep data associated with this memory object, even after virtual memory references to it are gone. copy strategy tells how the kernel should copy regions of the associated memory object. This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

Function kern_return_t memory object change attributes

(mach_port_t memory control, boolean_t may cache object,

    memory_object_copy_strategy_t copy strategy, mach_port_t reply to)

The function memory_object_change_attribute sets performance-related attributes for the specified memory object. If the caching attribute is asserted, the kernel is permitted (and encouraged) to maintain cached data for this memory object even after no virtual address space contains this data. There are three possible caching strategies: MEMORY_OBJECT_COPY_NONE which specifies that nothing special should be done when data in the object is copied; MEMORY_ OBJECT_COPY_CALL which specifies that the memory manager should be notified via a memory_object_copy call before any part of the object is copied; and MEMORY_ OBJECT_COPY_DELAY which guarantees that the memory manager does not externally modify the data so that the kernel can use its normal copy-on-write algorithms. MEMORY_OBJECT_COPY_DELAY is the strategy most commonly used. The argument memory control is the port, provided by the kernel in a memory_ object_init call, to which cache management requests may be issued. If may cache object is set, the kernel may keep data associated with this memory object, even after virtual memory references to it are gone. copy strategy tells how the kernel should copy regions of the associated memory object. reply to is a port on which a memory_ object_change_comleted call will be issued upon completion of the attribute change, or MACH_PORT_NULL if no acknowledgement is desired. This routine does not receive a reply message (and consequently has no return value), so only message transmission errors apply.

Function kern_return_t memory object change completed

(mach_port_t memory object, boolean_t may cache object,

    memory_object_copy_strategy_t copy strategy)

Function kern_return_t seqnos memory object change completed

    (mach_port_t memory object, mach_port_seqno_t seqno,

boolean_t may cache object,

    memory_object_copy_strategy_t copy strategy)

The function memory_object_change_completed indicates the completion of an attribute change call.

6.7 Default Memory Manager

Function kern_return_t vm set default memory manager

    (mach_port_t host, mach_port_t *default manager)

The function vm_set_default_memory_manager sets the kernel’s default memory manager. It sets the port to which newly-created temporary memory objects are delivered by memory_object_create to the host. The old memory manager port is returned. If default manager is MACH_PORT_NULL then this routine just returns the current default manager port without changing it. The argument host is a task port to the kernel whose default memory manager is to be changed. default manager is an in/out parameter. As input, default manager is the port that the new memory manager is listening on for memory_object_create calls. As output, it is the old default memory manager’s port. The function returns KERN_SUCCESS if the new memory manager is installed, and KERN_INVALID_ARGUMENT if this task does not have the privileges required for this call.

Function kern_return_t memory object create

(mach_port_t old memory object, mach_port_t new memory object, vm_size_t new object size, mach_port_t new control,

    mach_port_t new name, vm_size_t new page size)

Function kern_return_t seqnos memory object create

    (mach_port_t old memory object, mach_port_seqno_t seqno,

mach_port_t new memory object, vm_size_t new object size,

    mach_port_t new control, mach_port_t new name,

vm_size_t new page size) The function memory_object_create is a request that the given memory manager accept responsibility for the given memory object created by the kernel. This call will only be made to the system default memory manager. The memory object in question initially consists of zero-filled memory; only memory pages that are actually written will ever be provided to memory_object_data_request calls, the default memory manager must use memory_object_data_unavailable for any pages that have not previously been written. No reply is expected after this call. Since this call is directed to the default memory manager, the kernel assumes that it will be ready to handle data requests to this object and does not need the confirmation of a memory_object_set_attributes call. The argument old memory object is a memory object provided by the default memory manager on which the kernel can make memory_object_create calls. new memory object is a new memory object created by the kernel; see synopsis for further description. Note that all port rights (including receive rights) are included for the new memory object. new object size is the maximum size of the new object. new control is a port, created by the kernel, on which a memory manager may issue cache management requests for the new object. new name a port used by the kernel to refer to the new memory object data in response to vm_region calls. new page size is the page size to be used by this kernel. All data sizes in calls involving this kernel must be an integral multiple of the page size. Note that different kernels, indicated by different memory_controls, may have different page sizes. The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

Function kern_return_t memory object data initialize

(mach_port_t memory object, mach_port_t memory control,

    vm_offset_t offset, vm_offset_t data, vm_size_t data count)

Function kern_return_t seqnos memory object data initialize

    (mach_port_t memory object, mach_port_seqno_t seqno,
    mach_port_t memory control, vm_offset_t offset, vm_offset_t data,

vm_size_t data count) The function memory_object_data_initialize provides the memory manager with initial data for a kernel-created memory object. If the memory manager already

has been supplied data (by a previous memory_object_data_initialize or memory_

object_data_return), then this data should be ignored. Otherwise, this call behaves exactly as does memory_object_data_return on memory objects created by the kernel via memory_object_create and thus will only be made to default memory managers. This call will not be made on objects created via memory_object_copy. The argument memory object the port that represents the memory object data, as supplied by the kernel in a memory_object_create call. memory control is the request port to which a response is requested. (In the event that a memory object has been supplied to more than one the kernel that has made the request.) offset is the offset within a memory object to which this call refers. This will be page aligned. data os the data which has been modified while cached in physical memory. data count is the amount of data to be written, in bytes. This will be an integral number of memory object pages. The function should return KERN_SUCCESS, but since this routine is called by the kernel, which does not wait for a reply message, this value is ignored.

7 Threads and Tasks

7.1 Thread Interface

7.1.1 Thread Creation

Function kern_return_t thread create (mach_port_t parent task,

    mach_port_t *child thread)

The function thread_create creates a new thread within the task specified by parent task. The new thread has no processor state, and has a suspend count of 1. To get a new thread to run, first thread_create is called to get the new thread’s identifier, (child thread). Then thread_set_state is called to set a processor state, and finally thread_resume is called to get the thread scheduled to execute. When the thread is created send rights to its thread kernel port are given to it and returned to the caller in child thread. The new thread’s exception port is set to

MACH_PORT_NULL.

The function returns KERN_SUCCESS if a new thread has been created, KERN_INVALID_ ARGUMENT if parent task is not a valid task and KERN_RESOURCE_SHORTAGE if some critical kernel resource is not available.

7.1.2 Thread Termination

Function kern_return_t thread terminate (mach_port_t target thread)

The function thread_terminate destroys the thread specified by target thread. The function returns KERN_SUCCESS if the thread has been killed and KERN_INVALID_ ARGUMENT if target thread is not a thread.

7.1.3 Thread Information

Function mach_port_t mach thread self ()

The mach_thread_self system call returns the calling thread’s thread port. mach_thread_self has an effect equivalent to receiving a send right for the thread port. mach_thread_self returns the name of the send right. In particular, successive calls will increase the calling task’s user-reference count for the send right. As a special exception, the kernel will happily overrun the user reference count of the thread name port, so that this function can not fail for that reason. Because of this, the user should not deallocate the port right if an overrun might have happened. Otherwise the reference count could drop to zero and the send right be destroyed while the user still expects to be able to use it. As the kernel does not make use of the number of extant send rights anyway, this is safe to do (the thread port itself is not destroyed, even when there are no send rights anymore). The function returns MACH_PORT_NULL if a resource shortage prevented the reception of the send right or if the thread port is currently null and MACH_PORT_DEAD if the thread port is currently dead.

Function kern_return_t thread info (mach_port_t target thread, int flavor,

    thread_info_t thread info, mach_msg_type_number_t *thread infoCnt)

The function thread_info returns the selected information array for a thread, as specified by flavor. thread info is an array of integers that is supplied by the caller and returned filled with specified information. thread infoCnt is supplied as the maximum number of integers in thread info. On return, it contains the actual number of integers in thread info. The maximum number of integers by any flavor is THREAD_INFO_MAX. The type of information returned is defined by flavor, which can be one of the following:

THREAD_BASIC_INFO

The function returns basic information about the thread, as defined by thread_basic_info_t. This includes the user and system time, the run state, and scheduling priority. The number of integers returned is

          THREAD_BASIC_INFO_COUNT.
THREAD_SCHED_INFO

The function returns information about the schduling policy for the thread as defined by thread_sched_info_t. The number of integers

          returned is THREAD_SCHED_INFO_COUNT.

The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if target thread is not a thread or flavor is not recognized. The function returns MIG_ ARRAY_TOO_LARGE if the returned info array is too large for thread info. In this case, thread info is filled as much as possible and thread infoCnt is set to the number of elements that would have been returned if there were enough room.

Data type struct thread basic info

This structure is returned in thread info by the thread_info function and provides basic information about the thread. You can cast a variable of type thread_info_ t to a pointer of this type if you provided it as the thread info parameter for the THREAD_BASIC_INFO flavor of thread_info. It has the following members:

time_value_t user_time

user run time

time_value_t system_time

system run time

int cpu_usage

Scaled cpu usage percentage. The scale factor is TH_USAGE_SCALE.

int base_priority

base scheduling priority

int cur_priority

current scheduling priority

int run_state

The run state. The possible vlues of this field are:

TH_STATE_RUNNING

thread is running normally

TH_STATE_STOPPED

thread is suspended

TH_STATE_WAITING

thread is waiting normally

TH_STATE_UNINTERRUPTIBLE

thread is in an uninterruptible wait

TH_STATE_HALTED

thread is halted at a clean point

flags Various flags. The possible values of this field are:

TH_FLAGS_SWAPPED

thread is swapped out

TH_FLAGS_IDLE

thread is an idle thread

int suspend_count

suspend count for thread

int sleep_time

number of seconds that thread has been sleeping

time_value_t creation_time

time stamp of creation

Data type thread basic info t

This is a pointer to a struct thread_basic_info.

Data type struct thread sched info

This structure is returned in thread info by the thread_info function and provides schedule information about the thread. You can cast a variable of type thread_ info_t to a pointer of this type if you provided it as the thread info parameter for the THREAD_SCHED_INFO flavor of thread_info. It has the following members:

int policy

scheduling policy

int data associated data

int base_priority

base scheduling priority

int max_priority

max scheduling priority

int cur_priority

current scheduling priority

int depressed

depressed?

int depress_priority

priority depressed from

Data type thread sched info t

This is a pointer to a struct thread_sched_info.

7.1.4 Thread Settings

Function kern_return_t thread wire (mach_port_t host priv,

    mach_port_t thread, boolean_t wired)

The function thread_wire controls the VM privilege level of the thread thread. A VM-privileged thread never waits inside the kernel for memory allocation from the kernel’s free list of pages or for allocation of a kernel stack.

Threads that are part of the default pageout path should be VM-privileged, to prevent system deadlocks. Threads that are not part of the default pageout path should not be VM-privileged, to prevent the kernel’s free list of pages from being exhausted.

The functions returns KERN_SUCCESS if the call succeeded, KERN_INVALID_ARGUMENT if host priv or thread was invalid.

The thread_wire call is actually an RPC to host priv, normally a send right for a privileged host port, but potentially any send right. In addition to the normal diagnostic return codes from the call’s server (normally the kernel), the call may return mach_msg return codes.

7.1.5 Thread Execution

Function kern_return_t thread suspend (mach_port_t target thread)

Increments the thread’s suspend count and prevents the thread from executing any more user level instructions. In this context a user level instruction is either a machine instruction executed in user mode or a system trap instruction including page faults. Thus if a thread is currently executing within a system trap the kernel code may continue to execute until it reaches the system return code or it may supend within the kernel code. In either case, when the thread is resumed the system trap will return. This could cause unpredictible results if the user did a suspend and then altered the user state of the thread in order to change its direction upon a resume. The call thread_abort is provided to allow the user to abort any system call that is in progress in a predictable way.

The suspend count may become greater than one with the effect that it will take more than one resume call to restart the thread.

The function returns KERN_SUCCESS if the thread has been suspended and KERN_ INVALID_ARGUMENT if target thread is not a thread.

Function kern_return_t thread resume (thread_t target thread)

Decrements the threads’s suspend count. If the count becomes zero the thread is resumed. If it is still positive, the thread is left suspended. The suspend count may not become negative.

The function returns KERN_SUCCESS if the thread has been resumed, KERN_FAILURE if the suspend count is already zero and KERN_INVALID_ARGUMENT if target thread is not a thread.

Function kern_return_t thread abort (mach_port_t target thread)

The function thread_abort aborts the kernel primitives: mach_msg, msg_send, msg_ receive and msg_rpc and page-faults, making the call return a code indicating that it was interrupted. The call is interrupted whether or not the thread (or task containing it) is currently suspended. If it is supsended, the thread receives the interupt when it is resumed.

A thread will retry an aborted page-fault if its state is not modified before it

is resumed. msg_send returns SEND_INTERRUPTED; msg_receive returns RCV_
INTERRUPTED; msg_rpc returns either SEND_INTERRUPTED or RCV_INTERRUPTED, de-

pending on which half of the RPC was interrupted.

The main reason for this primitive is to allow one thread to cleanly stop another thread in a manner that will allow the future execution of the target thread to be controlled in a predictable way. thread_suspend keeps the target thread from executing any further instructions at the user level, including the return from a system call. thread_ get_state/thread_set_state allows the examination or modification of the user state of a target thread. However, if a suspended thread was executing within a system call, it also has associated with it a kernel state. This kernel state can not be modified by thread_set_state with the result that when the thread is resumed the system call may return changing the user state and possibly user memory. thread_ abort aborts the kernel call from the target thread’s point of view by resetting the kernel state so that the thread will resume execution at the system call return with the return code value set to one of the interrupted codes. The system call itself will either be entirely completed or entirely aborted, depending on the precise moment at which the abort was received. Thus if the thread’s user state has been changed by thread_set_state, it will not be modified by any unexpected system call side effects.

For example to simulate a Unix signal, the following sequence of calls may be used:

1. thread_suspend: Stops the thread.

2. thread_abort: Interrupts any system call in progress, setting the return value to ‘interrupted’. Since the thread is stopped, it will not return to user code.

3. thread_set_state: Alters thread’s state to simulate a procedure call to the signal handler

4. thread_resume: Resumes execution at the signal handler. If the thread’s stack has been correctly set up, the thread may return to the interrupted system call. (Of course, the code to push an extra stack frame and change the registers is VERY machine-dependent.)

Calling thread_abort on a non-suspended thread is pretty risky, since it is very difficult to know exactly what system trap, if any, the thread might be executing and whether an interrupt return would cause the thread to do something useful. The function returns KERN_SUCCESS if the thread received an interrupt and KERN_ INVALID_ARGUMENT if target thread is not a thread.

Function kern_return_t thread get state (mach_port_t target thread,

    int flavor, thread_state_t old state,
    mach_msg_type_number_t *old stateCnt)

The function thread_get_state returns the execution state (e.g. the machine registers) of target thread as specified by flavor. The old state is an array of integers that is provided by the caller and returned filled with the specified information. old stateCnt is input set to the maximum number of integers in old state and returned equal to the actual number of integers in old state. target thread may not be mach_thread_self(). The definition of the state structures can be found in ‘machine/thread_status.h’. The function returns KERN_SUCCESS if the state has been returned, KERN_INVALID_ ARGUMENT if target thread is not a thread or is thread_self or flavor is unrecogized for this machine. The function returns MIG_ARRAY_TOO_LARGE if the returned state is too large for old state. In this case, old state is filled as much as possible and old stateCnt is set to the number of elements that would have been returned if there were enough room.

Function kern_return_t thread set state (mach_port_t target thread,

    int flavor, thread_state_t new state,
    mach_msg_type_number_t new stateCnt)

The function thread_set_state sets the execution state (e.g. the machine registers) of target thread as specified by flavor. The new state is an array of integers. new stateCnt is the number of elements in new state. The entire set of registers is reset. This will do unpredictable things if target thread is not suspended. target thread may not be thread_self. The definition of the state structures can be found in ‘machine/thread_status.h’. The function returns KERN_SUCCESS if the state has been set and KERN_INVALID_ ARGUMENT if target thread is not a thread or is thread_self or flavor is unrecogized for this machine.

7.1.6 Scheduling

7.1.6.1 Thread Priority

Threads have three priorities associated with them by the system, a priority, a maximum priority, and a scheduled priority. The scheduled priority is used to make scheduling decisions about the thread. It is determined from the priority by the policy (for timesharing, this means adding an increment derived from cpu usage). The priority can be set under user control, but may never exceed the maximum priority. Changing the maximum priority requires presentation of the control port for the thread’s processor set; since the control port for the default processor set is privileged, users cannot raise their maximum priority to unfairly compete with other users on that set. Newly created threads obtain their priority from their task and their max priority from the thread.

Function kern_return_t thread priority (mach_port_t thread, int prority,

    boolean_t set max)

The function thread_priority changes the priority and optionally the maximum priority of thread. Priorities range from 0 to 31, where lower numbers denote higher priorities. If the new priority is higher than the priority of the current thread, preemption may occur as a result of this call. The maximum priority of the thread is also set if set max is TRUE. This call will fail if priority is greater than the current maximum priority of the thread. As a result, this call can only lower the value of a thread’s maximum priority.

The functions returns KERN_SUCCESS if the operation completed successfully, KERN_ INVALID_ARGUMENT if thread is not a thread or priority is out of range (not in 0..31), and KERN_FAILURE if the requested operation would violate the thread’s maximum priority (thread priority).

Function kern_return_t thread max priority (mach_port_t thread,

mach_port_t processor set, int priority) The function thread_max_priority changes the maximum priority of the thread. Because it requires presentation of the corresponding processor set port, this call can reset the maximum priority to any legal value.

The functions returns KERN_SUCCESS if the operation completed successfully, KERN_ INVALID_ARGUMENT if thread is not a thread or processor set is not a control port for a processor set or priority is out of range (not in 0..31), and KERN_FAILURE if the thread is not assigned to the processor set whose control port was presented.

7.1.6.2 Hand-Off Scheduling

Function kern_return_t thread switch (thread_t new thread, int option,

int time) The function thread_switch provides low-level access to the scheduler’s context switching code. new thread is a hint that implements hand-off scheduling. The operating system will attempt to switch directly to the new thread (by passing the normal logic that selects the next thread to run) if possible. Since this is a hint, it may be incorrect; it is ignored if it doesn’t specify a thread on the same host as the current thread or if that thread can’t be switched to (i.e., not runnable or already running on another processor). In this case, the normal logic to select the next thread to run is used; the current thread may continue running if there is no other appropriate thread to run.

Options for option are defined in ‘mach/thread_switch.h’ and specify the interpretation of time. The possible values for option are:

SWITCH_OPTION_NONE

No options, the time argument is ignored.

SWITCH_OPTION_WAIT

The thread is blocked for the specified time. This can be aborted by

          thread_abort.
SWITCH_OPTION_DEPRESS

The thread’s priority is depressed to the lowest possible value for the specified time. This can be aborted by thread_depress_abort. This depression is independent of operations that change the thread’s priority (e.g. thread_priority will not abort the depression). The minimum time and units of time can be obtained as the min_timeout value from host_info. The depression is also aborted when the current thread is next run (either via handoff scheduling or because the processor set has nothing better to do).

thread_switch is often called when the current thread can proceed no further for some reason; the various options and arguments allow information about this reason to be transmitted to the kernel. The new thread argument (handoff scheduling) is useful when the identity of the thread that must make progress before the current thread runs again is known. The WAIT option is used when the amount of time that the current thread must wait before it can do anything useful can be estimated and is fairly long. The DEPRESS option is used when the amount of time that must be waited is fairly short, especially when the identity of the thread that is being waited for is not known. Users should beware of calling thread_switch with an invalid hint (e.g. MACH_PORT_ NULL) and no option. Because the time-sharing scheduler varies the priority of threads based on usage, this may result in a waste of cpu time if the thread that must be run is of lower priority. The use of the DEPRESS option in this situation is highly recommended. thread_switch ignores policies. Users relying on the preemption semantics of a fixed time policy should be aware that thread_switch ignores these semantics; it will run the specified new thread indepent of its priority and the priority of any other threads that could be run instead. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_ARGUMENT if thread is not a thread or option is not a recognized option, and KERN_FAILURE if kern_depress_abort failed because the thread was not depressed.

Function kern_return_t thread depress abort (mach_port_t thread)

The function thread_depress_abort cancels any priority depression for thread

caused by a swtch_pri or thread_switch call.

The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if thread is not a valid thread.

Function boolean_t swtch ()

XXX FIXME

Function boolean_t swtch pri (int priority)

XXX FIXME

7.1.6.3 Scheduling Policy

Function kern_return_t thread policy (mach_port_t thread, int policy,

int data) The function thread_policy changes the scheduling policy for thread to policy.

data is policy-dependent scheduling information. There are currently two supported

policies: POLICY_TIMESHARE and POLICY_FIXEDPRI defined in ‘mach/policy.h’; this

file is included by ‘mach.h’. data is meaningless for timesharing, but is the quantum to be used (in milliseconds) for the fixed priority policy. To be meaningful, this quantum must be a multiple of the basic system quantum (min quantum) which can be obtained from host_info. The system will always round up to the next multiple of the quantum.

Processor sets may restrict the allowed policies, so this call will fail if the processor set to which thread is currently assigned does not permit policy.

The function returns KERN_SUCCESS if the call succeeded. KERN_INVALID_ARGUMENT if thread is not a thread or policy is not a recognized policy, and KERN_FAILURE if the processor set to which thread is currently assigned does not permit policy.

7.1.7 Thread Special Ports

Function kern_return_t thread get special port (thread_t thread,

int which port, mach_port_t *special port) The function thread_get_special_port returns send rights to one of a set of special ports for the thread specified by thread.

The possible values for which port are THREAD_KERNEL_PORT and THREAD_EXCEPTION_

PORT. A thread also has access to its task’s special ports.

The function returns KERN_SUCCESS if the port was returned and KERN_INVALID_ ARGUMENT if thread is not a thread or which port is an invalid port selector.

Function kern_return_t thread get kernel port (thread_t thread,

    mach_port_t *kernel port)

The function thread_get_kernel_port is equivalent to the function thread_get_

special_port with the which port argument set to THREAD_KERNEL_PORT.

Function kern_return_t thread get exception port (thread_t thread,

mach_port_t *exception port) The function thread_get_exception_port is equivalent to the function thread_

get_special_port with the which port argument set to THREAD_EXCEPTION_PORT.

Function kern_return_t thread set special port (thread_t thread,

int which port, mach_port_t special port) The function thread_set_special_port sets one of a set of special ports for the thread specified by thread.

The possible values for which port are THREAD_KERNEL_PORT and THREAD_EXCEPTION_

PORT. A thread also has access to its task’s special ports. The function returns KERN_SUCCESS if the port was set and KERN_INVALID_ARGUMENT if thread is not a thread or which port is an invalid port selector.

Function kern_return_t thread set kernel port (thread_t thread,

    mach_port_t kernel port)

The function thread_set_kernel_port is equivalent to the function thread_set_

special_port with the which port argument set to THREAD_KERNEL_PORT.

Function kern_return_t thread set exception port (thread_t thread,

mach_port_t exception port) The function thread_set_exception_port is equivalent to the function thread_

set_special_port with the which port argument set to THREAD_EXCEPTION_PORT.

7.1.8 Exceptions

Function kern_return_t catch exception raise

    (mach_port_t exception port, mach_port_t thread, mach_port_t task,

int exception, int code, int subcode) XXX NOT DOCUMENTED

Function kern_return_t exception raise (mach_port_t exception port,

    mach_port_t thread, mach_port_t task, int exception, int code,

int subcode) XXX NOT DOCUMENTED

Function kern_return_t evc wait (unsigned int event)

XXX NOT DOCUMENTED & NOT THE RIGHT PLACE

7.2 Task Interface

7.2.1 Task Creation

Function kern_return_t task create (mach_port_t parent task,

boolean_t inherit memory, mach_port_t *child task) The function task_create creates a new task from parent task; the resulting task (child task) acquires shared or copied parts of the parent’s address space (see vm_ inherit). The child task initially contains no threads.

If inherit memory is set, the child task’s address space is built from the parent task according to its memory inheritance values; otherwise, the child task is given an empty address space. The child task gets the three special ports created or copied for it at task creation. The TASK_KERNEL_PORT is created and send rights for it are given to the child and returned to the caller. The TASK_BOOTSTRAP_PORT and the TASK EXCEPTION PORT are inherited from the parent task. The new task can get send rights to these ports with

the call task_get_special_port.

The function returns KERN_SUCCESS if a new task has been created, KERN_INVALID_ ARGUMENT if parent task is not a valid task port and KERN_RESOURCE_SHORTAGE if some critical kernel resource is unavailable.

7.2.2 Task Termination

Function kern_return_t task terminate (mach_port_t target task)

The function task_terminate destroys the task specified by target task and all its threads. All resources that are used only by this task are freed. Any port to which this task has receive and ownership rights is destroyed. The function returns KERN_SUCCESS if the task has been killed, KERN_INVALID_ ARGUMENT if target task is not a task.

7.2.3 Task Information

Function mach_port_t mach task self ()

The mach_task_self system call returns the calling thread’s task port. mach_task_self has an effect equivalent to receiving a send right for the task port. mach_task_self returns the name of the send right. In particular, successive calls will increase the calling task’s user-reference count for the send right. As a special exception, the kernel will happily overrun the user reference count of the task name port, so that this function can not fail for that reason. Because of this, the user should not deallocate the port right if an overrun might have happened. Otherwise the reference count could drop to zero and the send right be destroyed while the user still expects to be able to use it. As the kernel does not make use of the number of extant send rights anyway, this is safe to do (the task port itself is not destroyed, even when there are no send rights anymore). The funcion returns MACH_PORT_NULL if a resource shortage prevented the reception of the send right, MACH_PORT_NULL if the task port is currently null, MACH_PORT_DEAD if the task port is currently dead.

Function kern_return_t task threads (mach_port_t target task,

    thread_array_t *thread list, mach_msg_type_number_t *thread count)

The function task_threads gets send rights to the kernel port for each thread contained in target task. thread list is an array that is created as a result of this call. The caller may wish to vm_deallocate this array when the data is no longer needed.

The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if target task is not a task.

Function kern_return_t task info (mach_port_t target task, int flavor,

    task_info_t task info, mach_msg_type_number_t *task infoCnt)

The function task_info returns the selected information array for a task, as specified by flavor. task info is an array of integers that is supplied by the caller, and filled with specified information. task infoCnt is supplied as the maximum number of integers in task info. On return, it contains the actual number of integers in task info. The maximum number of integers by any flavor is TASK_INFO_MAX. The type of information returned is defined by flavor, which can be one of the following:

TASK_BASIC_INFO

The function returns basic information about the task, as defined by task_basic_info_t. This includes the user and system time and memory consumption. The number of integers returned is TASK_BASIC_INFO_

          COUNT.
TASK_EVENTS_INFO

The function returns information about events for the task as defined by thread_sched_info_t. This includes statistics about virtual memory and IPC events like pageouts, pageins and messages sent and received. The number of integers returned is TASK_EVENTS_INFO_COUNT.

TASK_THREAD_TIMES_INFO

The function returns information about the total time for live threads as defined by task_thread_times_info_t. The number of integers re-

          turned is TASK_THREAD_TIMES_INFO_COUNT.

The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if target task is not a thread or flavor is not recognized. The function returns MIG_ ARRAY_TOO_LARGE if the returned info array is too large for task info. In this case, task info is filled as much as possible and task infoCnt is set to the number of elements that would have been returned if there were enough room.

Data type struct task basic info

This structure is returned in task info by the task_info function and provides basic information about the task. You can cast a variable of type task_info_t to a pointer of this type if you provided it as the task info parameter for the TASK_BASIC_INFO flavor of task_info. It has the following members:

integer_t suspend_count

suspend count for task

integer_t base_priority

base scheduling priority

vm_size_t virtual_size

number of virtual pages

vm_size_t resident_size

number of resident pages

time_value_t user_time

total user run time for terminated threads

time_value_t system_time

total system run time for terminated threads

time_value_t creation_time

creation time stamp

Data type task basic info t

This is a pointer to a struct task_basic_info.

Data type struct task events info

This structure is returned in task info by the task_info function and provides event statistics for the task. You can cast a variable of type task_info_t to a pointer of this type if you provided it as the task info parameter for the TASK_EVENTS_INFO flavor of task_info. It has the following members:

natural_t faults

number of page faults

natural_t zero_fills

number of zero fill pages

natural_t reactivations

number of reactivated pages

natural_t pageins

number of actual pageins

natural_t cow_faults

number of copy-on-write faults

natural_t messages_sent

number of messages sent

natural_t messages_received

number of messages received

Data type task events info t

This is a pointer to a struct task_events_info.

Data type struct task thread times info

This structure is returned in task info by the task_info function and provides event statistics for the task. You can cast a variable of type task_info_t to a pointer of this type if you provided it as the task info parameter for the TASK_THREAD_TIMES_INFO flavor of task_info. It has the following members:

time_value_t user_time

total user run time for live threads

time_value_t system_time

total system run time for live threads

Data type task thread times info t

This is a pointer to a struct task_thread_times_info.

7.2.4 Task Execution

Function kern_return_t task suspend (mach_port_t target task)

The function task_suspend increments the task’s suspend count and stops all threads in the task. As long as the suspend count is positive newly created threads will not run. This call does not return until all threads are suspended. The count may become greater than one, with the effect that it will take more than one resume call to restart the task. The function returns KERN_SUCCESS if the task has been suspended and KERN_ INVALID_ARGUMENT if target task is not a task.

Function kern_return_t task resume (mach_port_t target task)

The function task_resume decrements the task’s suspend count. If it becomes zero, all threads with zero suspend counts in the task are resumed. The count may not become negative. The function returns KERN_SUCCESS if the task has been resumed, KERN_FAILURE if the suspend count is already at zero and KERN_INVALID_ARGUMENT if target task is not a task.

Function kern_return_t task priority (mach_port_t task, int priority,

boolean_t change threads) The priority of a task is used only for creation of new threads; a new thread’s priority is set to the enclosing task’s priority. task_priority changes this task priority. It also sets the priorities of all threads in the task to this new priority if change threads is TRUE. Existing threads are not affected otherwise. If this priority change violates the maximum priority of some threads, as many threads as possible will be changed and an error code will be returned. The function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_ARGUMENT if task is not a task, or priority is not a valid priority and KERN_FAILURE if change threads was TRUE and the attempt to change the priority of at least one existing thread failed because the new priority would have exceeded that thread’s maximum priority.

Function kern_return_t task ras control (task_t target task,

    vm_offset_t start pc, vm_offset_t end pc, int flavor)

The function task_ras_control manipulates a task’s set of restartable atomic sequences. If a sequence is installed, and any thread in the task is preempted within the range [start pc,end pc], then the thread is resumed at start pc. This enables applications to build atomic sequences which, when executed to completion, will have executed atomically. Restartable atomic sequences are intended to be used on systems that do not have hardware support for low-overhead atomic primitives. As a thread can be rolled-back, the code in the sequence should have no side effects other than a final store at end pc. The kernel does not guarantee that the sequence is restartable. It assumes the application knows what it’s doing. A task may have a finite number of atomic sequences that is defined at compile time. The flavor specifices the particular operation that should be applied to this restartable atomic sequence. Possible values for flavor can be:

TASK_RAS_CONTROL_PURGE_ALL

Remove all registered sequences for this task.

TASK_RAS_CONTROL_PURGE_ONE

Remove the named registered sequence for this task.

TASK_RAS_CONTROL_PURGE_ALL_AND_INSTALL_ONE

Atomically remove all registered sequences and install the named sequence.

TASK_RAS_CONTROL_INSTALL_ONE

Install this sequence.

The function returns KERN_SUCCESS if the operation has been performed, KERN_ INVALID_ADDRESS if the start pc or end pc values are not a valid address for the requested operation (for example, it is invalid to purge a sequence that has not been registered), KERN_RESOURCE_SHORTAGE if an attempt was made to install more restartable atomic sequences for a task than can be supported by the kernel, KERN_ INVALID_VALUE if a bad flavor was specified, KERN_INVALID_ARGUMENT if target task is not a task and KERN_FAILURE if the call is not not supported on this configuration.

7.2.5 Task Special Ports

Function kern_return_t task get special port (mach_port_t task,

int which port, mach_port_t *special port) The function task_get_special_port returns send rights to one of a set of special ports for the task specified by task. The special ports associated with a task are the kernel port (TASK_KERNEL_PORT), the bootstrap port (TASK_BOOTSTRAP_PORT) and the exception port (TASK_EXCEPTION_ PORT). The bootstrap port is a port to which a The bootstrap port is a port to which a thread may send a message requesting other system service ports. This port is not used by the kernel. The task’s exception port is the port to which messages are sent by the kernel when an exception occurs and the thread causing the exception has no exception port of its own. The following macros to call task_get_special_port for a specific port are de-

fined in mach/task_special_ports.h: task_get_exception_port and task_get_
bootstrap_port.

The function returns KERN_SUCCESS if the port was returned and KERN_INVALID_ ARGUMENT if task is not a task or which port is an invalid port selector.

Function kern_return_t task get kernel port (task_t task,

    mach_port_t *kernel port)
The function task_get_kernel_port is equivalent to the function task_get_special_

port with the which port argument set to TASK_KERNEL_PORT.

Function kern_return_t task get exception port (task_t task,

mach_port_t *exception port) The function task_get_exception_port is equivalent to the function task_get_

special_port with the which port argument set to TASK_EXCEPTION_PORT.

Function kern_return_t task get bootstrap port (task_t task,

mach_port_t *bootstrap port) The function task_get_bootstrap_port is equivalent to the function task_get_

special_port with the which port argument set to TASK_BOOTSTRAP_PORT.

Function kern_return_t task set special port (mach_port_t task,

int which port, mach_port_t special port) The function thread_set_special_port sets one of a set of special ports for the task specified by task.

The special ports associated with a task are the kernel port (TASK_KERNEL_PORT), the bootstrap port (TASK_BOOTSTRAP_PORT) and the exception port (TASK_EXCEPTION_ PORT). The bootstrap port is a port to which a thread may send a message requesting other system service ports. This port is not used by the kernel. The task’s exception port is the port to which messages are sent by the kernel when an exception occurs and the thread causing the exception has no exception port of its own.

The function returns KERN_SUCCESS if the port was set and KERN_INVALID_ARGUMENT if task is not a task or which port is an invalid port selector.

Function kern_return_t task set kernel port (task_t task,

    mach_port_t kernel port)
The function task_set_kernel_port is equivalent to the function task_set_special_

port with the which port argument set to TASK_KERNEL_PORT.

Function kern_return_t task set exception port (task_t task,

mach_port_t exception port) The function task_set_exception_port is equivalent to the function task_set_

special_port with the which port argument set to TASK_EXCEPTION_PORT.

Function kern_return_t task set bootstrap port (task_t task,

mach_port_t bootstrap port) The function task_set_bootstrap_port is equivalent to the function task_set_

special_port with the which port argument set to TASK_BOOTSTRAP_PORT.

7.2.6 Syscall Emulation

Function kern_return_t task get emulation vector (mach_port_t task,

int *vector start, emulation_vector_t *emulation vector,

    mach_msg_type_number_t *emulation vectorCnt)

The function task_get_emulation_vector gets the user-level handler entry points for all emulated system calls.

Function kern_return_t task set emulation vector (mach_port_t task,

int vector start, emulation_vector_t emulation vector,

    mach_msg_type_number_t emulation vectorCnt)

The function task_set_emulation_vector establishes user-level handlers for the specified system calls. Non-emulated system calls are specified with an entry of EML_ROUTINE_NULL. System call emulation handlers are inherited by the childs of task.

Function kern_return_t task set emulation (mach_port_t task,

vm_address_t routine entry pt, int routine number) The function task_set_emulation establishes a user-level handler for the specified system call. System call emulation handlers are inherited by the childs of task.

7.3 Profiling

Function kern_return_t task enable pc sampling (task_t task,

    int *ticks, sampled_pc_flavor_t flavor)

Function kern_return_t thread enable pc sampling (thread_t thread,

    int *ticks, sampled_pc_flavor_t flavor)

The function task_enable_pc_sampling enables PC sampling for task, the function thread_enable_pc_sampling enables PC sampling for thread. The kernel’s idea of clock granularity is returned in ticks (this value should not be trusted). The sampling flavor is specified by flavor. The function returns KERN_SUCCESS if the operation is completed successfully and KERN_INVALID_ARGUMENT if thread is not a valid thread.

Function kern_return_t task disable pc sampling (task_t task,

int *sample cnt)

Function kern_return_t thread disable pc sampling (thread_t thread,

int *sample cnt) The function task_disable_pc_sampling disables PC sampling for task, the function thread_disable_pc_sampling disables PC sampling for thread. The number of sample elements in the kernel for the thread is returned in sample cnt. The function returns KERN_SUCCESS if the operation is completed successfully and KERN_INVALID_ARGUMENT if thread is not a valid thread.

Function kern_return_t task get sampled pcs (task_t task,

    unsigned *seqno, sampled_pc_t sampled pcs[], int *sample cnt)

Function kern_return_t thread get sampled pcs (thread_t thread,

    unsigned *seqno, sampled_pc_t sampled pcs[], int *sample cnt)

The function task_get_sampled_pcs extracts the PC samples for task, the function thread_get_sampled_pcs extracts the PC samples for thread. seqno is the sequence number of the sampled PCs. This is useful for determining when a collector thread has missed a sample. The sampled PCs for the thread are returned in sampled pcs. sample cnt contains the number of sample elements returned.

The function returns KERN_SUCCESS if the operation is completed successfully, KERN_ INVALID_ARGUMENT if thread is not a valid thread and KERN_FAILURE if thread is not sampled.

Data type sample pc t

This structure is returned in sampled pcs by the thread_get_sampled_pcs and task_get_sampled_pcs functions and provides pc samples for threads or tasks. It has the following members:

natural_t id

A thread-specific unique identifier.

vm_offset_t pc

A pc value.

sampled_pc_flavor_t sampletype

The type of the sample as per flavor.

Data type sample pc flavor t

This data type specifies a pc sample flavor, either as argument passed in flavor to the

thread_enable_pc_sample and thread_disable_pc_sample functions, or as mem-

ber sampletype in the sample_pc_t data type. The flavor is a bitwise-or of the possible flavors defined in ‘mach/pc_sample.h’:

SAMPLED_PC_PERIODIC

default

SAMPLED_PC_VM_ZFILL_FAULTS

zero filled fault

SAMPLED_PC_VM_REACTIVATION_FAULTS

reactivation fault

SAMPLED_PC_VM_PAGEIN_FAULTS

pagein fault

SAMPLED_PC_VM_COW_FAULTS

copy-on-write fault

SAMPLED_PC_VM_FAULTS_ANY

any fault

SAMPLED_PC_VM_FAULTS
the bitwise-or of SAMPLED_PC_VM_ZFILL_FAULTS, SAMPLED_PC_VM_REACTIVATION_
FAULTS, SAMPLED_PC_VM_PAGEIN_FAULTS and SAMPLED_PC_VM_COW_FAULTS.

80 The GNU Mach Reference Manual

8 Host Interface

This section describes the Mach interface to a host executing a Mach kernel. The intrface allows it to query statistics about a host and control its behaviour.

A host is represented by two ports, a name port host of type host_t used to query information about the host accessible to everyone and a control port host priv of type host_priv_t used to manipulate it. For example, you can query the current time over the name port, but to change the time you need to send a message to the host control port.

A send right to the name port of the host a task is running on is available with the mach_host_self system trap. A send right to the host control port is inserted into the first task at bootstrap.

Everything described in this section is declared in the header file ‘mach.h’.

8.1 Host Information

Function mach_port_t mach host self ()

The mach_host_self system call returns the calling thread’s host port. It has an effect equivalent to receiving a send right for the host port. mach_host_self returns the name of the send right. In particular, successive calls will increase the calling task’s user-reference count for the send right. As a special exception, the kernel will happily overrun the user reference count of the host name port, so that this function can not fail for that reason. Because of this, the user should not deallocate the port right if an overrun might have happened. Otherwise the reference count could drop to zero and the send right be destroyed while the user still expects to be able to use it. As the kernel does not make use of the number of extant send rights anyway, this is safe to do (the host port itself is never destroyed). The function returns MACH_PORT_NULL if a resource shortage prevented the reception of the send right. This function is also available in ‘mach/mach_traps.h’.

Function kern_return_t host info (mach_port_t host, int flavor,
    host_info_t host info, mach_msg_type_number_t *host infoCnt)

The host_info function returns various information about host. host info is an array of integers that is supplied by the caller, and filled with specified information. host infoCnt is supplied as the maximum number of integers in host info. On return, it contains the actual number of integers in host info. The type of information returned is defined by flavor, which can be one of the following:

HOST_BASIC_INFO

The function returns basic information about the host, as defined by host_basic_info_t. This includes the number of processors, their type, and the amount of memory installed in the system. The number of inte-

          gers returned is HOST_BASIC_INFO_COUNT.

HOST_PROCESSOR_SLOTS

The function returns the numbers of the slots with active processors in them. The number of integers returned can be up to max_cpus, as re-

          turned by the HOT_BASIC_INFO flavor of host_info.
HOST_SCHED_INFO

The function returns information of interest to schedulers as defined by host_sched_info_t. The number of integers returned is HOST_SCHED_

          INFO_COUNT.

The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if host is not a host or flavor is not recognized. The function returns MIG_ARRAY_ TOO_LARGE if the returned info array is too large for host info. In this case, host info is filled as much as possible and host infoCnt is set to the number of elements that would be returned if there were enough room.

Data type struct host basic info

A pointer to this structure is returned in host info by the host_info function and provides basic information about the host. You can cast a variable of type host_ info_t to a pointer of this type if you provided it as the host info parameter for the HOST_BASIC_INFO flavor of host_info. It has the following members:

int max_cpus

maximum possible cpus for which kernel is configured

int avail_cpus

number of cpus now available

vm_size_t memory_size

size of memory in bytes

cpu_type_t cpu_type

cpu type

cpu_subtype_t cpu_subtype

cpu subtype

Data type host basic info t

This is a pointer to a struct host_basic_info.

Data type struct host sched info

A pointer to this structure is returned in host info by the host_info function and provides information of interest to schedulers. You can cast a variable of type host_ info_t to a pointer of this type if you provided it as the host info parameter for the HOST_SCHED_INFO flavor of host_info. It has the following members:

int min_timeout

minimum timeout in milliseconds

int min_quantum

minimum quantum in milliseconds

Data type host sched info t

This is a pointer to a struct host_sched_info.

Function kern_return_t host kernel version (mach_port_t host,

    kernel_version_t *version)

The host_kernel_version function returns the version string compiled into the kernel executing on host at the time it was built in the character string version. This string describes the version of the kernel. The constant KERNEL_VERSION_MAX should be used to dimension storage for the returned string if the kernel_version_t declaration is not used. If the version string compiled into the kernel is longer than KERNEL VERSION MAX, the result is truncated and not necessarily null-terminated. If host is not a valid send right to a host port, the function returns KERN_INVALID_ ARGUMENT. If version points to inaccessible memory, it returns KERN_INVALID_

ADDRESS, and KERN_SUCCESS otherwise.

Function kern_return_t host get boot info (mach_port_t host priv,

    kernel_boot_info_t boot info)

The host_get_boot_info function returns the boot-time information string supplied by the operator to the kernel executing on host priv in the character string boot info. The constant KERNEL_BOOT_INFO_MAX should be used to dimension storage for the returned string if the kernel_boot_info_t declaration is not used. If the boot-time information string supplied by the operator is longer than KER-NEL BOOT INFO MAX, the result is truncated and not necessarily null-terminated.

8.2 Host Time

Data type time value t

This is the representation of a time in Mach. It is a struct time_value and consists of the following members:

integer_t seconds

The number of seconds.

integer_t microseconds

The number of microseconds.

The number of microseconds should always be smaller than TIME_MICROS_MAX (100000). A time with this property is normalized. Normalized time values can be manipulated with the following macros:

Macro time value add usec (time value t *val, integer t *micros)

Add micros microseconds to val. If val is normalized and micros smaller than TIME_ MICROS_MAX, val will be normalized afterwards.

Macro time value add (time value t *result, time value t *addend)

Add the values in addend to result. If both are normalized, result will be normalized afterwards.

A variable of type time_value_t can either represent a duration or a fixed point in time. In the latter case, it shall be interpreted as the number of seconds and microseconds after the epoch 1. Jan 1970.

Function kern_return_t host get time (mach_port_t host,
    time_value_t *current time)

Get the current time as seen by host. On success, the time passed since the epoch is returned in current time.

Function kern_return_t host set time (mach_port_t host priv,

    time_value_t new time)

Set the time of host priv to new time.

Function kern_return_t host adjust time (mach_port_t host priv,

time_value_t new adjustment, time_value_t *old adjustment) Arrange for the current time as seen by host priv to be gradually changed by the adjustment value new adjustment, and return the old adjustment value in old adjustment.

For efficiency, the current time is available through a mapped-time interface.

Data type mapped time value t

This structure defines the mapped-time interface. It has the following members:

integer_t seconds

The number of seconds.

integer_t microseconds

The number of microseconds.

integer_t check_seconds

This is a copy of the seconds value, which must be used to protect against a race condition when reading out the two time values.

Here is an example how to read out the current time using the mapped-time interface:

do
  {
    secs = mtime->seconds;
    usecs = mtime->microseconds;
  }
while (secs != mtime->check_seconds);

8.3 Host Reboot

Function kern_return_t host reboot (mach_port_t host priv, int options)

Reboot the host specified by host priv. The argument options specifies the flags. The available flags are specified in ‘sys/reboot.h’:

RB_HALT Do not reboot, but halt the machine.

Chapter 8: Host Interface 85

RB_DEBUGGER

Do not reboot, but enter kernel debugger from user space.

If successful, the function might not return.

86 The GNU Mach Reference Manual

9 Processor Interface

9.1 Processor Set Access

Function kern_return_t host processor sets (mach_port_t host,

    processor_set_name_array_t *processor sets,
    mach_msg_type_number_t *processor sets count)

The function host_processor_sets gets send rights to the name port for each processor set currently assigned to host. host_processor_set_priv can be used to obtain the control ports from these if desired. processor sets is an array that is created as a result of this call. The caller may wish to vm_deallocate this array when the data is no longer needed. processor sets count is set to the number of processor sets in the processor sets. This function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ ARGUMENT if host is not a host.

Function kern_return_t host processor set priv (mach_port_t host priv,

    mach_port_t set name, mach_port_t *set)

The function host_processor_set_priv allows a privileged application to obtain the control port set for an existing processor set from its name port set name. The privileged host port host priv is required. This function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ ARGUMENT if host priv is not a valid host control port.

Function kern_return_t processor set default (mach_port_t host,

    mach_port_t *default set)

The function processor_set_default returns the default processor set of host in default set. The default processor set is used by all threads, tasks, and processors that are not explicitly assigned to other sets. processor set default returns a port that can be used to obtain information about this set (e.g. how many threads are assigned to it). This port cannot be used to perform operations on that set. This function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_ARGUMENT if host is not a host and KERN_INVALID_ADDRESS if default set points to inaccessible memory.

9.2 Processor Set Creation

Function kern_return_t processor set create (mach_port_t host,

    mach_port_t *new set, mach_port_t *new name)

The function processor_set_create creates a new processor set on host and returns the two ports associated with it. The port returned in new set is the actual port representing the set. It is used to perform operations such as assigning processors, tasks, or threads. The port returned in new name identifies the set, and is used to obtain information about the set. This function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_ARGUMENT if host is not a host, KERN_INVALID_ADDRESS if new set or new name points to inaccessible memory and KERN_FAILURE is the operating system does not support processor allocation.

9.3 Processor Set Destruction

Function kern_return_t processor set destroy (mach_port_t processor set)

The function processor_set_destroy destroys the specified processor set. Any assigned processors, tasks, or threads are reassigned to the default set. The object port for the processor set is required (not the name port). The default processor set cannot be destroyed. This function returns KERN_SUCCESS if the set was destroyed, KERN_FAILURE if an attempt was made to destroy the default processor set, or the operating system does not support processor allocation, and KERN_INVALID_ARGUMENT if processor set is not a valid processor set control port.

9.4 Tasks and Threads on Sets

Function kern_return_t processor set tasks (mach_port_t processor set,

    task_array_t *task list, mach_msg_type_number_t *task count)

The function processor_set_tasks gets send rights to the kernel port for each task currently assigned to processor set. task list is an array that is created as a result of this call. The caller may wish to vm_deallocate this array when the data is no longer needed. task count is set to the number of tasks in the task list. This function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ ARGUMENT if processor set is not a processor set.

Function kern_return_t processor set threads (mach_port_t processor set,

    thread_array_t *thread list, mach_msg_type_number_t *thread count)

The function processor_set_thread gets send rights to the kernel port for each thread currently assigned to processor set. thread list is an array that is created as a result of this call. The caller may wish to vm_deallocate this array when the data is no longer needed. thread count is set to the number of threads in the thread list. This function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ ARGUMENT if processor set is not a processor set.

Function kern_return_t task assign (mach_port_t task,

mach_port_t processor set, boolean_t assign threads) The function task_assign assigns task the set processor set. This assignment is for the purposes of determining the initial assignment of newly created threads in task.

Any previous assignment of the task is nullified. Existing threads within the task are also reassigned if assign threads is TRUE. They are not affected if it is FALSE.

This function returns KERN_SUCCESS if the assignment has been performed and KERN_ INVALID_ARGUMENT if task is not a task, or processor set is not a processor set on the same host as task.

Function kern_return_t task assign default (mach_port_t task,

boolean_t assign threads) The function task_assign_default is a variant of task_assign that assigns the task to the default processor set on that task’s host. This variant exists because the control port for the default processor set is privileged and not ususally available to users.

This function returns KERN_SUCCESS if the assignment has been performed and KERN_ INVALID_ARGUMENT if task is not a task.

Function kern_return_t task get assignment (mach_port_t task,

mach_port_t *processor set) The function task_get_assignment returns the name of the processor set to which the thread is currently assigned in processor set. This port can only be used to obtain information about the processor set.

This function returns KERN_SUCCESS if the assignment has been performed, KERN_ INVALID_ADDRESS if processor set points to inaccessible memory, and KERN_INVALID_ ARGUMENT if task is not a task.

Function kern_return_t thread assign (mach_port_t thread,

mach_port_t processor set) The function thread_assign assigns thread the set processor set. After the assignment is completed, the thread only executes on processors assigned to the designated processor set. If there are no such processors, then the thread is unable to execute. Any previous assignment of the thread is nullified. Unix system call compatibility code may temporarily force threads to execute on the master processor.

This function returns KERN_SUCCESS if the assignment has been performed and KERN_ INVALID_ARGUMENT if thread is not a thread, or processor set is not a processor set on the same host as thread.

Function kern_return_t thread assign default (mach_port_t thread)

The function thread_assign_default is a variant of thread_assign that assigns the thread to the default processor set on that thread’s host. This variant exists because the control port for the default processor set is privileged and not ususally available to users.

This function returns KERN_SUCCESS if the assignment has been performed and KERN_ INVALID_ARGUMENT if thread is not a thread.

Function kern_return_t thread get assignment (mach_port_t thread,

    processor_set_name_t *processor set)

The function thread_get_assignment returns the name of the processor set to which the thread is currently assigned in processor set. This port can only be used to obtain information about the processor set. This function returns KERN_SUCCESS if the assignment has been performed, KERN_ INVALID_ADDRESS if processor set points to inaccessible memory, and KERN_INVALID_ ARGUMENT if thread is not a thread.

9.5 Processor Set Priority

Function kern_return_t processor set max priority

(mach_port_t processor set, int priority, boolean_t change threads) The function processor_set_max_priority is used to set the maximum priority for a processor set. The priority of a processor set is used only for newly created threads (thread’s maximum priority is set to processor set’s) and the assignment of threads to the set (thread’s maximum priority is reduced if it exceeds the set’s maximum priority, thread’s priority is similarly reduced). processor set max priority changes this priority. It also sets the maximum priority of all threads assigned to the processor set to this new priority if change threads is TRUE. If this maximum priority is less than the priorities of any of these threads, their priorities will also be set to this new value. This function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ ARGUMENT if processor set is not a processor set or priority is not a valid priority.

9.6 Processor Set Policy

Function kern_return_t processor set policy enable

(mach_port_t processor set, int policy)

Function kern_return_t processor set policy disable

(mach_port_t processor set, int policy, boolean_t change threads) Processor sets may restrict the scheduling policies to be used for threads assigned to them. These two calls provide the mechanism for designating permitted and forbidden policies. The current set of permitted policies can be obtained from processor_set_ info. Timesharing may not be forbidden by any processor set. This is a compromise to reduce the complexity of the assign operation; any thread whose policy is forbidden by the target processor set has its policy reset to timesharing. If the change threads argument to processor_set_policy_disable is true, threads currently assigned to this processor set and using the newly disabled policy will have their policy reset to timesharing. ‘mach/policy.h’ contains the allowed policies; it is included by ‘mach.h’. Not all policies (e.g. fixed priority) are supported by all systems. This function returns KERN_SUCCESS if the operation was completed successfully and KERN_INVALID_ARGUMENT if processor set is not a processor set or policy is not a valid policy, or an attempt was made to disable timesharing.

9.7 Processor Set Info

Function kern_return_t processor set info (mach_port_t processor set,

    int flavor, mach_port_t *host, processor_set_info_t processor set info,
    mach_msg_type_number_t *processor set infoCnt)

The function processor_set_info returns the selected information array for a processor set, as specified by flavor. host is set to the host on which the processor set resides. This is the non-privileged host port. processor set info is an array of integers that is supplied by the caller and returned filled with specified information. processor set infoCnt is supplied as the maximum number of integers in processor set info. On return, it contains the actual number of integers in processor set info. The maximum number of integers by any flavor is

PROCESSOR_SET_INFO_MAX.

The type of information returned is defined by flavor, which can be one of the following:

PROCESSOR_SET_BASIC_INFO

The function returns basic information about the processor set, as defined by processor_set_basic_info_t. This includes the number of tasks and threads assigned to the processor set. The number of integers

          returned is PROCESSOR_SET_BASIC_INFO_COUNT.
PROCESSOR_SET_SCHED_INFO

The function returns information about the schduling policy for the processor set as defined by processor_set_sched_info_t. The number of

          integers returned is PROCESSOR_SET_SCHED_INFO_COUNT.

Some machines may define additional (machine-dependent) flavors. The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if processor set is not a processor set or flavor is not recognized. The function returns MIG_ARRAY_TOO_LARGE if the returned info array is too large for processor set info. In this case, processor set info is filled as much as possible and processor set infoCnt is set to the number of elements that would have been returned if there were enough room.

Data type struct processor set basic info

This structure is returned in processor set info by the processor_set_info function and provides basic information about the processor set. You can cast a variable of type processor_set_info_t to a pointer of this type if you provided it as the processor set info parameter for the PROCESSOR_SET_BASIC_INFO flavor of processor_ set_info. It has the following members:

int processor_count

number of processors

int task_count

number of tasks

int thread_count

number of threads

int load_average

scaled load average

int mach_factor

scaled mach factor

Data type processor set basic info t

This is a pointer to a struct processor_set_basic_info.

Data type struct processor set sched info

This structure is returned in processor set info by the processor_set_info function and provides schedule information about the processor set. You can cast a variable of type processor_set_info_t to a pointer of this type if you provided it as the processor set info parameter for the PROCESSOR_SET_SCHED_INFO flavor of processor_set_info. It has the following members:

int policies

allowed policies

int max_priority

max priority for new threads

Data type processor set sched info t

This is a pointer to a struct processor_set_sched_info.

9.8 Processor

Function kern_return_t host processors (mach_port_t host priv,

    processor_array_t *processor list,
    mach_msg_type_number_t *processor count)

The function host_processors gets send rights to the processor port for each processor existing on host priv. This is the privileged port that allows its holder to control a processor. processor list is an array that is created as a result of this call. The caller may wish to vm_deallocate this array when the data is no longer needed. processor count is set to the number of processors in the processor list. This function returns KERN_SUCCESS if the call succeeded, KERN_INVALID_ARGUMENT if host priv is not a privileged host port, and KERN_INVALID_ADDRESS if processor count points to inaccessible memory.

Function kern_return_t processor start (mach_port_t processor) Function kern_return_t processor exit (mach_port_t processor) Function kern_return_t processor control (mach_port_t processor,

    processor_info_t *cmd, mach_msg_type_number_t count)

Some multiprocessors may allow privileged software to control processors. The

processor_start, processor_exit, and processor_control operations implement

this. The interpretation of the command in cmd is machine dependent. A newly started processor is assigned to the default processor set. An exited processor is removed from the processor set to which it was assigned and ceases to be active. count contains the length of the command cmd as a number of ints. Availability limited. All of these operations are machine-dependent. They may do nothing. The ability to restart an exited processor is also machine-dependent. This function returns KERN_SUCCESS if the operation was performed, KERN_FAILURE if the operation was not performed (a likely reason is that it is not supported on this processor), KERN_INVALID_ARGUMENT if processor is not a processor, and KERN_ INVALID_ADDRESS if cmd points to inaccessible memory.

Function kern_return_t processor assign (mach_port_t processor,

    mach_port_t processor set, boolean_t wait)

The function processor_assign assigns processor to the the set processor set. After the assignment is completed, the processor only executes threads that are assigned to that processor set. Any previous assignment of the processor is nullified. The master processor cannot be reassigned. All processors take clock interrupts at all times. The wait argument indicates whether the caller should wait for the assignment to be completed or should return immediately. Dedicated kernel threads are used to perform processor assignment, so setting wait to FALSE allows assignment requests to be queued and performed faster, especially if the kernel has more than one dedicated internal thread for processor assignment. Redirection of other device interrupts away from processors assigned to other than the default processor set is machine-dependent. Intermediaries that interpose on ports must be sure to interpose on both ports involved in this call if they interpose on either. This function returns KERN_SUCCESS if the assignment has been performed, KERN_ INVALID_ARGUMENT if processor is not a processor, or processor set is not a processor set on the same host as processor.

Function kern_return_t processor get assignment

    (mach_port_t processor, mach_port_t *assigned set)

The function processor_get_assignment obtains the current assignment of a processor. The name port of the processor set is returned in assigned set.

Function kern_return_t processor info (mach_port_t processor, int flavor,

    mach_port_t *host, processor_info_t processor info,
    mach_msg_type_number_t *processor infoCnt)

The function processor_info returns the selected information array for a processor, as specified by flavor. host is set to the host on which the processor set resides. This is the non-privileged host port. processor info is an array of integers that is supplied by the caller and returned filled with specified information. processor infoCnt is supplied as the maximum number of integers in processor info. On return, it contains the actual number of integers in processor info. The maximum number of integers by any flavor is PROCESSOR_INFO_

MAX.

The type of information returned is defined by flavor, which can be one of the following:

PROCESSOR_BASIC_INFO

The function returns basic information about the processor, as defined by processor_basic_info_t. This includes the slot number of the processor. The number of integers returned is PROCESSOR_BASIC_INFO_COUNT.

Machines which require more configuration information beyond the slot number are expected to define additional (machine-dependent) flavors. The function returns KERN_SUCCESS if the call succeeded and KERN_INVALID_ARGUMENT if processor is not a processor or flavor is not recognized. The function returns MIG_ ARRAY_TOO_LARGE if the returned info array is too large for processor info. In this case, processor info is filled as much as possible and processor infoCnt is set to the number of elements that would have been returned if there were enough room.

Data type struct processor basic info

This structure is returned in processor info by the processor_info function and provides basic information about the processor. You can cast a variable of type processor_info_t to a pointer of this type if you provided it as the processor info parameter for the PROCESSOR_BASIC_INFO flavor of processor_info. It has the following members:

cpu_type_t cpu_type

cpu type

cpu_subtype_t cpu_subtype

cpu subtype

boolean_t running

is processor running?

int slot_num

slot number

boolean_t is_master

is this the master processor

Data type processor basic info t

This is a pointer to a struct processor_basic_info.

10 Device Interface

The GNU Mach microkernel provides a simple device interface that allows the user space programs to access the underlying hardware devices. Each device has a unique name, which is a string up to 127 characters long. To open a device, the device master port has to be supplied. The device master port is only available through the bootstrap port. Anyone who has control over the device master port can use all hardware devices.

For each device opened, a port is created that represants the device. Operations on the device are implemented as remote procedure calls to the device port. Each device provides a sequence of records. The length of a record is specific to the device. Data can be transferred “out-of-band” or “inband”.

10.1 Device Reply Server

Beside the usual synchronous interface, an asynchronous interface is provided. For this, the caller has to receive and handle the reply messages seperately from the function call.

Function boolean_t device reply server (msg_header_t *in msg,

    msg_header_t *out msg)

The function device_reply_server is produced by the remote procedure call generator to to handle a received message. This function does all necessary argument handling, and actually calls one of the following functions: ds_device_open_reply,

ds_device_read_reply, ds_device_read_reply_inband, ds_device_write_reply
and ds_device_write_reply_inband.

The in msg argument is the message that has been received from the kernel. The out msg is a reply message, but this is not used for this server. The function returns TRUE to indicate that the message in question was applicable to this interface, and that the appropriate routine was called to interpret the message. It returns FALSE to indicate that the message did not apply to this interface, and that no other action was taken.

10.2 Device Open

Function kern_return_t device open (mach_port_t master port,

    dev_mode_t mode, dev_name_t name, mach_port_t *device)

The function device_open opens the device name and returns a port to it in device. The open count for the device is incremented by one. If the open count was 0, the open handler for the device is invoked. master port must hold the master device port. name specifies the device to open, and is a string up to 128 characters long. mode is the open mode. It is a bitwise-or of the following constants:

D_READ Request read access for the device.

D_WRITE Request write access for the device.

D_NODELAY

Do not delay an open.

The function returns D_SUCCESS if the device was successfully opened, D_INVALID_ OPERATION if master port is not the master device port, D_WOULD_BLOCK is the device is busy and D_NOWAIT was specified in mode, D_ALREADY_OPEN if the device is already open in an incompatible mode and D_NO_SUCH_DEVICE if name does not denote a know device.

Function kern_return_t device open request (mach_port_t master port,

    mach_port_t reply port, dev_mode_t mode, dev_name_t name)

Function kern_return_t ds device open reply (mach_port_t reply port,

    kern_return_t return, mach_port_t *device)

This is the asynchronous form of the device_open function. device_open_request performs the open request. The meaning for the parameters is as in device_open. Additionally, the caller has to supply a reply port to which the ds_device_open_ reply message is sent by the kernel when the open has been performed. The return value of the open operation is stored in return code. As neither function receives a reply message, only message transmission errors apply. If no error occurs, KERN_SUCCESS is returned.

10.3 Device Close

Function kern_return_t device close (mach_port_t device)

The function device_close decrements the open count of the device by one. If the open count drops to zero, the close handler for the device is called. The device to close is specified by its port device. The function returns D_SUCCESS if the device was successfully closed and D_NO_SUCH_ DEVICE if device does not denote a device port.

10.4 Device Read

Function kern_return_t device read (mach_port_t device,
    dev_mode_t mode, recnum_t recnum, int bytes wanted,
    io_buf_ptr_t *data, mach_msg_type_number_t *data count)

The function device_read reads bytes wanted bytes from device, and stores them in a buffer allocated with vm_allocate, which address is returned in data. The caller must deallocated it if it is no longer needed. The number of bytes actually returned is stored in data count. If mode is D_NOWAIT, the operation does not block. Otherwise mode should be 0. recnum is the record number to be read, its meaning is device specific. The function returns D_SUCCESS if some data was successfully read, D_WOULD_BLOCK if no data is currently available and D_NOWAIT is specified, and D_NO_SUCH_DEVICE if device does not denote a device port.

Function kern_return_t device read inband (mach_port_t device,

    dev_mode_t mode, recnum_t recnum, int bytes wanted,
    io_buf_ptr_inband_t *data, mach_msg_type_number_t *data count)

The device_read_inband function works as the device_read function, except that the data is returned “inband” in the reply IPC message.

Function kern_return_t device read request (mach_port_t device,

    mach_port_t reply port, dev_mode_t mode, recnum_t recnum,

int bytes wanted)

Function kern_return_t ds device read reply (mach_port_t reply port,

    kern_return_t return code, io_buf_ptr_t data,
    mach_msg_type_number_t data count)

This is the asynchronous form of the device_read function. device_read_request performs the read request. The meaning for the parameters is as in device_read. Additionally, the caller has to supply a reply port to which the ds_device_read_ reply message is sent by the kernel when the read has been performed. The return value of the read operation is stored in return code.

As neither function receives a reply message, only message transmission errors apply. If no error occurs, KERN_SUCCESS is returned.

Function kern_return_t device read request inband

    (mach_port_t device, mach_port_t reply port, dev_mode_t mode,

recnum_t recnum, int bytes wanted)

Function kern_return_t ds device read reply inband

    (mach_port_t reply port, kern_return_t return code, io_buf_ptr_t data,
    mach_msg_type_number_t data count)
The device_read_request_inband and ds_device_read_reply_inband functions
work as the device_read_request and ds_device_read_reply functions, except

that the data is returned “inband” in the reply IPC message.

10.5 Device Write

Function kern_return_t device write (mach_port_t device,

    dev_mode_t mode, recnum_t recnum, io_buf_ptr_t data,
    mach_msg_type_number_t data count, int *bytes written)

The function device_write writes data count bytes from the buffer data to device. The number of bytes actually written is returned in bytes written.

If mode is D_NOWAIT, the function returns without waiting for I/O completion. Otherwise mode should be 0. recnum is the record number to be written, its meaning is device specific.

The function returns D_SUCCESS if some data was successfully written and D_NO_ SUCH_DEVICE if device does not denote a device port or the device is dead or not completely open.

Function kern_return_t device write inband (mach_port_t device,

    dev_mode_t mode, recnum_t recnum, int bytes wanted,
    io_buf_ptr_inband_t *data, mach_msg_type_number_t *data count)

The device_write_inband function works as the device_write function, except that the data is sent “inband” in the request IPC message.

Function kern_return_t device write request (mach_port_t device,

    mach_port_t reply port, dev_mode_t mode, recnum_t recnum,
    io_buf_ptr_t data, mach_msg_type_number_t data count)

Function kern_return_t ds device write reply (mach_port_t reply port,

kern_return_t return code, int bytes written) This is the asynchronous form of the device_write function. device_write_request performs the write request. The meaning for the parameters is as in device_write. Additionally, the caller has to supply a reply port to which the ds_device_write_ reply message is sent by the kernel when the write has been performed. The return value of the write operation is stored in return code.

As neither function receives a reply message, only message transmission errors apply. If no error occurs, KERN_SUCCESS is returned.

Function kern_return_t device write request inband

    (mach_port_t device, mach_port_t reply port, dev_mode_t mode,
    recnum_t recnum, io_buf_ptr_t data,
    mach_msg_type_number_t data count)

Function kern_return_t ds device write reply inband

(mach_port_t reply port, kern_return_t return code, int bytes written)

The device_write_request_inband and ds_device_write_reply_inband func-
tions work as the device_write_request and ds_device_write_reply functions,

except that the data is sent “inband” in the request IPC message.

10.6 Device Map

Function kern_return_t device map (mach_port_t device, vm_prot_t prot,
    vm_offset_t offset, vm_size_t size, mach_port_t *pager, int unmap)

The function device_map creates a new memory manager for device and returns a port to it in pager. The memory manager is usable as a memory object in a vm_map call. The call is device dependant.

The protection for the memory object is specified by prot. The memory object starts at offset within the device and extends size bytes. unmap is currently unused.

The function returns D_SUCCESS if some data was successfully written and D_NO_ SUCH_DEVICE if device does not denote a device port or the device is dead or not completely open.

10.7 Device Status

Function kern_return_t device set status (mach_port_t device,

    dev_flavor_t flavor, dev_status_t status,
    mach_msg_type_number_t status count)

The function device_set_status sets the status of a device. The possible values for flavor and their interpretation is device specific. The function returns D_SUCCESS if some data was successfully written and D_NO_ SUCH_DEVICE if device does not denote a device port or the device is dead or not completely open.

Function kern_return_t device get status (mach_port_t device,

    dev_flavor_t flavor, dev_status_t status,
    mach_msg_type_number_t *status count)

The function device_get_status gets the status of a device. The possible values for flavor and their interpretation is device specific. The function returns D_SUCCESS if some data was successfully written and D_NO_ SUCH_DEVICE if device does not denote a device port or the device is dead or not completely open.

10.8 Device Filter

Function kern_return_t device set filter (mach_port_t device,

    mach_port_t receive port, mach_msg_type_name_t receive port type,
    int priority, filter_array_t filter, mach_msg_type_number_t filter count)

The function device_set_filter makes it possible to filter out selected data arriving at the device and forward it to a port. filter is a list of filter commands, which are applied to incoming data to determine if the data should be sent to receive port. The IPC type of the send right is specified by receive port right, it is either MACH_MSG_

TYPE_MAKE_SEND or MACH_MSG_TYPE_MOVE_SEND. The priority value is used to order

multiple filters. There can be up to NET_MAX_FILTER commands in filter. The actual number of commands is passed in filter count. For the purpose of the filter test, an internal stack is provided. After all commands have been processed, the value on the top of the stack determines if the data is forwarded or the next filter is tried. Each word of the command list specifies a data (push) operation (high order NETF NBPO bits) as well as a binary operator (low order NETF NBPA bits). The value to be pushed onto the stack is chosen as follows.

ETF_PUSHLIT

Use the next short word of the filter as the value.

NETF_PUSHZERO

Use 0 as the value.

NETF_PUSHWORD+N

Use short word N of the “data” portion of the message as the value.

NETF_PUSHHDR+N

Use short word N of the “header” portion of the message as the value.

NETF_PUSHIND+N

Pops the top long word from the stack and then uses short word N of the “data” portion of the message as the value.

NETF_PUSHHDRIND+N

Pops the top long word from the stack and then uses short word N of the “header” portion of the message as the value.

NETF_PUSHSTK+N

Use long word N of the stack (where the top of stack is long word 0) as the value.

NETF_NOPUSH

Don’t push a value.

The unsigned value so chosen is promoted to a long word before being pushed. Once a value is pushed (except for the case of NETF_NOPUSH), the top two long words of the stack are popped and a binary operator applied to them (with the old top of stack as the second operand). The result of the operator is pushed on the stack. These operators are: NETF_NOP Don’t pop off any values and do no operation. NETF_EQ Perform an equal comparison. NETF_LT Perform a less than comparison. NETF_LE Perform a less than or equal comparison. NETF_GT Perform a greater than comparison. NETF_GE Perform a greater than or equal comparison. NETF_AND Perform a bitise boolean AND operation. NETF_OR Perform a bitise boolean inclusive OR operation. NETF_XOR Perform a bitise boolean exclusive OR operation. NETF_NEQ Perform a not equal comparison. NETF_LSH Perform a left shift operation. NETF_RSH Perform a right shift operation. NETF_ADD Perform an addition. NETF_SUB Perform a subtraction. NETF_COR Perform an equal comparison. If the comparison is TRUE, terminate the

filter list. Otherwise, pop the result of the comparison off the stack.

NETF_CAND

Perform an equal comparison. If the comparison is FALSE, terminate the filter list. Otherwise, pop the result of the comparison off the stack.

NETF_CNOR

Perform a not equal comparison. If the comparison is FALSE, terminate the filter list. Otherwise, pop the result of the comparison off the stack.

NETF_CNAND

Perform a not equal comparison. If the comparison is TRUE, terminate the filter list. Otherwise, pop the result of the comparison off the stack. The scan of the filter list terminates when the filter list is emptied, or a NETF_C... operation terminates the list. At this time, if the final value of the top of the stack is TRUE, then the message is accepted for the filter.

The function returns D_SUCCESS if some data was successfully written, D_INVALID_ OPERATION if receive port is not a valid send right, and D_NO_SUCH_DEVICE if device does not denote a device port or the device is dead or not completely open.

102 The GNU Mach Reference Manual

11 Kernel Debugger

The GNU Mach kernel debugger ddb is a powerful built-in debugger with a gdb like syntax. It is enabled at compile time using the ‘--enable-kdb’ option. Whenever you want to enter the debugger while running the kernel, you can press the key combination ⟨\underline{\overline{Ctrl-Alt-D}}^{⟩}.

11.1 Operation

The current location is called dot. The dot is displayed with a hexadecimal format at a prompt. Examine and write commands update dot to the address of the last line examined or the last location modified, and set next to the address of the next location to be examined or changed. Other commands don’t change dot, and set next to be the same as dot.

The general command syntax is:

command[/modifier] address [,count] !! repeats the previous command, and a blank line repeats from the address next with

count 1 and no modifiers. Specifying address sets dot to the address. Omitting address uses dot. A missing count is taken to be 1 for printing commands or infinity for stack traces.

Current ddb is enhanced to support multi-thread debugging. A break point can be set only for a specific thread, and the address space or registers of non current thread can be examined or modified if supported by machine dependent routines. For example,

     break/t mach_msg_trap $task11.0

sets a break point at mach_msg_trap for the first thread of task 11 listed by a show all threads command.

In the above example, $task11.0 is translated to the corresponding thread structure’s address by variable translation mechanism described later. If a default target thread is set in a variable $thread, the $task11.0 can be omitted. In general, if t is specified in a modifier of a command line, a specified thread or a default target thread is used as a target thread instead of the current one. The t modifier in a command line is not valid in evaluating expressions in a command line. If you want to get a value indirectly from a specific thread’s address space or access to its registers within an expression, you have to specify a default target thread in advance, and to use :t modifier immediately after the indirect access or the register reference like as follows:

     set $thread $task11.0
     print $eax:t *(0x100):tuh

No sign extension and indirection size(long, half word, byte) can be specified with u, l, h and b respectively for the indirect access.

Note: Support of non current space/register access and user space break point depend on the machines. If not supported, attempts of such operation may provide incorrect information or may cause strange behavior. Even if supported, the user space access is limited to the pages resident in the main memory at that time. If a target page is not in the main memory, an error will be reported.

ddb has a feature like a command more for the output. If an output line exceeds the number set in the $lines variable, it displays ‘--db_more--’ and waits for a response. The valid responses for it are:

⟨\underline{\overline{SPC}}^{⟩} one more page ⟨\underline{\overline{RET}}^{⟩} one more line q abort the current command, and return to the command input mode

11.2 Commands

examine(x) [/modifier] addr[,count] [ thread ] Display the addressed locations according to the formats in the modifier. Multiple modifier formats display multiple locations. If no format is specified, the last formats specified for this command is used. Address space other than that of the current thread can be specified with t option in the modifier and thread parameter. The format characters are b look at by bytes(8 bits) h look at by half words(16 bits) l look at by long words(32 bits) a print the location being displayed , skip one unit producing no output A print the location with a line number if possible x display in unsigned hex z display in signed hex o display in unsigned octal d display in signed decimal u display in unsigned decimal r display in current radix, signed c display low 8 bits as a character. Non-printing characters are dis-

played as an octal escape code (e.g. ’\000’).

s display the null-terminated string at the location. Non-printing

characters are displayed as octal escapes.

m display in unsigned hex with character dump at the end of each line. The location is also displayed in hex at the beginning of each line.

i display as an instruction I display as an instruction with possible alternate formats depending

on the machine:

vax don’t assume that each external label is a procedure

entry mask

i386 don’t round to the next long word boundary

mips print register contents

xf Examine forward. It executes an examine command with the last specified parameters to it except that the next address displayed by it is used as the start address.

xb Examine backward. It executes an examine command with the last specified parameters to it except that the last start address subtracted by the size displayed by it is used as the start address.

print[/axzodurc] addr1 [ addr2 ... ]

Print addr’s according to the modifier character. Valid formats are: a x z o d u r c. If no modifier is specified, the last one specified to it is used. addr can be a string, and it is printed as it is. For example,

     print/x "eax = " $eax "\necx = " $ecx "\n"

will print like

     eax = xxxxxx
     ecx = yyyyyy

write[/bhlt] addr [ thread ] expr1 [ expr2 ... ] Write the expressions at succeeding locations. The write unit size can be specified in the modifier with a letter b (byte), h (half word) or l(long word) respectively. If omitted, long word is assumed. Target address space can also be specified with t option in the modifier and thread parameter. Warning: since there is no delimiter between expressions, strange things may happen. It’s best to enclose each expression in parentheses.

set $variable [=] expr Set the named variable or register with the value of expr. Valid variable names are described below.

break[/tuTU] addr[,count] [ thread1 ... ]

Set a break point at addr. If count is supplied, continues (count-1) times before stopping at the break point. If the break point is set, a break point number is printed with ‘#’. This number can be used in deleting the break point or adding conditions to it.

t Set a break point only for a specific thread. The thread is specified by thread parameter, or default one is used if the parameter is omitted.

u Set a break point in user space address. It may be combined with t or T option to specify the non-current target user space. Without u option, the address is considered in the kernel space, and wrong space address is rejected with an error message. This option can be used only if it is supported by machine dependent routines.

T Set a break point only for threads in a specific task. It is like t option except that the break point is valid for all threads which belong to the same task as the specified target thread.

U Set a break point in shared user space address. It is like u option, except that the break point is valid for all threads which share the same address space even if t option is specified. t option is used only to specify the target shared space. Without t option, u and U have the same meanings. U is useful for setting a user space break point in non-current address space with t option such as in an emulation library space. This option can be used only if it is supported by machine dependent routines.

Warning: if a user text is shadowed by a normal user space debugger, user space break points may not work correctly. Setting a break point at the low-level code paths may also cause strange behavior.

delete[/tuTU] addr|#number [ thread1 ... ]

Delete the break point. The target break point can be specified by a break point number with #, or by addr like specified in break command.

cond #number [ condition commands ] Set or delete a condition for the break point specified by the number. If the condition and commands are null, the condition is deleted. Otherwise the condition is set for it. When the break point is hit, the condition is evaluated. The commands will be executed if the condition is true and the break point count set by a break point command becomes zero. commands is a list of commands separated by semicolons. Each command in the list is executed in that order, but if a continue command is executed, the command execution stops there, and the stopped thread resumes execution. If the command execution reaches the end of the list, and it enters into a command input mode. For example,

     set $work0 0
     break/Tu xxx_start $task7.0
     cond #1  (1)  set $work0 1; set $work1 0; cont
     break/T  vm_fault $task7.0
     cond #2  ($work0) set $work1 ($work1+1); cont
     break/Tu xxx_end $task7.0
     cond #3  ($work0) print $work1 " faults\n"; set $work0 0
     cont

will print page fault counts from xxx_start to xxx_end in task7.

step[/p] [,count]

Single step count times. If p option is specified, print each instruction at each step. Otherwise, only print the last instruction. Warning: depending on machine type, it may not be possible to single-step through some low-level code paths or user space code. On machines with software-emulated single-stepping (e.g., pmax), stepping through code executed by interrupt handlers will probably do the wrong thing.

continue[/c]

Continue execution until a breakpoint or watchpoint. If /c, count instructions while executing. Some machines (e.g., pmax) also count loads and stores. Warning: when counting, the debugger is really silently single-stepping. This means that single-stepping on low-level code may cause strange behavior. until Stop at the next call or return instruction.

next[/p] Stop at the matching return instruction. If p option is specified, print the call nesting depth and the cumulative instruction count at each call or return. Otherwise, only print when the matching return is hit.

match[/p]

A synonym for next.

trace[/tu] [ frame addr|thread ][,count] Stack trace. u option traces user space; if omitted, only traces kernel space. If t option is specified, it shows the stack trace of the specified thread or a default target thread. Otherwise, it shows the stack trace of the current thread from the frame address specified by a parameter or from the current frame. count is the number of frames to be traced. If the count is omitted, all frames are printed. Warning: If the target thread’s stack is not in the main memory at that time, the stack trace will fail. User space stack trace is valid only if the machine dependent code supports it.

search[/bhl] addr value [mask] [,count] Search memory for a value. This command might fail in interesting ways if it doesn’t find the searched-for value. This is because ddb doesn’t always recover from touching bad memory. The optional count argument limits the search.

macro name commands Define a debugger macro as name. commands is a list of commands to be associated with the macro. In the expressions of the command list, a variable $argxx can be used to get a parameter passed to the macro. When a macro is called, each argument is evaluated as an expression, and the value is assigned to each parameter, $arg1, $arg2, … respectively. 10 $arg variables are reserved to each level of macros, and they can be used as local variables. The nesting of macro can be allowed up to 5 levels. For example,

     macro xinit set $work0 $arg1
     macro xlist examine/m $work0,4; set $work0 *($work0)
     xinit *(xxx_list)
     xlist
     ....

will print the contents of a list starting from xxx_list by each xlist command.

dmacro name

Delete the macro named name.

show all threads[/ul]

Display all tasks and threads information. This version of ddb prints more information than previous one. It shows UNIX process information like ps for each task. The UNIX process information may not be shown if it is not supported in the machine, or the bottom of the stack of the target task is not in the main memory at that time. It also shows task and thread identification numbers. These numbers can be used to specify a task or a thread symbolically in various commands. The numbers are valid only in the same debugger session. If the execution is resumed again, the numbers may change. The current thread can be distinguished from others by a # after the thread id instead of :. Without l option, it only shows thread id, thread structure address and the status for each thread. The status consists of 5 letters, R(run), W(wait), S(sus pended), O(swapped out) and N(interruptible), and if corresponding status bit is off, . is printed instead. If l option is specified, more detail information is printed for each thread.

show task [ addr ]

Display the information of a task specified by addr. If addr is omitted, current task information is displayed.

show thread [ addr ]

Display the information of a thread specified by addr. If addr is omitted, current thread information is displayed.

show registers[/tu [ thread ]]

Display the register set. Target thread can be specified with t option and thread parameter. If u option is specified, it displays user registers instead of kernel or currently saved one. Warning: The support of t and u option depends on the machine. If not supported, incorrect information will be displayed.

show map addr

Prints the vm_map at addr.

show object addr

Prints the vm_object at addr.

show page addr

Prints the vm_page structure at addr.

show port addr

Prints the ipc_port structure at addr.

show ipc_port[/t [ thread ]]

Prints all ipc_port structure’s addresses the target thread has. The target thread is a current thread or that specified by a parameter.

show macro [ name ]

Show the definitions of macros. If name is specified, only the definition of it is displayed. Otherwise, definitions of all macros are displayed.

show watches

Displays all watchpoints.

watch[/T] addr,size [ task ]

Set a watchpoint for a region. Execution stops when an attempt to modify the region occurs. The size argument defaults to 4. Without T option, addr is assumed to be a kernel address. If you want to set a watch point in user space, specify T and task parameter where the address belongs to. If the task parameter is omitted, a task of the default target thread or a current task is assumed. If you specify a wrong space address, the request is rejected with an error message. Warning: Attempts to watch wired kernel memory may cause unrecoverable error in some systems such as i386. Watchpoints on user addresses work best.

11.3 Variables

The debugger accesses registers and variables as $name. Register names are as in the show registers command. Some variables are suffixed with numbers, and may have some modifier following a colon immediately after the variable name. For example, register variables can have u and t modifier to indicate user register and that of a default target thread instead of that of the current thread (e.g. $eax:tu).

Built-in variables currently supported are:

taskxx[.yy]

Task or thread structure address. xx and yy are task and thread identification numbers printed by a show all threads command respectively. This variable is read only.

thread The default target thread. The value is used when t option is specified without explicit thread structure address parameter in command lines or expression evaluation.

radix Input and output radix

maxoff Addresses are printed as symbol+offset unless offset is greater than maxoff.

maxwidth The width of the displayed line.

lines The number of lines. It is used by more feature.

tabstops Tab stop width.

argxx Parameters passed to a macro. xx can be 1 to 10.

workxx Work variable. xx can be 0 to 31.

11.4 Expressions

Almost all expression operators in C are supported except ~, ^, and unary &. Special rules in ddb are:

identifier name of a symbol. It is translated to the address(or value) of it. . and : can be used in the identifier. If supported by an object format dependent routine, [file name:]func[:line number] [file name:]variable, and file name[:line number] can be accepted as a symbol. The symbol may be prefixed with symbol table name:: like emulator::mach_msg_trap to specify other than kernel symbols.

number radix is determined by the first two letters:

0x hex

0o octal 0t decimal otherwise, follow current radix.

. dot + next .. address of the start of the last line examined. Unlike dot or next, this is only

changed by examine or write command. last address explicitly specified.

$variable register name or variable. It is translated to the value of it. It may be followed

by a : and modifiers as described above.

a multiple of right hand side. *expr indirection. It may be followed by a : and modifiers as described above.

Appendix A GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc. 59 Temple Place – Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

A.0.1 Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

A.0.2 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

1. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”.

Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

2. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program.

You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

3. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions:

a. You must cause the modified files to carry prominent notices stating that you changed the files and the date of any change.

b. You must cause any work that you distribute or publish, that in whole or in part contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

4. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code,

which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

5. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

6. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

7. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

8. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

9. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

10. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

11. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

12. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLI-CABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPY-RIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MER-CHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

13. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

Copyright (C) 19yy name of author
This program is free software; you can redistribute it and/or
modify it under the terms of the GNU General Public License
as published by the Free Software Foundation; either version 2
of the License, or (at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License along
with this program; if not, write to the Free Software Foundation, Inc.,
59 Temple Place, Suite 330, Boston, MA 02111-1307, USA.

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright
interest in the program `Gnomovision'
(which makes passes at compilers) written
by James Hacker.

signature of Ty Coon, 1 April 1989

Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

118 The GNU Mach Reference Manual

Appendix B Documentation License

This manual is copyrighted and licensed under the GNU Free Documentation license. Parts of this manual are derived from the Mach manual packages originally provided by

Carnegie Mellon University.

B.1 GNU Free Documentation License

Version 1.1, March 2000

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other written document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (For example, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, whose contents can be viewed and edited directly and straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup has been designed to thwart or discourage subsequent modification by readers is not Transparent. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, sgml or xml using a publicly available dtd, and standard-conforming simple html designed for human modification. Opaque formats include PostScript, pdf, proprietary formats that can be read and edited only by proprietary word processors, sgml or xml for which the dtd and/or processing tools are not generally available, and the machine-generated html produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies of the Document numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a publicly-accessible computer-network location containing a complete Transparent copy of the Document, free of added material, which the general network-using public has access to download anonymously at no charge using public-standard network protocols. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has less than five).

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License.

I. Preserve the section entitled “History”, and its title, and add to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. In any section entitled “Acknowledgments” or “Dedications”, preserve the section’s title, and preserve in the section all the substance and tone of each of the contributor acknowledgments and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section as “Endorsements” or to conflict in title with

any Invariant Section.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections entitled “History” in the various original documents, forming one section entitled “History”; likewise combine any sections entitled “Acknowledgments”, and any sections entitled “Dedications”. You must delete all sections entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, does not as a whole count as a Modified Version of the Document, provided no compilation copyright is claimed for the compilation. Such a compilation is called an “aggregate”, and this License does not apply to the other self-contained works thus compiled with the Document, on account of their being thus compiled, if they are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one quarter of the entire aggregate, the Document’s Cover Texts may be placed on covers that surround only the Document within the aggregate. Otherwise they must appear on covers around the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License provided that you also include the original English version of this License. In case of a disagreement between the translation and the original English version of this License, the original English version will prevail.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

B.1.0.1 ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

  Copyright (C) year your name.
  Permission is granted to copy, distribute and/or modify this document
  under the terms of the GNU Free Documentation License, Version 1.1
  or any later version published by the Free Software Foundation;
  with the Invariant Sections being list their titles, with the
  Front-Cover Texts being list, and with the Back-Cover Texts being list.
  A copy of the license is included in the section entitled ``GNU
  Free Documentation License''.

If you have no Invariant Sections, write “with no Invariant Sections” instead of saying which ones are invariant. If you have no Front-Cover Texts, write “no Front-Cover Texts” instead of “Front-Cover Texts being list”; likewise for Back-Cover Texts.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

B.2 CMU License

Mach Operating System

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991,1990,1989 Carnegie Mellon University All Rights Reserved.

Permission to use, copy, modify and distribute this software and its documentation is hereby granted, provided that both the copyright notice and this permission notice appear in all copies of the software, derivative works or modified versions, and any portions thereof, and that both notices appear in supporting documentation. carnegie mellon allows free use of this software in its “as is” condition. carnegie mellon disclaims any liability of any kind for any damages whatsoever resulting from the use of this software. Carnegie Mellon requests users of this software to return to

Software Distribution Coordinator School of Computer Science Carnegie Mellon University Pittsburgh PA 15213-3890

or Software.Distribution@CS.CMU.EDU any improvements or extensions that they make and grant Carnegie Mellon the rights to redistribute these changes.

126 The GNU Mach Reference Manual Concept Index 127

Concept Index

F

FDL, GNU Free Documentation License . . . . . . 119

G

GPL, GNU General Public License. . . . . . . . . . . 111 GRand Unified Bootloader . . . . . . . . . . . . . . . . . . . 11 GRUB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

S

serverboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

128 The GNU Mach Reference Manual Function and Data Index 129

Function and Data Index

C

catch_exception_raise . . . . . . . . . . . . . . . . . . . . . 70

D

device_close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 device_get_status . . . . . . . . . . . . . . . . . . . . . . . . . . 99 device_map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 device_open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 device_open_request. . . . . . . . . . . . . . . . . . . . . . . . 96 device_read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 device_read_inband . . . . . . . . . . . . . . . . . . . . . . . . . 97 device_read_request. . . . . . . . . . . . . . . . . . . . . . . . 97

device_read_request_inband . . . . . . . . . . . . . . . . 97

device_reply_server. . . . . . . . . . . . . . . . . . . . . . . . 95 device_set_filter . . . . . . . . . . . . . . . . . . . . . . . . . . 99 device_set_status . . . . . . . . . . . . . . . . . . . . . . . . . . 99 device_write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 device_write_inband. . . . . . . . . . . . . . . . . . . . . . . . 98 device_write_request. . . . . . . . . . . . . . . . . . . . . . . 98

device_write_request_inband . . . . . . . . . . . . . . . 98

ds_device_open_reply. . . . . . . . . . . . . . . . . . . . . . . 96 ds_device_read_reply. . . . . . . . . . . . . . . . . . . . . . . 97

ds_device_read_reply_inband . . . . . . . . . . . . . . . 97

ds_device_write_reply . . . . . . . . . . . . . . . . . . . . . 98

ds_device_write_reply_inband . . . . . . . . . . . . . . 98

E

evc_wait. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 exception_raise . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

H

host_adjust_time . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 host_basic_info_t . . . . . . . . . . . . . . . . . . . . . . . . . . 82 host_get_boot_info . . . . . . . . . . . . . . . . . . . . . . . . . 83 host_get_time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 host_info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 host_kernel_version. . . . . . . . . . . . . . . . . . . . . . . . 83

host_processor_set_priv . . . . . . . . . . . . . . . . . . . 87

host_processor_sets. . . . . . . . . . . . . . . . . . . . . . . . 87 host_processors . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 host_reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 host_sched_info_t . . . . . . . . . . . . . . . . . . . . . . . . . . 83 host_set_time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

M

mach_host_self . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 mach_msg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 mach_msg_bits_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 mach_msg_header_t . . . . . . . . . . . . . . . . . . . . . . . . . . 15 mach_msg_id_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 mach_msg_size_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 mach_msg_type_long_t. . . . . . . . . . . . . . . . . . . . . . . 19 mach_msg_type_name_t. . . . . . . . . . . . . . . . . . . . . . . 17

mach_msg_type_number_t . . . . . . . . . . . . . . . . . . . . 17
MACH_MSG_TYPE_PORT_ANY . . . . . . . . . . . . . . . . . . . . 19
MACH_MSG_TYPE_PORT_ANY_RIGHT . . . . . . . . . . . . . . 19
MACH_MSG_TYPE_PORT_ANY_SEND . . . . . . . . . . . . . . . 19

mach_msg_type_size_t. . . . . . . . . . . . . . . . . . . . . . . 17 mach_msg_type_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 MACH_MSGH_BITS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 MACH_MSGH_BITS_LOCAL. . . . . . . . . . . . . . . . . . . . . . . 17 MACH_MSGH_BITS_OTHER. . . . . . . . . . . . . . . . . . . . . . . 17 MACH_MSGH_BITS_PORTS. . . . . . . . . . . . . . . . . . . . . . . 17 MACH_MSGH_BITS_REMOTE . . . . . . . . . . . . . . . . . . . . . 17 mach_port_allocate . . . . . . . . . . . . . . . . . . . . . . . . . 29

mach_port_allocate_name . . . . . . . . . . . . . . . . . . . 30

mach_port_deallocate. . . . . . . . . . . . . . . . . . . . . . . 31 mach_port_destroy . . . . . . . . . . . . . . . . . . . . . . . . . . 31

mach_port_extract_right . . . . . . . . . . . . . . . . . . . 35
mach_port_get_receive_status . . . . . . . . . . . . . . 37

mach_port_get_refs . . . . . . . . . . . . . . . . . . . . . . . . . 33

mach_port_get_set_status . . . . . . . . . . . . . . . . . . 38
mach_port_insert_right . . . . . . . . . . . . . . . . . . . . 35

mach_port_mod_refs . . . . . . . . . . . . . . . . . . . . . . . . . 34 mach_port_move_member . . . . . . . . . . . . . . . . . . . . . 38 mach_port_mscount_t. . . . . . . . . . . . . . . . . . . . . . . . 36 mach_port_msgcount_t. . . . . . . . . . . . . . . . . . . . . . . 36 mach_port_names . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mach_port_rename . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

mach_port_request_notification . . . . . . . . . . . . 39

mach_port_rights_t . . . . . . . . . . . . . . . . . . . . . . . . . 36 mach_port_seqno_t . . . . . . . . . . . . . . . . . . . . . . . . . . 36 mach_port_set_mscount . . . . . . . . . . . . . . . . . . . . . 37 mach_port_set_qlimit. . . . . . . . . . . . . . . . . . . . . . . 37 mach_port_set_seqno. . . . . . . . . . . . . . . . . . . . . . . . 38 mach_port_status_t . . . . . . . . . . . . . . . . . . . . . . . . . 36 mach_port_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 mach_port_type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mach_reply_port . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 mach_task_self . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mach_thread_self . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 mapped_time_value_t. . . . . . . . . . . . . . . . . . . . . . . . 84

130 The GNU Mach Reference Manual

memory_object_change_attributes . . . . . . . . . . . 58
memory_object_change_completed . . . . . . . . . . . . 59

memory_object_copy . . . . . . . . . . . . . . . . . . . . . . . . . 55 memory_object_create. . . . . . . . . . . . . . . . . . . . . . . 59

memory_object_data_error . . . . . . . . . . . . . . . . . . 54
memory_object_data_initialize . . . . . . . . . . . . . 60
memory_object_data_provided . . . . . . . . . . . . . . . 56
memory_object_data_request . . . . . . . . . . . . . . . . 52
memory_object_data_return . . . . . . . . . . . . . . . . . 52
memory_object_data_supply . . . . . . . . . . . . . . . . . 53
memory_object_data_unavailable . . . . . . . . . . . . 54
memory_object_data_unlock . . . . . . . . . . . . . . . . . 57
memory_object_default_server . . . . . . . . . . . . . . 49

memory_object_destroy . . . . . . . . . . . . . . . . . . . . . 51

memory_object_get_attributes . . . . . . . . . . . . . . 58

memory_object_init . . . . . . . . . . . . . . . . . . . . . . . . . 49

memory_object_lock_completed . . . . . . . . . . . . . . 57
memory_object_lock_request . . . . . . . . . . . . . . . . 56

memory_object_ready. . . . . . . . . . . . . . . . . . . . . . . . 50 memory_object_server. . . . . . . . . . . . . . . . . . . . . . . 49

memory_object_supply_completed . . . . . . . . . . . . 53
memory_object_terminate . . . . . . . . . . . . . . . . . . . 51

N

natural_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

P

processor_assign . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

processor_basic_info_t . . . . . . . . . . . . . . . . . . . . 94

processor_control . . . . . . . . . . . . . . . . . . . . . . . . . . 92 processor_exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

processor_get_assignment . . . . . . . . . . . . . . . . . . 93

processor_info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

processor_set_basic_info_t . . . . . . . . . . . . . . . . 92

processor_set_create. . . . . . . . . . . . . . . . . . . . . . . 87 processor_set_default . . . . . . . . . . . . . . . . . . . . . 87 processor_set_destroy . . . . . . . . . . . . . . . . . . . . . 88 processor_set_info . . . . . . . . . . . . . . . . . . . . . . . . . 91

processor_set_max_priority . . . . . . . . . . . . . . . . 90
processor_set_policy_disable . . . . . . . . . . . . . . 90
processor_set_policy_enable . . . . . . . . . . . . . . . 90
processor_set_sched_info_t . . . . . . . . . . . . . . . . 92

processor_set_tasks. . . . . . . . . . . . . . . . . . . . . . . . 88 processor_set_threads . . . . . . . . . . . . . . . . . . . . . 88 processor_start . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

S

sample_pc_flavor_t . . . . . . . . . . . . . . . . . . . . . . . . . 78 sample_pc_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

seqnos_memory_object_change_completed . . . . 59
seqnos_memory_object_copy . . . . . . . . . . . . . . . . . 55
seqnos_memory_object_create . . . . . . . . . . . . . . . 59
seqnos_memory_object_data_initialize . . . . . 60
seqnos_memory_object_data_request. . . . . . . . . 52
seqnos_memory_object_data_return. . . . . . . . . . 52
seqnos_memory_object_data_unlock. . . . . . . . . . 57
seqnos_memory_object_default_server . . . . . . 49
seqnos_memory_object_init . . . . . . . . . . . . . . . . . 50
seqnos_memory_object_lock_completed . . . . . . 57
seqnos_memory_object_server . . . . . . . . . . . . . . . 49
seqnos_memory_object_supply_completed . . . . 54
seqnos_memory_object_terminate . . . . . . . . . . . . 51

struct host_basic_info . . . . . . . . . . . . . . . . . . . . . 82 struct host_sched_info . . . . . . . . . . . . . . . . . . . . . 82

struct processor_basic_info . . . . . . . . . . . . . . . 94
struct processor_set_basic_info . . . . . . . . . . . 91
struct processor_set_sched_info . . . . . . . . . . . 92

struct task_basic_info . . . . . . . . . . . . . . . . . . . . . 72

struct task_events_info . . . . . . . . . . . . . . . . . . . . 73
struct task_thread_times_info . . . . . . . . . . . . . 73
struct thread_basic_info. . . . . . . . . . . . . . . . . . . 62
struct thread_sched_info. . . . . . . . . . . . . . . . . . . 63

swtch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 swtch_pri . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

T

task_assign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 task_assign_default. . . . . . . . . . . . . . . . . . . . . . . . 89 task_basic_info_t . . . . . . . . . . . . . . . . . . . . . . . . . . 73 task_create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

task_disable_pc_sampling . . . . . . . . . . . . . . . . . . 77
task_enable_pc_sampling . . . . . . . . . . . . . . . . . . . 77

task_events_info_t . . . . . . . . . . . . . . . . . . . . . . . . . 73 task_get_assignment. . . . . . . . . . . . . . . . . . . . . . . . 89

task_get_bootstrap_port . . . . . . . . . . . . . . . . . . . 76
task_get_emulation_vector . . . . . . . . . . . . . . . . . 77
task_get_exception_port . . . . . . . . . . . . . . . . . . . 76

task_get_kernel_port. . . . . . . . . . . . . . . . . . . . . . . 76 task_get_sampled_pcs. . . . . . . . . . . . . . . . . . . . . . . 78 task_get_special_port . . . . . . . . . . . . . . . . . . . . . 75 task_info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 task_priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 task_ras_control . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

Function and Data Index 131

task_resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

task_set_bootstrap_port . . . . . . . . . . . . . . . . . . . 76

task_set_emulation . . . . . . . . . . . . . . . . . . . . . . . . . 77

task_set_emulation_vector . . . . . . . . . . . . . . . . . 77
task_set_exception_port . . . . . . . . . . . . . . . . . . . 76

task_set_kernel_port. . . . . . . . . . . . . . . . . . . . . . . 76 task_set_special_port . . . . . . . . . . . . . . . . . . . . . 76 task_suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 task_terminate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

task_thread_times_info_t . . . . . . . . . . . . . . . . . . 74

task_threads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 thread_abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 thread_assign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 thread_assign_default . . . . . . . . . . . . . . . . . . . . . 89 thread_basic_info_t. . . . . . . . . . . . . . . . . . . . . . . . 63 thread_create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 thread_depress_abort. . . . . . . . . . . . . . . . . . . . . . . 68

thread_disable_pc_sampling . . . . . . . . . . . . . . . . 77
thread_enable_pc_sampling . . . . . . . . . . . . . . . . . 77

thread_get_assignment . . . . . . . . . . . . . . . . . . . . . 89

thread_get_exception_port . . . . . . . . . . . . . . . . . 69
thread_get_kernel_port . . . . . . . . . . . . . . . . . . . . 69
thread_get_sampled_pcs . . . . . . . . . . . . . . . . . . . . 78
thread_get_special_port . . . . . . . . . . . . . . . . . . . 69

thread_get_state . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 thread_info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 thread_max_priority. . . . . . . . . . . . . . . . . . . . . . . . 67 thread_policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 thread_priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 thread_resume . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 thread_sched_info_t. . . . . . . . . . . . . . . . . . . . . . . . 64

                                                thread_set_exception_port . . . . . . . . . . . . . . . . . 70
                                                thread_set_kernel_port . . . . . . . . . . . . . . . . . . . . 70
                                                thread_set_special_port . . . . . . . . . . . . . . . . . . . 70

thread_set_state . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 thread_suspend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 thread_switch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 thread_terminate . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 thread_wire . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 time_value_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 time_value_add_usec. . . . . . . . . . . . . . . . . . . . . . . . 83 time_value_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

V

vm_allocate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 vm_copy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 vm_deallocate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 vm_inherit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 vm_machine_attribute. . . . . . . . . . . . . . . . . . . . . . . 45 vm_map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 vm_protect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 vm_read. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 vm_region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

                                                vm_set_default_memory_manager . . . . . . . . . . . . . 59

vm_statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 vm_statistics_data_t. . . . . . . . . . . . . . . . . . . . . . . 47 vm_wire. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 vm_write. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

132 The GNU Mach Reference Manual

Short Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Bootstrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4 Inter Process Communication . . . . . . . . . . . . . . . . . . . . . . 13 5 Virtual Memory Interface. . . . . . . . . . . . . . . . . . . . . . . . . 41 6 External Memory Management . . . . . . . . . . . . . . . . . . . . . 49 7 Threads and Tasks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 8 Host Interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 9 Processor Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 10 Device Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 11 Kernel Debugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Appendix A GNU GENERAL PUBLIC LICENSE . . . . . . . . . 111 Appendix B Documentation License . . . . . . . . . . . . . . . . . . . 119 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Function and Data Index . . . . . . . . . . . . . . . . . . . . . . . . . . . 129

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Audience . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Installing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.1 Binary Distributions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Cross-Compilation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3 Bootstrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.1 Bootloader. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2 Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Inter Process Communication . . . . . . . . . . . . . 13

4.1 Major Concepts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.2 Messaging Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

4.2.1 Mach Message Call . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2.2 Message Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.2.3 Exchanging Port Rights . . . . . . . . . . . . . . . . . . . . . . . . 20 4.2.4 Memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.2.5 Message Send . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.2.6 Message Receive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 4.2.7 Atomicity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

4.3 Port Manipulation Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.3.1 Port Creation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.3.2 Port Destruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 4.3.3 Port Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.3.4 Port Rights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.3.5 Ports and other Tasks. . . . . . . . . . . . . . . . . . . . . . . . . . 35 4.3.6 Receive Rights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.3.7 Port Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.3.8 Request Notifications . . . . . . . . . . . . . . . . . . . . . . . . . . 39

5 Virtual Memory Interface . . . . . . . . . . . . . . . . . 41

5.1 Memory Allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.2 Memory Deallocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.3 Data Transfer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.4 Memory Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.5 Mapping Memory Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.6 Memory Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

6 External Memory Management . . . . . . . . . . . . 49

6.1 Memory Object Server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 6.2 Memory Object Creation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 6.3 Memory Object Termination . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 6.4 Memory Objects and Data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 6.5 Memory Object Locking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 6.6 Memory Object Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.7 Default Memory Manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

7 Threads and Tasks . . . . . . . . . . . . . . . . . . . . . . . . 61

7.1 Thread Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.1.1 Thread Creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.1.2 Thread Termination . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.1.3 Thread Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.1.4 Thread Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.1.5 Thread Execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.1.6 Scheduling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

7.1.6.1 Thread Priority . . . . . . . . . . . . . . . . . . . . . . . 66 7.1.6.2 Hand-Off Scheduling . . . . . . . . . . . . . . . . . . 67 7.1.6.3 Scheduling Policy . . . . . . . . . . . . . . . . . . . . . 69

7.1.7 Thread Special Ports. . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.1.8 Exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

7.2 Task Interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.2.1 Task Creation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.2.2 Task Termination. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.2.3 Task Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.2.4 Task Execution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 7.2.5 Task Special Ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 7.2.6 Syscall Emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

7.3 Profiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

8 Host Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

8.1 Host Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 8.2 Host Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 8.3 Host Reboot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

9 Processor Interface . . . . . . . . . . . . . . . . . . . . . . . 87

9.1 Processor Set Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.2 Processor Set Creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 9.3 Processor Set Destruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.4 Tasks and Threads on Sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 9.5 Processor Set Priority . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 9.6 Processor Set Policy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 9.7 Processor Set Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 9.8 Processor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

10 Device Interface . . . . . . . . . . . . . . . . . . . . . . . . . 95

10.1 Device Reply Server . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 10.2 Device Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 10.3 Device Close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 10.4 Device Read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 10.5 Device Write. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 10.6 Device Map. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 10.7 Device Status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 10.8 Device Filter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

11 Kernel Debugger . . . . . . . . . . . . . . . . . . . . . . . 103

11.1 Operation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 11.2 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 11.3 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 11.4 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

Appendix A GNU GENERAL PUBLIC

LICENSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111

A.0.1 Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 A.0.2 TERMS AND CONDITIONS FOR COPYING,

DISTRIBUTION AND MODIFICATION . . . . . . . . 111

How to Apply These Terms to Your New Programs . . . . . . . . . . . 116

Appendix B Documentation License . . . . . . . . 119

B.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . 119

B.1.0.1 ADDENDUM: How to use this License for

your documents. . . . . . . . . . . . . . . . . . . . . . . . . . 125

B.2 CMU License. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

Function and Data Index . . . . . . . . . . . . . . . . . . . . 129